### v1

# FRNet: Frustum-Range Networks for Scalable LiDAR Segmentation

Abstract
LiDAR segmentation is crucial for autonomous driving systems. The recent range-view approaches are promising for real-time processing. However, they suffer inevitably from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement. In this work, we propose a simple yet powerful FRNet that restores the contextual information of the range image pixels with corresponding frustum LiDAR points. Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions. Next, a frustum-point fusion module is introduced to update per-point features hierarchically, which enables each point to extract more surrounding information via the frustum features. Finally, a head fusion module is used to fuse features at different levels for final semantic prediction. Extensive experiments on four popular LiDAR segmentation benchmarks under various task setups demonstrate our superiority. FRNet achieves competitive performance while maintaining high efficiency. The code is publicly available.
1 Introduction
Figure 1: Scalability comparisons among state-of-the-art LiDAR segmentation models on the SemanticKITTI [5] leaderboard. The size of the circular represents the number of model parameters. Our FRNet achieves competitive performance with current arts while maintaining satisfactory efficiency for real-time processing.
Real-time and accurate LiDAR semantic segmentation is crucial in autonomous driving, robotics, and other safety-critical applications [18]. Many recent approaches have made significant progress in this area, which represent LiDAR point clouds from different perspectives, including point view [48, 21], voxel view [47, 65], pseudo-image view [50, 35, 61], and multi-view representation [31, 54, 2], depending on different pre-processing techniques.
Point-view methods manipulate the origin point cloud to preserve raw spatial information, following the paradigm of PointNet++ [40]. Nevertheless, they require computationally expensive neighborhood searches to construct local neighborhoods, and the computation overhead is considerable for large scene point clouds. Voxel-view methods transform scatter points into regular voxel representations and apply 3D convolutions for label prediction, but the computation grows cubically when voxel resolution increases. Multi-view methods extract features from different representations separately and fuse these features from different levels for more accurate prediction. Consequently, the computation of all representations is also gathered. Albeit these representation methods achieve some satisfactory performance, their computation overhead hinders them from real-time online applications.
Recently, pseudo-image methods propose a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [61, 1] and range-view [50, 51, 43, 4, 24]. Under such representations, some popular image segmentation approaches [33, 62] can be directly applied to such pseudo-images to balance the efficiency and accuracy of segmentation. However, the 3D-to-2D projection procedure tends to result in corrupted neighborhood contextual information, decreasing the precision. Detailed analysis of this problem can be summarized into two aspects.
Firstly, some 3D points are discarded in the projection procedure. Under such a paradigm, only the projected points are involved in convolutions, while the contextual information of discarded points is absent. Consequently, the pixel in the pseudo-image can hardly preserve the high-fidelity 3D geometric context. More importantly, these discarded 3D points still need to predict 3D labels but no features correspond to them. To reconstruct these labels, some post-processing techniques, such as CRF [22] and KNN [35], are used for these points based on 2D predictions. However, these techniques are unsupervised and fail to extract the local features of discarded points in a learnable manner, which significantly limits the improvement of pseudo-image methods.
Secondly, the pseudo-image suffers from many empty pixels due to the sparsity of LiDAR points. Under such a non-compact representation, the projected points will extract much noisy information as the convolutions go deeper. In summary, pseudo-image methods fail to preserve the contextual information of all the points, which is significant for LiDAR segmentation. So, it is important for points and pseudo-images to cooperate with each other. Although existing multi-view methods have successfully preserved the context of all points via fusing features of different representations, they are always computationally expensive.
To this end, we study a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which can preserve the contextual information of all points through the cooperation of point and pseudo-image and still maintain high efficiency. Specifically, FRNet comprises three main components. To preserve all points and predict semantic labels for these points in an end-to-end manner, a Frustum Feature Encoder is used to group all points with the same frustum region into the corresponding range-view pixel in relation to the range image via several multi-layer perceptions. The point features can be pooled to represent the frustum region and formatted into a 2D representation, followed by traditional convolutions. Then, a Frustum-Point (FP) Fusion Module is introduced to update the features of each point hierarchically during each stage of the convolutions, which consists of frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features. In this way, all the points extract larger local features based on the frustum region. Finally, a Fusion Head module is designed to utilize different level features to generate individual features for each point, so that the results can be predicted end-to-end without any post-processing techniques. As shown in Fig. 1, the proposed FRNet achieves great improvements among range-view methods while still maintaining high efficiency.
To alleviate empty pixels in 2D representation, we introduce a new range interpolation augmentation method that adds several potential points based on the information of surrounding points. Specifically, we first transform the point clouds into range images, as done in previous works. For empty pixels in the range images, we then calculate their potential coordinates when both the nearest left and right pixels have projected points as they often share similar attributes. With this augmentation, the 2D representation will have fewer empty pixels, allowing convolutions to learn more semantically coherent information.
The key contributions are summarized as follows:
  • We introduce a novel Frustum-Range representation for LiDAR segmentation. In this representation, each pixel in the 2D representation is treated as a corresponding frustum region in the point clouds. All points within the same frustum region are preserved for semantic prediction without the need for any post-processing techniques.
  • We propose a novel range interpolation augmentation that synthesizes possibly lost LiDAR points based on surrounding information, helping to fill empty pixels and learn more coherent range-view representations.
  • Extensive experiments across four LiDAR segmentation benchmarks demonstrate that FRNet can achieve the optimal trade-off between efficiency and accuracy.
2 Related Work
Scalable LiDAR Segmentation. Although Point-view [21, 48], voxel-view [47, 65], and multi-view [31, 32] have achieved great success in LiDAR segmentation, they are often time-consuming and cannot achieve real-time processing. Recently, various methods proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation. SqueezeSeg [50] and SqueezeSegV2 [51] used the lightweight model SqueezeNet [22] to retain information. SqueezeSegV3 [53] introduced a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images. RangeNet++ [35] integrated Darknet into SqueezeSeg [50] and proposed an efficient KNN post-processing for segmentation. RangeViT [4] and RangeFormer [24] introduced popular transformer blocks to extract local and global information of the range images. These methods only learn the features of projected points while ignoring the occluded points.
LiDAR Post-Processing. Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results. Many of these methods follow the CRF [50] or KNN [35] approach to smooth the semantic results of occluded points. FIDNet [64] proposed a Nearest Label Assignment (NLA) method for label smoothing. Both of these techniques are unsupervised and can not extract the features of all points in a learnable manner. RangeFormer [24] created sub-clouds from full point clouds and inferred each subset in a supervised manner, but it still relies on 2D prediction results. In contrast, our method involves the frustum region and predicts the semantic label of each point directly without any information loss, making it a fully end-to-end approach.
LiDAR Data Augmentation. Recent works have explored different data augmentation techniques on point clouds [7, 60], which can help improve performance when it comes to indoor scene understanding but has limited generalization to outdoor scenes. Mix3D [36] created new training samples by combining two scenes out of context. PolarMix [52] proposed mixing two LiDAR scenes based on yaw angles and cropping instances from one scene to another, while LaserMix [26] proposed exchanging labeled scans with unlabeled scans along the pitch axis for efficient semi-supervised learning. RangeFormer [24] presented a lightweight augmentation method that only operated on projected points in range images, disregarding occluded points. In this work, we prevent a novel augmentation technique based on our proposed frustum region, which preserves all points within the frustum region, encompassing techniques such as FrustumMix and Range-Interpolation.
Figure 2: A pilot study on the performance degradation of post-processing in existing range-view methods [35, 12, 64, 8] on SemanticKITTI [5]. We choose various $K$ values as hyperparameters in KNN post-processing. Compared to their performance at 2D, a severe drop in performance occurs with different $K$ values.
3 FRNet: A Scalable LiDAR Segmentor
3.1 Pilot Study
Figure 3: Architecture overview. The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features in the frustum region. 2) Frustum-Point (FP) Fusion Module is used to update per-point features hierarchically at each stage of the 2D backbone. 3) Fusion Head fuses different levels of features to predict final results.
LiDAR segmentation aims to assign semantic labels to each point in the point clouds. A common approach is to perform operations directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [40]. However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes. To address this issue, many recent works propose projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency. Given point clouds with coordinates $(x,y,z)$ and the expected projected resolution $(H,W)$ of range images, each point can be mapped to its correspondence coordinates $(u,v)$ in the range image via the following transformation:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (1)
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. However, such a projection process can cause the loss of occluded points and often requires post-processing techniques to reduce the blurry prediction of lost points [24].
We conduct a pilot study on different range-view methods, including RangeNet++ [35], SalsaNeXt [12], FIDNet [64], and CENet [8], and compare their performance with different KNN post-processing parameters and 2D performance. As shown in Fig. 2, the post-processing procedures are often unsupervised, and the hyperparameters can significantly influence the final performance. To address this issue, we propose a scalable Frustum-Range Network (FRNet) that can directly predict semantic labels for each point while maintaining high efficiency.
3.2 Frustum-Range Architecture
Given a point cloud $\mathbf{P}$ acquired by the LiDAR sensor, FRNet is expected to learn a feed-forward network $\Phi$ to predict semantic labels $\mathbf{Y}$ for each point:
$\mathbf{Y}=\Phi(\mathbf{P},\theta)~{},$ (2)
where $\theta$ denotes the learnable parameters in the network. As shown in Fig. 3, FRNet comprises a Frustum Feature Encoder that carries out per-point feature extraction, a Frustum-Point Fusion Module that fuses per-point features and frustum features hierarchically, and a Fusion Head Module that predicts semantic labels for each point.
Frustum Feature Encoder. LiDAR segmentation involves predicting semantic labels for each point, so it is important to extract individual features for each point. Before converting point clouds into 2D representations, Frustum Feature Encoder first extracts per-point features via several MLPs, which is crucial for accurate label prediction. For points $\mathbf{P}_{i}=\{p_{i}^{1},p_{i}^{2},...,p_{i}^{N}\}$ in the $i$-th frustum region, the points are manually embedded with $(p_{i}^{k},p_{i}^{k}-p_{i}^{mean})$, where $p_{i}^{k}$ is the $k$-th point in the frustum, $N$ is the total number of points in the frustum region, and $p_{i}^{mean}=\frac{1}{N}\sum_{k=1}^{N}p_{i}^{k}$ represents the mean information of the frustum region. The per-point features are then obtained as follows:
$f_{i}^{k}=\text{MLP}([p_{i}^{k};p_{i}^{mean}])~{},$ (3)
[…]
Frustum-Point Fusion Module. As shown in Fig. 4, the Frustum-Point (FP) Fusion Module comprises two components: 1) A Frustum-to-Point fusion that hierarchically updates per-point features, and 2) A Point-to-Frustum fusion to fuse individual features of each point within its corresponding frustum region. Specifically, for the point features $F_{p}$ and frustum features $F_{f}$, $F_{f}$ is back-project to the point-level features $F_{p}^{f}$ based on the projection index. This means that points within the same frustum region share the same frustum features. Then, the two features are concatenated together, followed by an MLP to update per-point features:
$\hat{F}_{p}=\text{MLP}([F_{p}^{f};F_{p}])~{}.$ (4)
Next, the updated per-point features $\hat{F}_{p}$ are pooled into the corresponding frustum region to further enhance the frustum features $F_{f}$ in a residual-attentive manner. The pooled features are denoted as $F_{f}^{p}$, which are calculated with the max operation based on the frustum region. The pooled features $F_{f}^{p}$ are first concatenated with $F_{f}$, followed by a simple convolutional layer to reduce the number of channels:
$F^{fuse}=g([F_{f}^{p};F_{f}],\theta_{1})~{},$ (5)
where $g(\cdot)$ is the mapping function and $\theta_{1}$ is the learnable parameters. Finally, the fused features are used to further enhance the frustum features in a residual-attentive manner:
$\hat{F}_{f}=F_{f}+\sigma(h(F^{fuse},\theta_{2}))\odot F^{fuse}~{},$ (6)
where $\sigma$ represents the sigmoid function, $h$ is the mapping function, $\theta_{2}$ is the learnable parameters, and $\odot$ indicates element-wise multiplication. The output features $\hat{F}_{f}$ are then used to extract larger local frustum features with subsequent convolutions.
Fusion Head Module. Since the aim of FRNet is to classify categories for each point, the most straightforward method is to directly use the updated per-point features after the FP Fusion Module for prediction. Although this can also achieve great performance, we observe that this is not the optimal solution. The output of our FP Fusion Module consists of two features: per-point features and frustum features represented in a 2D manner, and our prediction head expects to fully utilize both of these two features. Additionally, the per-point features generated from the Frustum Feature Encoder Module directly embed coordinate information, which can also provide more low-level geometric information. The Fusion Head Module progressively fuses three different level features. Firstly, the frustum features $F_{f}$ are mapped to point-level features $F_{f}^{p}$. It is worth noting that points in the same frustum regions share the same features at this stage. Then, we map the features $F_{f}^{p}$ with MLPs to ensure they share the same channels as point features. Finally, the add operation is applied to fuse the two features. The same operation is also used to fuse with low-level point features $F_{p}^{l}$. The overall operation can be formulated as:
$F_{out}=\text{MLP}(\text{MLP}(F_{f}^{p})+F_{p})+F_{p}^{l}~{}.$ (7)
Here, $F_{out}$ is used to generate the final semantic scores for each point.
Optimization. To improve the optimization of FRNet, we not only utilize the commonly used point-level supervised loss function but also supervise the frustum-level features with generated pseudo frustum labels. Typically, a frustum region may contain multiple points, and these points often do not belong to the same category. Based on this observation, we generate a pseudo label for each frustum in a statistical manner. We count the frequency of each category in the frustum region and choose the category with the highest frequency as the pseudo label for the frustum. Finally, the frustum-level features will be supervised using the generated pseudo labels. The overall loss can be calculated as:
$L=L_{p}+\lambda L_{f}~{},$ (8)
where $L_{p}$ and $L_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision. $L_{p}$ is calculated by cross-entropy loss and $L_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42].
Figure 4: Frustum-Point fusion module comprises two steps: 1) A Frustum-to-Point fusion to update per-point features. 2) A Point-to-Frustum fusion to update frustum features.
3.3 Frustum-Range Operators
Data augmentation enables the model to learn more robust representations. Commonly used techniques in LiDAR segmentation often operate on the global view, including rotation, flipping, scaling, jittering, etc. In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
FrustumMix. Different from the previous LaserMix [26] and PolarMix [52] augmentations which mix two LiDAR scenes in the point level, we propose a FrustumMix to mix two scenes based on frustum units. Points in the scan will be first grouped into different frustum regions along inclination and azimuth directions. We then randomly split frustum regions equally from two directions and switch the corresponding frustum regions from two scans.
[…]
Method Venue Representation Param SemanticKITTI nuScenes ScribbleKITTI SemanticPOSS
FPS normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$
MinkUNet [10] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M
$\mathbf{9.1}$
$63.7$
$55.0$
$\mathbf{53.1}$
$\mathbf{68.3}$
SPVNAS [47] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$
$\underline{66.4}$
$\underline{56.9}$
$48.4$
$61.5$
Cylinder3D [65] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M
$6.2$
$\mathbf{65.9}$
$\mathbf{67.8}$
$\mathbf{76.1}$
$\mathbf{77.9}$
$\mathbf{57.0}$
$\underline{52.9}$
$\underline{64.9}$
RPVNet [54] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$
$65.5$
$70.3$
$77.6$
2DPASS [56] ECCV’22 Multi-View $\bullet$ $26.5$ M
$\mathbf{8.4}$
$69.3$
$72.2$
$\mathbf{79.4}$
$80.8$
SphereFormer [27] CVPR’23 Multi-View $\bullet$ $32.3$ M
$4.9$
$67.8$
$74.8$ $78.4$
$\underline{81.9}$
UniSeg [32] ICCV’23 Multi-View $\bullet$ $147.6$ M
$6.9$
$\mathbf{71.3}$
$\mathbf{75.2}$
$\mathbf{83.5}$
RangeNet++ [35] IROS’19 Range View $\bullet$ $50.4$ M
$12.0$
$50.3$
$52.2$
$65.5$
$44.6$
$57.8$
$30.9$
FIDNet [64] IROS’20 Range View $\bullet$ $6.1$ M
$31.8$
$58.9$
$59.5$
$54.1$
$65.4$
$46.4$
CENet [8] ICME’22 Range View $\bullet$ $6.8$ M $33.4$
$62.6$
$64.7$
$55.7$
$66.8$
$50.3$
RangeViT [4] CVPR’23 Range View $\bullet$ $23.7$ M
$10.0$
$60.7$
$64.0$
$75.2$
$53.6$
$66.5$
RangeFormer [24] ICCV’23 Range View $\bullet$ $24.3$ M
$6.2$
$67.6$
$\mathbf{73.3}$
$78.1$
$80.1$
$63.0$
Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M
$\mathbf{33.8}$
$67.1$
$\underline{72.5}$
$78.8$ $82.1$
$62.4$
$71.2$ $52.5$ $67.1$
FRNet Ours Frustum-Range $\bullet$ $10.0$ M
$29.1$
$\mathbf{68.7}$
$\mathbf{73.3}$
$\mathbf{79.0}$
$\mathbf{82.5}$
$\mathbf{63.1}$
$\mathbf{72.3}$
$\mathbf{53.5}$
$\mathbf{68.1}$
[…]
Method Venue Backbone SemanticKITTI nuScenes ScribbleKITTI
$\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$ $\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$ $\mathbf{1\%}$ $\mathbf{10\%}$ $\mathbf{20\%}$ $\mathbf{50\%}$
Sup.-only - FIDNet [64] $36.2$ $52.2$ $55.9$ $57.2$ $38.3$ $57.5$ $62.7$ $67.6$ $33.1$ $47.7$ $49.9$ $52.5$
LaserMix [26] CVPR’23 $43.4$ $58.8$ $59.4$ $61.4$ $49.5$ $68.2$ $70.6$ $73.0$ $38.3$ $54.4$ $55.6$ $58.7$
Sup.-only - Cylinder3D [65] $45.4$ $56.1$ $57.8$ $58.7$ $50.9$ $65.9$ $66.6$ $71.2$ $39.2$ $48.0$ $52.1$ $53.8$
CRB [49] CVPR’22 - $58.7$ $59.1$ $60.9$ $54.2$ $56.5$ $58.9$
LaserMix [26] CVPR’23 $50.6$ $60.0$ $61.9$ $62.3$ $55.3$ $69.9$ $71.8$ $73.2$ $44.2$ $53.7$ $55.1$ $56.8$
LiM3D [29] CVPR’23 - $61.6$ $62.6$ $62.8$ $\mathbf{60.3}$ $\mathbf{60.5}$ $60.9$
ImageTo360 [44] ICCVW’23 $54.5$ $58.6$ $61.4$ $64.2$
FrustumMix Ours $55.7$ $62.5$ $63.0$ $64.9$ $60.0$ $70.0$ $72.6$ $74.1$ $45.6$ $55.7$ $58.2$ $60.8$
Sup.-only - FRNet (Ours) $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
LaserMix [26] CVPR’23 $52.9$ $62.9$ $63.2$ $65.0$ $58.7$ $71.5$ $72.3$ $75.0$ $45.8$ $56.8$ $57.7$ $59.0$
FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $\mathbf{65.2}$ $\mathbf{65.4}$ $\mathbf{61.2}$ $\mathbf{72.2}$ $\mathbf{74.6}$ $\mathbf{75.4}$ $\mathbf{46.6}$ $57.0$ $59.5$ $\mathbf{61.2}$
Range-Interpolation. Previous range-view methods often suffer from large number of empty pixels, which still exist in FRNet. To alleviate this problem, we propose to interpolate some potential points based on the range image. Specifically, LiDAR points will be first projected to range images as previous works do. We then scan all pixels in the range images and find the empty pixels. For each empty pixel, we will judge whether there are projected points in the pixels along its left and right directions. When both the two pixels have projected points, the interpolated points of this pixel can be calculated by averaging the coordinates of two points. During the training procedure, we also need to generate semantic labels for the interpolated points, which will face two cases. If both right and left points have the same semantic labels, we will set the semantic label of the interpolated point to be the same as them. For another case with different semantic labels, we consider the interpolated points as boundaries of the object, and it will be set to an ignored label, which will not be used to calculate loss.
4 Experiments
4.1 Settings
Datasets. We conduct experiments on four popular LiDAR segmentation datasets. SemanticKITTI [5] consists of $22$ sequences with a total of $43,552$ LiDAR scans. Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing with $19$ semantic classes. nuScenes [13] provides $1,000$ driving scenes captured with much sparser points by 32-beam LiDAR sensors, and 16 classes are evaluated during inference. SemanticPOSS [37] is a more challenging benchmark collected by Peking University with much smaller and sparser point clouds. It consists of $2,988$ LiDAR scenes with a large number of dynamic instances. Officially, it is divided into $6$ parts, and part $2$ is used to evaluate the performance with $13$ categories. ScribbleKITTI [49] shares the same scenes with SemanticKITTI [5] but only provides weak annotations with line scribbles, which means around $8.06\%$ labeled points available during training.
Evaluation Metrics. In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to access the performance. IoU and Acc can be calculated as:
$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},~{}~{}\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (9)
where $\text{TP}_{i},\text{TP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively. For robustness evaluation, we follow the practice of Robo3D [25] to use Corruption Error (CE) and Resilience Rate (RR) to evaluate the robustness of FRNet. The CE and RR can be calculated as:
$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},~{}~{}\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (10)
where $\text{IoU}_{i,l}$ is the task-specific IoU; $\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ are scores of the baseline model and scores on the clean evaluation set.
Implementation Details. We implement FRNet using the MMDetection3D [11] framework and utilize CENet [8] as our backbone to extract frustum features. We set the shape of the 2D representation to $64\times 512$ for SemanticKITTI [5] and ScribbelKITTI [49] and $32\times 480$ for nuScenes [13] and SemanticPOSS [37], respectively. We choose AdamW [34] as our default optimizer and utilize the OneCycle scheduler [45] with $lr=0.01$. The models are trained for $50$ epochs on SemanticKITTI [5] and ScribbelKITTI [49], and $80$ epochs on nuScenes [13] and SemanticPOSS [37], respectively, with a batch size of $2$ with $4$ GPUs.
4.2 Comparative Study
Comparisons to State of the Arts. Firstly, we compare the proposed FRNet with state-of-the-art LiDAR segmentation methods from different representations on both fully-supervised and weakly-supervised benchmarks. Tab. 1 summarizes all the results and the FPS of all the models is evaluated with a single GeForce RTX 2080Ti GPU for fair comparison. Fast-FRNet takes resnet18 [19] as the 2D backbone and the shapes of 2D representation on four datasets are $32\times 360$. Compared with multi-view methods such as SphereFormer [27] and UniSeg [32], although FRNet does not have a significant advantage in performance, its parameters and FPS are far superior to theirs. Compared with voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency, and Fast-FRNet achieves the fastest inference speed with only a little drop in performance. Fig. 6 also shows qualitative results among state-of-the-art methods, including CENet [8] and RangeViT [4] on the val set of SemanticKITTI [5]. FRNet reduces prediction errors over large areas, especially for vegetation and parking.
Label-Efficient LiDAR Segmentation. Recently, semi-supervised learning has been explored in LiDAR segmentation. In this work, we follow the paradigm of LaserMix [26] to conduct FRNet on the semi-supervised segmentation task. Here, we use FRNet as the backbone and the proposed FrustumMix to mix labeled and unlabeled. Tab. 2 summarizes the results of recent methods. Results show that our FRNet backbone achieves appealing performance among state-of-the-art methods. Meanwhile, we compare the proposed FrustumMix with LaserMix among different backbones, including Cylinder3D [65] and FRNet. Results show that the proposed FrustumMix outperforms LaserMix, even in voxel representation.
Robustness. Robustness reflects the model’s ability to generalization to a certain extent, which is particularly important in autonomous driving, especially in extreme weather conditions. To this end, we also evaluate the robustness of FRNet using the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [25], which include 8 different conditions and each condition has 3 different levels. The models trained with SemanticKITTI [5] and nuScenes [13] are used to evaluate the performance on SemanticKITTI-C and nuScenes-C [25], respectively. As shown in Tab. 3, FRNet demonstrates strong robustness, especially under extreme weather conditions.
[…]
Method Venue SemKITTI-C nuScenes-C
mCE normal-$\downarrow$ mRR normal-$\uparrow$ mCE normal-$\downarrow$ mRR normal-$\uparrow$
MinkUNet [10] CVPR’19 $100.0$
$\mathbf{81.9}$
$100.0$
$74.4$
Cylinder3D [65] CVPR’21
$103.3$
$80.1$
$111.8$
$72.9$
2DPASS [56] ECCV’22
$106.1$
$77.5$
$\mathbf{98.6}$
$75.2$
FIDNet [64] IROS’20
$113.8$
$77.0$
$122.4$
$73.3$
CENet [8] ICME’22
$103.4$
$81.3$
$112.8$
$76.0$
FRNet Ours
$\mathbf{96.8}$
$80.0$
$\mathbf{98.6}$
$\mathbf{77.5}$
Figure 5: Ablation study on different configurations in FRNet. Tables show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [5]. b) Different mixing strategies on the val set of nuScenes [13]. c) Different mixing strategies on the val set of SemanticKITTI [5]. Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c).
Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [8, 4]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
4.3 Ablation Study
In this section, we will discuss each component in FRNet on the val set of SemanticKITTI [5] and nuScenes [13].
Model Analysis. We first analyze different components in FRNet, including the Frustum Feature Encoder, Frustum-Point (FP) Fusion, Fusion Head, Frustum-level Supervision, and Range-Interpolation. The results are summarized in Tab. 4. As the baseline of FRNet, the Frustum Feature Encoder module extracts per-point features in the frustum region and predicts semantic labels of each point with the supervision of point-level labels, and the performance is $62.7\%$ and $72.8\%$ on SemanticKITTI [5] and nuScenes [13], respectively. When we use the FP Fusion module to update per-point features hierarchically, the performance improves by $1.6\%$ and $0.8\%$ on the two datasets, respectively. With the frustum-level supervision, the performance further improves by $1.9\%$ and $2.8\%$. Next, we utilize a Fusion Head module to fuse different level features for predicting, and it improves by $0.8\%$ on both datasets. To alleviate the empty pixels in 2D representations, we propose a Range-Interpolation method to generate some potential points. Results show that this augmentation can also help to improve the performance by $0.6\%$ and $0.5\%$. Finally, following most works, we also use Test Time Augmentation (TTA) during inference, and it will bring an improvement of $1.1\%$ and $1.3\%$, respectively.
[…]
FFE FP FH FS RI TTA SemKITTI nuScenes
mIoU mAcc mIoU mAcc
$62.7$
$72.3$
$72.8$
$82.7$
$64.3$
$73.5$
$73.6$
$83.4$
$66.2$
$73.9$
$76.4$
$84.6$
$67.0$
$74.1$
$77.2$
$84.8$
$67.6$
$74.2$
$77.7$
$\mathbf{85.2}$
$\mathbf{68.7}$
$\mathbf{74.9}$
$\mathbf{79.0}$
$85.0$
Mixing Strategy. Although we have validated the effectiveness of our proposed FrustumMix in semi-supervised segmentation, we still compare FrustumMix with different popular outdoor mix strategies, including LaserMix [26], PolarMix [52]. As shown in Fig. 5 (b) and (c), the proposed FrustumMix significantly improves the performance and outperforms existing mixing strategies on both datasets.
Range-Interpolation Strategy. Range-Interpolation synthesizes possible points to reduce empty pixels in 2D representations. For the empty pixel in the range image, we initially use a $3\times 3$ window to pick its neighboring pixels and average valid pixel information as the interpolated point. Unfortunately, this results in a $0.4\%$ drop in performance. When we simply average the effective pixels, the generated points might fall out of the expected frustum region due to the empty surroundings. Based on this observation, we re-think the interpolation strategy. For an empty pixel, only when there are projection points on both the left and right or above and below, the calculated interpolation point can fall in the expected region. To this end, we propose two interpolation strategies along horizontal and vertical directions and improve the results by $0.4\%$ and $0.6\%$, respectively. It seems that points along horizontal directions share similar attributes as they line in the same laser beams.
Backbones. To validate the versatility of our method, we select different range-view methods as our 2D backbone to extract frustum features, including SalsaNeXt [12], FIDNet [64], and CENet [8]. As shown in Fig. 5 (a), our proposed frustum-range representation can bring an improvement of $5.0\%$ to $6.8\%$ on different range-view methods.
5 Conclusion
In this work, we presented a scalable Frustum-Range network (FRNet) for LiDAR segmentation, which can be seamlessly integrated into any range-view method. FRNet comprises three main components: the Frustum Feature Encoder Module to extract per-point features in low-level, the Frustum-Point Fusion Module to update per-point features hierarchically, and the Fusion Head Module to fuse different level features to predict more accurate labels. We also introduced two efficient augmentations, FurstumMix and Range-Interpolation, to improve the generalization of FRNet. Extensive experiments showed that our FRNet achieves competitive performance while maintaining high efficiency.
Appendix
[…]
Appendix A Additional Implementation Details
In this section, we provide additional details regarding the implementation and reproduction details of the proposed FRNet and Fast-FRNet.
A.1 Datasets
We adopt the following four LiDAR segmentation datasets to validate the effectiveness and superiority of the proposed approach.
  • SemanticKITTI [5] is a large-scale outdoor dataset which is based on KITTI odometry benchmark [14]. It consists of $22$ sequences, collected from different scenes in Karlsruhe, Germany by Velodyne HDL-64E. Commonly, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) is used for validation, and sequences $11$ to $21$ ($20,351$ scans) are used for testing. The original dataset is annotated with $28$ classes and only $19$ semantic classes are used for single-scan evaluation. The vertical field-of-view is $[-25^{\circ},3^{\circ}]$.
  • nuScenes [13] is a multimodal dataset that is widely used in autonomous driving. The LiDAR point clouds are collected around the streets in Boston and Singapore by Velodyne HDL-32E. It comprises $1,000$ driving scenes, in which $700$ scenes ($28,130$ scans) for training, $150$ scenes ($6,019$ scans) for validation, and $150$ scenes ($6,008$ scans) for testing. The dataset is annotated with $32$ classes and only $16$ semantic classes are used for evaluation. The vertical field-of-view is $[-30^{\circ},10^{\circ}]$.
  • ScribbleKITTI [49] is a scribble annotated dataset with line drawings, which shares the same LiDAR point clouds with SemanticKITTI [5]. It contains $189$ million labeled points, around $8.06\%$ of the total training points. This dataset is suitable for conducting weakly-supervised LiDAR segmentation, which is useful for reducing the annotation cost in existing LiDAR segmentation systems.
  • SemanticPOSS [37] is a more challenging dataset collected in Peking University by Pandora. The dataset includes $6$ sequences. Commonly, sequence $02$ ($500$ scans) is used for validation, and others ($2,488$ scans) are used for training. The dataset is annotated with $14$ classes and only $13$ semantic classes are used for evaluation. The vertical field-of-view is $[-16^{\circ},7^{\circ}]$.
A.2 Hyperparameters
For each point with coordinate $(x,y,z)$, its corresponding frustum region $(u,v)$ in the 2D representation with shape of $(H,W)$ is calculated as:
$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (11)
where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively. Points with the same $(u,v)$ are grouped into the same frustum region. The hyperparameter configuration of each module in our proposed FRNet framework is summarized as follows.
  • Frustum Feature Encoder. The input of frustum feature encoder for each point is $(x,y,z,i,d,x-x_{m},y-y_{m},z-z_{m})$, where $(x_{m},y_{m},z_{m})$ is the mean coordinate in the corresponding frustum region. The features are fed into four MLPs that map the channel from $8$ to $64$, $128$, $256$, and $256$, respectively, with batch normalization and ReLU function on the first three layers. Then the point features are pooled into frustum features, which are fed into an MLP that maps the channel from $256$ to $16$ with a ReLU function.
  • 2D Backbone. We adopt a variant of ResNet18 and ResNet34 used in CENet [8] as the backbone of Fast-FRNet and FRNet, respectively. The spatial shape of each stage is $[(H,W),(\frac{H}{2},\frac{W}{2}),(\frac{H}{4},\frac{W}{4}),(\frac{H}{8},\frac{W}{8})]$ and the output channel is $[128,128,128,128]$. The four features are interpolated to the shape of $(H,W)$ for feature concatenation, with two convolutional layers to reduce the channel dimension to $[256,128]$ progressively.
  • Frustum-Point Fusion Module. Point features are updated hierarchically at each stage of the 2D backbone. Accordingly, the channel dimension of point features at each stage is $[128,128,128,128]$. We also concatenate point features at each stage and reduce the channel dimension to $[256,128]$ with two MLPs progressively.
  • Optimization. For point-level supervision, we use the cross-entropy loss. For frustum-level supervision, we use cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42], and the weights for each loss are $1.0$, $1.5$, and $1.0$, respectively. The frustum-level loss is also used at each stage of the 2D backbone to supervise the middle frustum features. The weights for point-level loss and frustum-level loss are $1.0$ and $1.0$, respectively.
Algorithm 1 FrustumMix, NumPy-style
# lidar_list: the whole dataset.
# h: height of the 2D representation.
# w: width of the 2D representation.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# num_areas: List of split regions along inclination direction. We set [3, 4, 5, 6] here.
# sample: function to sample the lidar points and labels from the whole dataset.
# group: function to calculate frustum regions for points, refer to Eq. (1).
def FrustumMix(xa, ya, xb, yb):
    # group frustum regions
    ua, va = group(xa, h, w, f_u, f_d)
    ub, vb = group(xb, h, w, f_u, f_d)
    out_points = []
    out_labels = []
    if np.random.rand() > 0.5:
        # mix two scans along inclination direction
        n = np.random.choice(num_areas, size=1)[0]
        inc_list = np.linspace(0, h, n+1, dtype=np.int64)
        for i in range(n):
            start = inc_list[i]
            end = inc_list[i+1]
            if i % 2 == 0:
                idx = (va>=start) & (va<end)
                out_points.append(xa[idx])
                out_labels.append(ya[idx])
            else:
                idx = (vb>=start) & (vb<end)
                out_points.append(xb[idx])
                out_labels.append(yb[idx])
    else:
        # mix two scans along azimuth direction
        start = np.random.randint(0, w//2)
        end = start + w//2
        idxa = (ua<start) | (ua>=end)
        idxb = (ub>=start) & (ub<end)
        out_points = [xa[idxa], xb[idxb]]
        out_labels = [ya[idxa], yb[idxb]]
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
# sample two sans
idx_a = np.random.randint(0, len(lidar_list))
idx_b = np.random.randint(0, len(lidar_list))
points_a, labels_a = sample(idx_a)
points_b, labels_b = sample(idx_b)
# mix the two scans
points, labels = FrustumMix(points_a, labels_a, points_b, labels_b)
Algorithm 2 Range-Interpolation, NumPy-style
# h: height of the range image.
# w: width of the range image.
# f_u: upward pitch angles of the field-of-view.
# f_d: downward pitch angles of the field-of-view.
# points: lidar points with shape [m, c].
# labels: semantic labels with shape: [m,].
# ignore_label: ignored label that is not used to calculate loss.
# laser_scan: function to convert LiDAR points to range image.
def RangeInterpolation(p, l):
    # convert to range images
    # r_m means the valid mask for each pixel. 0 is invalid, while 1 is valid.
    r_p, r_l, r_m = laser_scan(p, l, h, w, f_u, f_d)
    out_points = [p]
    out_labels = [l]
    for y in range(h):
        for x in range(w):
            if r_m[y, x]:
                continue
            if (x-1>=0) and (x+1<w):
                if r_m[y, x-1] and r_m[y, x+1]:
                    # both right and left pixels are valid
                    point = (r_p[y, x-1]+r_p[y, x+1]) / 2
                    if r_l[y, x-1] == r_l[y, x+1]:
                        label = r_l[y, x-1]
                    else:
                        label = ignore_label
                    out_points.append(point)
                    out_labels.append(label)
                    r_p[y, x] = point
                    r_l[y, x] = label
                    r_m[y, x] = 1
    out_points = np.concatenate(out_points)
    out_labels = np.concatenate(out_labels)
    return out_points, out_labels
points, labels = RangeInterpolation(points, labels)
A.3 Training Configurations
The common augmentation methods, including random dropping, random flipping, global rotation, global scale, global translation, and instance copy, are used in FRNet. Additionally, we also use the following proposed augmentation methods to further improve the performance.
  • FrustumMix. Based on the calculated frustum region $(u,v)$, the two scans are randomly mixed along inclination and azimuth directions. For inclination direction, the frustum regions are randomly split into several subsets sampled from $[3,4,5,6]$, while for azimuth direction, the frustum regions are split into $2$ subsets.
  • Range-Interpolation. Range-Interpolation aims to synthesize some possible points based on range images, which helps to fill empty pixels in the 2D representations. We first scan all the pixels in the range image. For each empty pixel, the interpolated points are calculated only when its right and left pixels both have projected points.
During the validation and testing phase, only the Range-Interpolation is kept to fill the empty pixels in 2D representations. To make the final results correspond to the original points, we set the first $N$ labels as the predicted labels as the last multiple points are interpolated points. Moreover, we apply the test time augmentation to improve the performance. Specifically, the input points are flipped along different directions, with randomly rotated, scaled, and jittered. The total number of voted point cloud scenes is $4$. For each scene, FRNet generates confidence scores for each class. We add the confidence scores of the $4$ scenes and then predict the final semantic labels.
Appendix B Additional Experimental Results
In this section, we provide the complete experimental results and qualitative visual examples among the four popular LiDAR segmentation benchmarks used in this work.
(a) Speed vs. Accuracy
(b) Speed vs. Robustness
Figure 7: The scalability analysis of existing LiDAR semantic segmentation approaches. Subfigure (a) The inference speed vs. segmentation accuracy (the higher the better). Subfigure (b) The inference speed vs. corruption error (the lower the better).
B.1 Scalability and Robustness
In this section, we provide additional analysis of existing LiDAR segmentation approaches regarding their trade-offs in-between the inference speed (FPS), segmentation accuracy (mIoU), and out-of-training-distribution robustness (mCE). As shown in Fig. 7, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed. Those more accurate LiDAR segmentation models often contain larger parameter sets and are less efficient during the inference stage. This is particularly predominant for the voxel-based (as shown in yellow circulars) and point-voxel fusion (as shown in blue circulars) approaches. The range-view approaches, on the contrary, are faster in terms of inference speed. However, existing range-view models only achieve sub-par performance compared to the voxel-based and point-voxel fusion approaches. It is worth highlighting again that the proposed FRNet and Fast-FRNet provide a better trade-off in-between speed, accuracy, and robustness. Our models achieve competitive segmentation results and superior robustness over the voxel-based and point-voxel fusion counterparts, while still maintaining high efficiency for real-time LiDAR segmentation.
B.2 Per-Class Results
In this section, we supplement the complete (class-wise) segmentation results of the baselines, prior works, and our proposed FRNet and Fast-FRNet.
B.2.1 SemanticKITTI
Tab. 5 shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [5]. We compare FRNet with different LiDAR representations, including point-view, range-view, voxel-view, and multi-view methods. Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods. Compared with SphereFormer [27] and UniSeg [32], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy. More results about efficiency and accuracy can be found in the main body. It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
B.2.2 nuScenes
Tab. 6 and Tab. 7 show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [13], respectively. The results demonstrate the great advantages of Fast-FRNet and FRNet. For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters. Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
B.2.3 ScribbleKITTI
To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [49], including RangeNet++ [35], SalsaNext [12], FIRNet [64], CENet [8], and RangeViT [4]. All the models are trained with their official settings. As shown in Tab. 8, FRNet achieves state-of-the-art performance. However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
B.2.4 SemanticPOSS
We also re-implement some popular voxel-view methods, including SPVCNN [47], Cylinder3D [65], and MinkNet [10], on SemanticPOSS [37]. All the experiments are conducted based on MMDetection3D framework [11]. Tab. 9 summarizes the class-wise IoU scores among these methods and some other popular range-view methods. Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods. For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
B.3 Qualitative Results
In this section, we provide additional qualitative examples to further validate the effectiveness and superiority of the proposed FRNet framework.
B.3.1 Visual Comparisons
We supplement visual comparison examples of the proposed FRNet with prior works [4] and show the results in Fig. 9. Compared with range-view methods, FRNet can handle better for objects with complex structures, such as vegetation and terrain. In addition, the objects have much clearer boundaries.
B.3.2 Failure Cases
Although FRNet achieves great improvements among popular range-view methods, there are still some limitations that need to be considered. As shown in Fig. 8, FRNet tends to misclassify objects with similar structures to a large extent, i.e., bicyclist and motorcyclist, sidewalk, and other-ground. Spatial-temporal methods will help to alleviate this problem and yield better performance.
B.3.3 Video Demos
In addition to the figures shown in the main body and supplementary material, we also provide some video demos to support the superiority of FRNet. Specifically, we attached three video clips in the supplementary file, named demo1.mp4, demo2.mp4, and demo3.mp4. The first two demos include the error maps of the proposed FRNet and the prior art RangeViT [4]. The third demo contains the error maps and semantic predictions of FRNet. Each video demo shows LiDAR semantic segmentation sequences with hundreds of frames. All three demos validate again the effectiveness and superiority of FRNet in tackling the challenging LiDAR semantic segmentation task. We hope these video demos can provide a more comprehensive qualitative assessment of our proposed approach.
Table 5: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [5] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign PointNet [39] $14.6$ $46.3$ $1.3$ $0.3$ $0.1$ $0.8$ $0.2$ $0.2$ $0.0$ $61.6$ $15.8$ $35.7$ $1.4$ $41.4$ $12.9$ $31.0$ $4.6$ $17.6$ $2.4$ $3.7$ PointNet++ [40] $20.1$ $53.7$ $1.9$ $0.2$ $0.9$ $0.2$ $0.9$ $1.0$ $0.0$ $72.0$ $18.7$ $41.8$ $5.6$ $62.3$ $16.9$ $46.5$ $13.8$ $30.0$ $6.0$ $8.9$ SqSeg [50] $30.8$ $68.3$ $18.1$ $5.1$ $4.1$ $4.8$ $16.5$ $17.3$ $1.2$ $84.9$ $28.4$ $54.7$ $4.6$ $61.5$ $29.2$ $59.6$ $25.5$ $54.7$ $11.2$ $36.3$ SqSegV2 [51] $39.6$ $82.7$ $21.0$ $22.6$ $14.5$ $15.9$ $20.2$ $24.3$ $2.9$ $88.5$ $42.4$ $65.5$ $18.7$ $73.8$ $41.0$ $68.5$ $36.9$ $58.9$ $12.9$ $41.0$ RandLA-Net [21] $50.3$ $94.0$ $19.8$ $21.4$ $42.7$ $38.7$ $47.5$ $48.8$ $4.6$ $90.4$ $56.9$ $67.9$ $15.5$ $81.1$ $49.7$ $78.3$ $60.3$ $59.0$ $44.2$ $38.1$ RangeNet++ [35] $52.2$ $91.4$ $25.7$ $34.4$ $25.7$ $23.0$ $38.3$ $38.8$ $4.8$ $91.8$ $65.0$ $75.2$ $27.8$ $87.4$ $58.6$ $80.5$ $55.1$ $64.6$ $47.9$ $55.9$ PolarNet [61] $54.3$ $93.8$ $40.3$ $30.1$ $22.9$ $28.5$ $43.2$ $40.2$ $5.6$ $90.8$ $61.7$ $74.4$ $21.7$ $90.0$ $61.3$ $84.0$ $65.5$ $67.8$ $51.8$ $57.5$ MPF [2] $55.5$ $93.4$ $30.2$ $38.3$ $26.1$ $28.5$ $48.1$ $46.1$ $18.1$ $90.6$ $62.3$ $74.5$ $30.6$ $88.5$ $59.7$ $83.5$ $59.7$ $69.2$ $49.7$ $58.1$ 3D-MiniNet [3] $55.8$ $90.5$ $42.3$ $42.1$ $28.5$ $29.4$ $47.8$ $44.1$ $14.5$ $91.6$ $64.2$ $74.5$ $25.4$ $89.4$ $60.8$ $82.8$ $60.8$ $66.7$ $48.0$ $56.6$ SqSegV3 [53] $55.9$ $92.5$ $38.7$ $36.5$ $29.6$ $33.0$ $45.6$ $46.2$ $20.1$ $91.7$ $63.4$ $74.8$ $26.4$ $89.0$ $59.4$ $82.0$ $58.7$ $65.4$ $49.6$ $58.9$ KPConv [48] $58.8$ $96.0$ $32.0$ $42.5$ $33.4$ $44.3$ $61.5$ $61.6$ $11.8$ $88.8$ $61.3$ $72.7$ $31.6$ $95.0$ $64.2$ $84.8$ $69.2$ $69.1$ $56.4$ $47.4$ SalsaNext [12] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $19.4$ $91.7$ $63.7$ $75.8$ $29.1$ $90.2$ $64.2$ $81.8$ $63.6$ $66.5$ $54.3$ $62.1$ FIDNet [64] $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $23.7$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$ FusionNet [59] $61.3$ $95.3$ $47.5$ $37.7$ $41.8$ $34.5$ $59.5$ $56.8$ $11.9$ $91.8$ $68.8$ $77.1$ $30.8$ $92.5$ $69.4$ $84.5$ $69.8$ $68.5$ $60.4$ $66.5$ PCSCNet [38] $62.7$ $95.7$ $48.8$ $46.2$ $36.4$ $40.6$ $55.5$ $68.4$ $55.9$ $89.1$ $60.2$ $72.4$ $23.7$ $89.3$ $64.3$ $84.2$ $68.2$ $68.1$ $60.5$ $63.9$ KPRNet [23] $63.1$ $95.5$ $54.1$ $47.9$ $23.6$ $42.6$ $65.9$ $65.0$ $16.5$ $93.2$ $73.9$ $80.6$ $30.2$ $91.7$ $68.4$ $85.7$ $69.8$ $71.2$ $58.7$ $64.1$ TornadoNet [16] $63.1$ $94.2$ $55.7$ $48.1$ $40.0$ $38.2$ $63.6$ $60.1$ $34.9$ $89.7$ $66.3$ $74.5$ $28.7$ $91.3$ $65.6$ $85.6$ $67.0$ $71.5$ $58.0$ $65.9$ LiteHDSeg [43] $63.8$ $92.3$ $40.0$ $55.4$ $37.7$ $39.6$ $59.2$ $71.6$ $54.3$ $93.0$ $68.2$ $78.3$ $29.3$ $91.5$ $65.0$ $78.2$ $65.8$ $65.1$ $59.5$ $67.7$ RangeViT [4] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $38.1$ $93.1$ $70.2$ $80.0$ $32.5$ $92.0$ $69.0$ $85.3$ $70.6$ $71.2$ $60.8$ $64.7$ CENet [8] $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $43.5$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $70.0$ $61.5$ $67.6$ SVASeg [63] $65.2$ $96.7$ $56.4$ $57.0$ $49.1$ $56.3$ $70.6$ $67.0$ $15.4$ $92.3$ $65.9$ $76.5$ $23.6$ $91.4$ $66.1$ $85.2$ $72.9$ $67.8$ $63.9$ $65.2$ AMVNet [31] $65.3$ $96.2$ $59.9$ $54.2$ $48.8$ $45.7$ $71.0$ $65.7$ $11.0$ $90.1$ $71.0$ $75.8$ $32.4$ $92.4$ $69.1$ $85.6$ $71.7$ $69.6$ $62.7$ $67.2$ GFNet [41] $65.4$ $96.0$ $53.2$ $48.3$ $31.7$ $47.3$ $62.8$ $57.3$ $44.7$ $\mathbf{93.6}$ $72.5$ $\mathbf{80.8}$ $31.2$ $\mathbf{94.0}$ $\mathbf{73.9}$ $85.2$ $71.1$ $69.3$ $61.8$ $68.0$ JS3C-Net [55] $66.0$ $95.8$ $59.3$ $52.9$ $54.3$ $46.0$ $69.5$ $65.4$ $39.9$ $88.9$ $61.9$ $72.1$ $31.9$ $92.5$ $70.8$ $84.5$ $69.8$ $67.9$ $60.7$ $68.7$ MaskRange [17] $66.1$ $94.2$ $56.0$ $55.7$ $59.2$ $52.4$ $67.6$ $64.8$ $31.8$ $91.7$ $70.7$ $77.1$ $29.5$ $90.6$ $65.2$ $84.6$ $68.5$ $69.2$ $60.2$ $66.6$ SPVNAS [47] $66.4$ $97.3$ $51.5$ $50.8$ $59.8$ $58.8$ $65.7$ $65.2$ $43.7$ $90.2$ $67.6$ $75.2$ $16.9$ $91.3$ $65.9$ $86.1$ $73.4$ $71.0$ $64.2$ $66.9$ MSSNet [46] $66.7$ $96.8$ $52.2$ $48.5$ $54.4$ $56.3$ $67.0$ $70.9$ $49.3$ $90.1$ $65.5$ $74.9$ $30.2$ $90.5$ $64.9$ $84.9$ $72.7$ $69.2$ $63.2$ $65.1$ Cylinder3D [65] $68.9$ $97.1$ $67.6$ $63.8$ $50.8$ $58.5$ $73.7$ $69.2$ $48.0$ $92.2$ $65.0$ $77.0$ $32.3$ $90.7$ $66.5$ $85.6$ $72.5$ $69.8$ $62.4$ $66.2$ AF2S3Net [9] $69.7$ $94.5$ $65.4$ $\mathbf{86.8}$ $39.2$ $41.1$ $\mathbf{80.7}$ $80.4$ $74.3$ $91.3$ $68.8$ $72.5$ $\mathbf{53.5}$ $87.9$ $63.2$ $70.2$ $68.5$ $53.7$ $61.5$ $71.0$ RPVNet [54] $70.3$ $97.6$ $68.4$ $68.7$ $44.2$ $61.1$ $75.9$ $74.4$ $73.4$ $93.4$ $70.3$ $80.7$ $33.3$ $93.5$ $72.1$ $86.5$ $75.1$ $71.7$ $64.8$ $61.4$ SDSeg3D [28] $70.4$ $97.4$ $58.7$ $54.2$ $54.9$ $65.2$ $70.2$ $74.4$ $52.2$ $90.9$ $69.4$ $76.7$ $41.9$ $93.2$ $71.1$ $86.1$ $74.3$ $71.1$ $65.4$ $70.6$ GASN [58] $70.7$ $96.9$ $65.8$ $58.0$ $59.3$ $61.0$ $80.4$ $\mathbf{82.7}$ $46.3$ $89.8$ $66.2$ $74.6$ $30.1$ $92.3$ $69.6$ $87.3$ $73.0$ $72.5$ $66.1$ $71.6$ PVKD [20] $71.2$ $97.0$ $67.9$ $69.3$ $53.5$ $60.2$ $75.1$ $73.5$ $50.5$ $91.8$ $70.9$ $77.5$ $41.0$ $92.4$ $69.4$ $86.5$ $73.8$ $71.9$ $64.9$ $65.8$ Fast-FRNet $72.5$ $97.1$ $66.0$ $73.1$ $59.3$ $65.7$ $76.0$ $78.9$ $54.5$ $91.8$ $72.6$ $77.6$ $42.0$ $92.4$ $70.6$ $86.5$ $71.9$ $72.5$ $63.1$ $66.7$ 2DPASS [56] $72.9$ $97.0$ $63.6$ $63.4$ $61.1$ $61.5$ $77.9$ $81.3$ $74.1$ $89.7$ $67.4$ $74.7$ $40.0$ $93.5$ $72.9$ $86.2$ $73.9$ $71.0$ $65.0$ $70.4$ RangeFormer [24] $73.3$ $96.7$ $69.4$ $73.7$ $59.9$ $66.2$ $78.1$ $75.9$ $58.1$ $92.4$ $73.0$ $78.8$ $42.4$ $92.3$ $70.1$ $86.6$ $73.3$ $72.8$ $66.4$ $66.6$ FRNet $73.3$ $97.3$ $67.9$ $74.6$ $59.4$ $66.3$ $78.1$ $79.2$ $57.3$ $92.1$ $73.0$ $78.1$ $41.8$ $92.7$ $71.0$ $86.7$ $73.2$ $72.5$ $64.7$ $67.3$ SphereFormer [27] $74.8$ $97.5$ $70.1$ $70.5$ $59.6$ $67.7$ $79.0$ $80.4$ $\mathbf{75.3}$ $91.8$ $69.7$ $78.2$ $41.3$ $93.8$ $72.8$ $86.7$ $75.1$ $72.4$ $66.8$ $\mathbf{72.9}$ UniSeg [32] $\mathbf{75.2}$ $\mathbf{97.9}$ $\mathbf{71.9}$ $75.2$ $\mathbf{63.6}$ $\mathbf{74.1}$ $78.9$ $74.8$ $60.6$ $92.6$ $\mathbf{74.0}$ $79.5$ $46.1$ $93.4$ $72.7$ $\mathbf{87.5}$ $\mathbf{76.3}$ $\mathbf{73.1}$ $\mathbf{68.3}$ $68.5$
Table 6: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [9] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [35] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [61] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [38] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [12] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [63] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [4] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [65] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [31] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [54] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [24] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [27] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
Table 7: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [13]. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [61] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [55] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [66] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [65] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [31] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [47] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [9] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [15] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [24] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [58] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [56] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [57] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [27] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [32] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
Table 8: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [49] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign RangeNet++ [35] $44.6$ $84.6$ $24.6$ $38.9$ $13.7$ $12.9$ $29.0$ $51.4$ $0.0$ $86.0$ $35.2$ $72.4$ $4.5$ $80.2$ $41.6$ $80.1$ $51.2$ $66.9$ $43.9$ $34.5$ SalsaNext [12] $50.3$ $85.6$ $33.8$ $42.7$ $8.4$ $20.8$ $64.3$ $71.5$ $0.2$ $85.6$ $31.6$ $72.0$ $1.1$ $84.0$ $39.9$ $81.0$ $62.0$ $63.6$ $60.5$ $46.3$ RangeViT [4] $53.6$ $85.6$ $31.6$ $50.1$ $40.3$ $36.3$ $57.6$ $68.7$ $0.0$ $86.1$ $32.6$ $75.2$ $0.3$ $87.9$ $49.3$ $83.6$ $62.8$ $67.5$ $59.6$ $43.7$ FIDNet [64] $54.1$ $85.6$ $36.7$ $48.7$ $60.8$ $38.4$ $63.3$ $68.2$ $0.0$ $84.1$ $25.9$ $71.2$ $0.4$ $85.6$ $41.3$ $81.7$ $64.1$ $62.7$ $61.5$ $48.0$ MinkNet [10] $55.0$ $88.1$ $13.2$ $55.1$ $72.3$ $36.9$ $61.3$ $77.1$ $0.0$ $83.4$ $32.7$ $71.0$ $0.3$ $90.0$ $50.0$ $84.1$ $66.6$ $65.8$ $61.6$ $35.2$ CENet [8] $55.7$ $86.1$ $39.4$ $53.2$ $61.0$ $46.1$ $69.2$ $72.2$ $0.0$ $85.7$ $28.7$ $72.6$ $1.1$ $85.8$ $43.1$ $81.8$ $64.2$ $63.8$ $59.6$ $45.0$ SPVCNN [47] $56.9$ $88.6$ $25.7$ $55.9$ $67.4$ $48.8$ $65.0$ $78.2$ $0.0$ $82.6$ $30.4$ $70.1$ $0.3$ $\mathbf{90.5}$ $49.6$ $84.4$ $\mathbf{67.6}$ $66.1$ $61.6$ $\mathbf{48.7}$ Cylinder3D [65] $57.0$ $88.5$ $39.9$ $58.0$ $58.4$ $48.1$ $68.6$ $77.0$ $\mathbf{0.5}$ $84.4$ $30.4$ $72.2$ $2.5$ $89.4$ $48.4$ $81.9$ $64.6$ $59.8$ $61.2$ $\mathbf{48.7}$ Fast-FRNet $62.4$ $90.8$ $41.0$ $\mathbf{66.8}$ $81.7$ $\mathbf{64.0}$ $70.5$ $84.2$ $0.0$ $91.3$ $35.1$ $78.3$ $0.0$ $89.4$ $64.4$ $85.0$ $65.5$ $69.5$ $60.5$ $47.7$ RangeFormer [24] $63.0$ $\mathbf{92.6}$ $\mathbf{51.6}$ $65.7$ $74.4$ $49.6$ $71.6$ $82.1$ $0.0$ $\mathbf{94.8}$ $\mathbf{44.4}$ $\mathbf{80.6}$ $\mathbf{11.4}$ $85.6$ $56.9$ $\mathbf{87.2}$ $64.1$ $\mathbf{77.0}$ $\mathbf{62.7}$ $45.1$ FRNet $\mathbf{63.1}$ $90.5$ $42.3$ $\mathbf{66.8}$ $\mathbf{82.4}$ $63.0$ $\mathbf{73.4}$ $\mathbf{86.2}$ $0.0$ $92.1$ $35.1$ $79.1$ $0.3$ $90.4$ $\mathbf{64.7}$ $85.4$ $65.8$ $70.7$ $61.8$ $48.6$
Appendix C Broader Impact
In this section, we elaborate on the positive societal influence and potential limitations of the proposed FRNet.
C.1 Positive Societal Impacts
Conducting efficient and accurate range view LiDAR semantic segmentation for autonomous driving has several positive societal impacts, particularly in terms of enhancing road safety, improving transportation efficiency, and contributing to broader technological advancements.
  • Enhanced Road Safety. One of the most significant benefits is the improvement in road safety. Autonomous vehicles equipped with advanced LiDAR semantic segmentation can accurately detect and classify objects in their environment, such as other vehicles, pedestrians, cyclists, and road obstacles. This precise perception capability allows for safer navigation and decision-making, potentially reducing accidents caused by human error.
  • Reduced Traffic Congestion. Autonomous vehicles with advanced perception systems can optimize driving patterns, leading to smoother traffic flow. This can significantly reduce traffic congestion, especially in urban areas, thereby saving time for commuters and reducing stress associated with driving.
  • Emergency Response and Healthcare. In emergency situations, autonomous vehicles can be used to quickly and safely transport patients or deliver medical supplies. The precision of LiDAR segmentation ensures that these vehicles can navigate through complex environments effectively.
  • Advancements in Smart City Infrastructure. The integration of autonomous vehicles with smart city initiatives can lead to more efficient urban planning and infrastructure development. LiDAR data can be used not just for navigation but also for gathering urban data, which can inform city planning and maintenance.
C.2 Potential Limitations
Although the proposed FRNet and Fast-FRNet are capable of providing a better trade-off between LiDAR segmentation network efficiency and accuracy, there are several limitations within the current framework. First, frustum-level supervision counts the high-frequency semantic labels in the frustum region, which will cover the objects with few points. Such supervision weakens the performance of FRNet on tiny objects. Second, FRNet cannot handle objects with similar structures well. Some objects share a similar appearance with different semantic attributes, which limits the ability of FRNet to distinguish them in a 2D manner.
Appendix D Public Resources Used
In this section, we acknowledge the use of public resources, during the course of this work.
D.1 Public Codebase Used
We acknowledge the use of the following public codebase, during the course of this work:
D.2 Public Datasets Used
We acknowledge the use of the following public datasets, during the course of this work:
D.3 Public Implementations Used
We acknowledge the use of the following public implementations, during the course of this work:
Table 9: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [37] leaderboard. All IoU scores are given in percentage (%). The best and second best scores for each class are highlighted in bold and underline.
Method (year)
mIoU
person
rider
car
truck
plants
traffic sign
pole
trashcan
building
cone/stone
walk
fence
bike
SqSeg [50] [’18] $18.9$ $14.2$ $1.0$ $13.2$ $10.4$ $28.0$ $5.1$ $5.7$ $2.3$ $43.6$ $0.2$ $15.6$ $31.0$ $75.0$
SqSegV2 [51] [’19] $30.0$ $48.0$ $9.4$ $48.5$ $11.3$ $50.1$ $6.7$ $6.2$ $14.8$ $60.4$ $5.2$ $22.1$ $36.1$ $71.3$
RangeNet++ [35] [’19] $30.9$ $57.3$ $4.6$ $35.0$ $14.1$ $58.3$ $3.9$ $6.9$ $24.1$ $66.1$ $6.6$ $23.4$ $28.6$ $73.5$
MINet [30] [’21] $43.2$ $62.4$ $12.1$ $63.8$ $22.3$ $68.6$ $16.7$ $30.1$ $28.9$ $75.1$ $28.6$ $32.2$ $44.9$ $76.3$
FIDNet [64] [’21] $46.4$ $72.2$ $23.1$ $72.7$ $23.0$ $68.0$ $22.2$ $28.6$ $16.3$ $73.1$ $34.0$ $40.9$ $50.3$ $79.1$
SPVCNN [47] [’20] $48.4$ $72.5$ $24.7$ $72.1$ $31.4$ $72.7$ $10.8$ $41.3$ $31.8$ $78.4$ $23.8$ $42.6$ $51.7$ $75.3$
CENet [8] [’22] $50.3$ $75.5$ $22.0$ $77.6$ $25.3$ $72.2$ $18.2$ $31.5$ $\mathbf{48.1}$ $76.3$ $27.7$ $47.7$ $51.4$ $\mathbf{80.3}$
Fast-FRNet $52.5$ $76.9$ $28.3$ $79.9$ $28.8$ $73.8$ $30.2$ $32.9$ $28.5$ $80.9$ $40.7$ $47.3$ $55.9$ $79.0$
Cylinder3D [65] [’21] $52.9$ $75.9$ $\mathbf{30.0}$ $75.8$ $28.7$ $\mathbf{75.7}$ $29.5$ $37.2$ $36.7$ $\mathbf{82.3}$ $34.1$ $47.5$ $53.9$ $80.1$
MinkNet [10] [’19] $53.1$ $\mathbf{77.8}$ $29.1$ $\mathbf{81.3}$ $\mathbf{33.9}$ $75.2$ $22.0$ $\mathbf{42.5}$ $36.4$ $80.7$ $23.9$ $\mathbf{51.2}$ $\mathbf{57.5}$ $79.1$
FRNet $\mathbf{53.5}$ $\underline{77.7}$ $28.7$ $81.2$ $28.9$ $74.3$ $\mathbf{30.4}$ $34.5$ $29.8$ $81.1$ $\mathbf{45.6}$ $47.9$ $56.4$ $79.2$
Figure 8: Failure case examples in FRNet. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
Figure 9: Qualitative comparisons with state-of-the-art LiDAR segmentation methods [4] on the val set of nuScenes [13]. To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively. Each point cloud scene is visualized under both the bird’s-eye-view and range-view.

### v2

@@
 Abstract
-LiDAR segmentation is crucial for autonomous driving systems.
-The recent range-view approaches are promising for real-time processing.
-However, they suffer inevitably from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement.
-In this work, we propose a simple yet powerful FRNet that restores the contextual information of the range image pixels with corresponding frustum LiDAR points.
+LiDAR segmentation has become a crucial component in advanced autonomous driving systems.
+Recent range-view LiDAR segmentation approaches show promise for real-time processing.
+However, they inevitably suffer from corrupted contextual information and rely heavily on post-processing techniques for prediction refinement.
+In this work, we propose FRNet, a simple yet powerful method aimed at restoring the contextual information of range image pixels using corresponding frustum LiDAR points.
 Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions.
-Next, a frustum-point fusion module is introduced to update per-point features hierarchically, which enables each point to extract more surrounding information via the frustum features.
+Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information via the frustum features.
 Finally, a head fusion module is used to fuse features at different levels for final semantic prediction.
-Extensive experiments on four popular LiDAR segmentation benchmarks under various task setups demonstrate our superiority.
-FRNet achieves competitive performance while maintaining high efficiency.
-The code is publicly available.
-1 Introduction
-Figure 1: Scalability comparisons among state-of-the-art LiDAR segmentation models on the SemanticKITTI [5] leaderboard.
-The size of the circular represents the number of model parameters.
-Our FRNet achieves competitive performance with current arts while maintaining satisfactory efficiency for real-time processing.
-Real-time and accurate LiDAR semantic segmentation is crucial in autonomous driving, robotics, and other safety-critical applications [18].
-Many recent approaches have made significant progress in this area, which represent LiDAR point clouds from different perspectives, including point view [48, 21], voxel view [47, 65], pseudo-image view [50, 35, 61], and multi-view representation [31, 54, 2], depending on different pre-processing techniques.
-Point-view methods manipulate the origin point cloud to preserve raw spatial information, following the paradigm of PointNet++ [40].
-Nevertheless, they require computationally expensive neighborhood searches to construct local neighborhoods, and the computation overhead is considerable for large scene point clouds.
-Voxel-view methods transform scatter points into regular voxel representations and apply 3D convolutions for label prediction, but the computation grows cubically when voxel resolution increases.
-Multi-view methods extract features from different representations separately and fuse these features from different levels for more accurate prediction.
-Consequently, the computation of all representations is also gathered.
-Albeit these representation methods achieve some satisfactory performance, their computation overhead hinders them from real-time online applications.
-Recently, pseudo-image methods propose a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [61, 1] and range-view [50, 51, 43, 4, 24].
-Under such representations, some popular image segmentation approaches [33, 62] can be directly applied to such pseudo-images to balance the efficiency and accuracy of segmentation.
-However, the 3D-to-2D projection procedure tends to result in corrupted neighborhood contextual information, decreasing the precision.
-Detailed analysis of this problem can be summarized into two aspects.
-Firstly, some 3D points are discarded in the projection procedure.
-Under such a paradigm, only the projected points are involved in convolutions, while the contextual information of discarded points is absent.
-Consequently, the pixel in the pseudo-image can hardly preserve the high-fidelity 3D geometric context.
-More importantly, these discarded 3D points still need to predict 3D labels but no features correspond to them.
-To reconstruct these labels, some post-processing techniques, such as CRF [22] and KNN [35], are used for these points based on 2D predictions.
-However, these techniques are unsupervised and fail to extract the local features of discarded points in a learnable manner, which significantly limits the improvement of pseudo-image methods.
-Secondly, the pseudo-image suffers from many empty pixels due to the sparsity of LiDAR points.
-Under such a non-compact representation, the projected points will extract much noisy information as the convolutions go deeper.
-In summary, pseudo-image methods fail to preserve the contextual information of all the points, which is significant for LiDAR segmentation.
-So, it is important for points and pseudo-images to cooperate with each other.
-Although existing multi-view methods have successfully preserved the context of all points via fusing features of different representations, they are always computationally expensive.
-To this end, we study a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which can preserve the contextual information of all points through the cooperation of point and pseudo-image and still maintain high efficiency.
-Specifically, FRNet comprises three main components.
-To preserve all points and predict semantic labels for these points in an end-to-end manner, a Frustum Feature Encoder is used to group all points with the same frustum region into the corresponding range-view pixel in relation to the range image via several multi-layer perceptions.
-The point features can be pooled to represent the frustum region and formatted into a 2D representation, followed by traditional convolutions.
-Then, a Frustum-Point (FP) Fusion Module is introduced to update the features of each point hierarchically during each stage of the convolutions, which consists of frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features.
-In this way, all the points extract larger local features based on the frustum region.
-Finally, a Fusion Head module is designed to utilize different level features to generate individual features for each point, so that the results can be predicted end-to-end without any post-processing techniques.
+Extensive experiments conducted on four popular LiDAR segmentation benchmarks under various task setups demonstrate the superiority of FRNet.
+Notably, FRNet achieves 73.3% and 82.5% mIoU scores on the testing sets of SemanticKITTI and nuScenes.
+While achieving competitive performance, FRNet operates 5 times faster than state-of-the-art approaches.
+Such high efficiency opens up new possibilities for more scalable LiDAR segmentation.
+The code has been made publicly available at https://github.com/Xiangxu-0103/FRNet.
+Index Terms:
+LiDAR Semantic Segmentation; Autonomous Driving; Real-Time Processing; Frustum-Range Representation
+I Introduction
+LiDAR segmentation, an crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4].
+The principal challenge lies in devising a scalable LiDAR segmentation system that strikes a delicate balance between efficiency and accuracy, especially in resource-constrained operational scenarios [5, 6].
+Existing LiDAR segmentation methods adopt various data representation perspectives, each with its own set of trade-offs in terms of accuracy and computational efficiency.
+The point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve raw spatial information, but entail computationally expensive neighborhood searches to construct local structures.
+This demand results in considerable computation overhead and limits their capability to deal with large-scale scene point clouds [13, 14, 10].
+Sparse-voxel-based methods [15, 16, 17, 18] transform scattered LiDAR points into regular voxel representations and leverage popular sparse convolutions to extract features.
+However, this process also demands heavy computations, especially for applications that require high voxel resolutions [16, 15, 19].
+Multi-view methods [20, 21, 22, 23] extract features from various representations to enhance prediction accuracy [24, 25, 26].
+While this strategy can improve performance, it aggregates the computation required for all representations.
+Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [27, 23].
+Figure 1: A study on the scalability of state-of-the-art LiDAR segmentation models on the SemanticKITTI [28] leaderboard.
+The size of the circular representation corresponds to the number of model parameters.
+FRNet achieves competitive performance with current state-of-the-art models while still maintaining satisfactory efficiency for real-time processing.
+Recently, pseudo-image-based methods [29, 30, 31, 32, 33, 34, 35, 36] have appeared as a simple yet efficient intermediate representation for LiDAR processing, such as bird’s-eye-view [33, 34, 35] and range-view [29, 30, 31, 32], which are more computationally tractable.
+These representations enable the direct application of popular 2D segmentation approaches [37, 38, 39] to the pseudo-images, offering a promising solution for real-time LiDAR segmentation.
+However, the 3D-to-2D projection inevitably introduces corrupted neighborhood contextual information, which poses limitations on the further development of pseudo-image approaches [40].
+The detailed analysis of this problem can be further summarized into two aspects as follows.
+Firstly, a fundamental concern with pseudo-image methods is the inadvertent exclusion of certain 3D points during the conversion to 2D projections [41].
+This leads to an incomplete representation where only the points that are projected are involved in subsequent convolutional operations.
+Such a scenario disregards the contextual information of the unprojected points, causing the pseudo-images to lose crucial high-fidelity 3D geometric context [42].
+Furthermore, the omitted points, which still necessitate 3D label predictions, lack associated feature information, thereby necessitating post-processing to infer these labels.
+Unfortunately, current post-processing methods struggle to incorporate local features of these omitted points effectively, capping the potential advancements in pseudo-image approaches.
+Secondly, the challenge with pseudo-images is exacerbated by the significant presence of empty pixels due to the inherent sparsity of LiDAR data.
+For instance, in the nuScenes dataset, nearly 60% of range image pixels are found to be void [43, 32].
+This vacancy not only skews the scene representation but also compromises the quality of semantic segmentation.
+As the network processes deeper layers, these sparse pseudo-images can become riddled with noise, obstructing the extraction of meaningful patterns.
+In essence, although pseudo-image methods excel in terms of real-time processing capabilities, their reduced dimensional perspective can overlook crucial 3D details, leading to degraded scalability.
+Observing the above issues, in this work, we propose a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy.
+FRNet consists of three main components.
+Firstly, a frustum feature encoder (FFE) is utilized to group all points with the same frustum region into corresponding range-view pixels in relation to the range image using multiple multi-layer perceptrons (MLPs).
+This allows for the preservation of all points and the prediction of semantic labels for these points in an end-to-end manner.
+Subsequently, the point features are pooled to represent the frustum region and formatted into a 2D representation, which is then subjected to traditional convolutions.
+Secondly, a frustum-point (FP) fusion module is employed to efficiently update the hierarchical features of each point during each convolutional stage.
+This module includes frustum-to-point fusion to update per-point features and point-to-frustum fusion to enhance frustum features.
+As a result, all points extract larger local features based on the frustum region.
+Finally, a fusion head (FH) module is designed to leverage features from different levels to generate individual features for each point, facilitating end-to-end prediction without the need for post-processing techniques.
 As shown in Fig.
-1, the proposed FRNet achieves great improvements among range-view methods while still maintaining high efficiency.
-To alleviate empty pixels in 2D representation, we introduce a new range interpolation augmentation method that adds several potential points based on the information of surrounding points.
-Specifically, we first transform the point clouds into range images, as done in previous works.
-For empty pixels in the range images, we then calculate their potential coordinates when both the nearest left and right pixels have projected points as they often share similar attributes.
-With this augmentation, the 2D representation will have fewer empty pixels, allowing convolutions to learn more semantically coherent information.
-The key contributions are summarized as follows:
-  • We introduce a novel Frustum-Range representation for LiDAR segmentation.
-In this representation, each pixel in the 2D representation is treated as a corresponding frustum region in the point clouds.
-All points within the same frustum region are preserved for semantic prediction without the need for any post-processing techniques.
-  • We propose a novel range interpolation augmentation that synthesizes possibly lost LiDAR points based on surrounding information, helping to fill empty pixels and learn more coherent range-view representations.
-  • Extensive experiments across four LiDAR segmentation benchmarks demonstrate that FRNet can achieve the optimal trade-off between efficiency and accuracy.
-2 Related Work
-Scalable LiDAR Segmentation.
-Although Point-view [21, 48], voxel-view [47, 65], and multi-view [31, 32] have achieved great success in LiDAR segmentation, they are often time-consuming and cannot achieve real-time processing.
-Recently, various methods proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation.
-SqueezeSeg [50] and SqueezeSegV2 [51] used the lightweight model SqueezeNet [22] to retain information.
-SqueezeSegV3 [53] introduced a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images.
-RangeNet++ [35] integrated Darknet into SqueezeSeg [50] and proposed an efficient KNN post-processing for segmentation.
-RangeViT [4] and RangeFormer [24] introduced popular transformer blocks to extract local and global information of the range images.
-These methods only learn the features of projected points while ignoring the occluded points.
-LiDAR Post-Processing.
+1, the proposed FRNet achieves great improvement among range-view methods while still maintaining high efficiency.
+Data augmentation plays an important role in training robust and generalizable models.
+Inspired by the success of mixing strategies [44, 45, 46], we introduce a novel mixing method dubbed FrustumMix, which operates based on frustum units.
+Specifically, we randomly divide point cloud scenes into several frustum regions along inclination and azimuth directions and swap the corresponding regions from different scenes, with the neighborhood of each frustum region coming from the other scene.
+FrustumMix provides a straightforward and lightweight method to adapt FRNet by focusing on the frustum region.
+Furthermore, to address the issue of empty pixels in the 2D representation, we propose RangeInterpolation to reconstruct the semantic surface based on surrounding point information.
+It first transforms the LiDAR point cloud into a range image, as done in previous works [29, 30, 31, 32].
+For empty pixels in the range image, we then aggregate surrounding range information within a pre-defined window using an average pooling operation to generate a new point.
+RangeInterpolation is able to assist in creating a more compact 2D representation with fewer empty pixels, enabling convolutions to learn more semantically coherent information.
+In summary, the key contributions of this work can be listed as follows:
+  • We introduce a novel Frustum-Range representation for scalable LiDAR segmentation.
+In our framework, the point-level geometric information is integrated into the pseudo-image representation, which not only retains the complete geometric structure of the point cloud, but also takes advantage of the efficiency of the 2D network.
+  • We propose two novel 3D data augmentation techniques to assist in training robust and generalizable models.
+FrustumMix generates more complex scenes by mixing two different scans, while RangeInterpolation reconstructs the semantic surfaces based on surrounding range information.
+These techniques help generate a compact 2D representation and learn more context-aware features.
+  • Extensive experiments across four prevailing LiDAR segmentation benchmarks demonstrate our superiority.
+FRNet achieves $73.3\%$ and $82.5\%$ mIoU scores, respectively, on SemanticKITTI and nuScenes, while still maintaining promising scalability for real-time LiDAR segmentation.
+II Related Work
+II-A Scalable LiDAR Segmentation
+Although point-view [7, 8, 10, 11, 12], sparse-voxel-view [15, 16, 17, 18], and multi-view [20, 21, 22, 23, 47, 48] methods have achieved significant success in LiDAR segmentation, they often suffer from high computational costs and are unable to achieve real-time processing.
+Recently, various methods have been proposed to project LiDAR points into range images along inclination and azimuth directions to balance the efficiency and accuracy of segmentation [49].
+SqueezeSeg [29] and SqueezeSegV2 [50] use the lightweight model SqueezeNet [51] to retain information.
+SqueezeSegV3 [52] introduces a Spatially-Adaptive Convolution module to adjust kernel weights based on the locations of range images.
+RangeNet++ [30] integrates DarkNet into SqueezeSeg [29] and proposes an efficient KNN post-processing for segmentation.
+RangeViT [31] and RangeFormer [32] introduce transformer blocks to extract both local and global information from the range image.
+However, these methods only involve projected points and fail to extract contextual information for discarded points.
+To address this problem, WaffleIron [11] proposes to mix point and image features with residual connections.
+Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predicting over the entire point cloud.
+In this work, we propose to integrate point and image features efficiently, preserving both 3D geometric information from raw point features and efficiency from 2D convolutions, without the need for any pre- or post-processing techniques, which yields superior scalability for real-time LiDAR segmentation.
+II-B LiDAR Data Augmentation
+Recent works have explored various data augmentation techniques on point clouds [53, 54], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes.
+Mix3D [46] creates new training samples by combining two scenes out of context, which requires a larger memory cost.
+PolarMix [45] proposes to mix two LiDAR scenes based on azimuth angles and cropping instances from one scene to another, while LaserMix [44] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning.
+RangeFormer [32] presents a lightweight augmentation method that operates only on projected points in range images, disregarding occluded points.
+In this work, we introduce augmentation techniques aimed at training a robust and scalable model for LiDAR segmentation, including FrustumMix and RangeInterpolation.
+II-C LiDAR Post-Processing
 Prior works have proposed various post-processing techniques to reconstruct the semantic labels of occluded points based on 2D results.
-Many of these methods follow the CRF [50] or KNN [35] approach to smooth the semantic results of occluded points.
-FIDNet [64] proposed a Nearest Label Assignment (NLA) method for label smoothing.
-Both of these techniques are unsupervised and can not extract the features of all points in a learnable manner.
-RangeFormer [24] created sub-clouds from full point clouds and inferred each subset in a supervised manner, but it still relies on 2D prediction results.
-In contrast, our method involves the frustum region and predicts the semantic label of each point directly without any information loss, making it a fully end-to-end approach.
-LiDAR Data Augmentation.
-Recent works have explored different data augmentation techniques on point clouds [7, 60], which can help improve performance when it comes to indoor scene understanding but has limited generalization to outdoor scenes.
-Mix3D [36] created new training samples by combining two scenes out of context.
-PolarMix [52] proposed mixing two LiDAR scenes based on yaw angles and cropping instances from one scene to another, while LaserMix [26] proposed exchanging labeled scans with unlabeled scans along the pitch axis for efficient semi-supervised learning.
-RangeFormer [24] presented a lightweight augmentation method that only operated on projected points in range images, disregarding occluded points.
-In this work, we prevent a novel augmentation technique based on our proposed frustum region, which preserves all points within the frustum region, encompassing techniques such as FrustumMix and Range-Interpolation.
-Figure 2: A pilot study on the performance degradation of post-processing in existing range-view methods [35, 12, 64, 8] on SemanticKITTI [5].
+SqueezeSeg [29] employs the traditional conditional random field (CRF) as a recurrent neural network (RNN).
+RangeNet++ [30] introduces a fast K-Nearest-Neighbor (KNN) search to aggregate final results across the entire point cloud.
+FIDNet [55] uses a Nearest Label Assignment (NLA) method for label smoothing.
+RangeFormer [32] divides full point clouds into sub-clouds and infers each subset in a supervised manner.
+However, these methods heavily rely on 2D results and fail to extract point features over entire points in a learnable manner.
+In contrast, KPRNet [56] and RangeViT [31] incorporate a KPConv [7] block to extract point features in an end-to-end manner.
+Nevertheless, this approach requires substantial computational resources to search neighborhoods in 3D space, which hinders them from real-time processing.
+In this work, we partition points into frustum regions and extract local features of points in a 2D manner, avoiding time-consuming neighbor searches.
+Based on the learned point features, FRNet predicts results over the entire point cloud in an end-to-end manner with high efficiency.
+III FRNet: A Scalable LiDAR Segmentor
+Figure 2: Pilot study on the performance degradation of post-processing in existing range-view methods [30, 57, 55, 58] on the val set of SemanticKITTI [28].
 We choose various $K$ values as hyperparameters in KNN post-processing.
-Compared to their performance at 2D, a severe drop in performance occurs with different $K$ values.
-3 FRNet: A Scalable LiDAR Segmentor
-3.1 Pilot Study
+Compared to their performance at 2D (i.e., the range image), a severe drop in performance occurs with different $K$ values.
+In this section, we first conduct a pilot study on the most popular range-view LiDAR segmentation methods, wherein we observe a significant impact of KNN post-processing on performance (Sec.
+III-A).
+We then elaborate on the technical details of frustum-range architecture (Sec.
+III-B) and frustum-range operators (Sec.
+III-C).
+The overall pipeline of the proposed FRNet framework is depicted in Fig.
+3.
+III-A Pilot Study
 Figure 3: Architecture overview.
-The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features in the frustum region.
-2) Frustum-Point (FP) Fusion Module is used to update per-point features hierarchically at each stage of the 2D backbone.
+The proposed FRNet comprises three main components: 1) Frustum Feature Encoder is used to embed per-point features within the frustum region.
+2) Frustum-Point (FP) Fusion Module updates per-point features hierarchically at each stage of the 2D backbone.
 3) Fusion Head fuses different levels of features to predict final results.
-LiDAR segmentation aims to assign semantic labels to each point in the point clouds.
-A common approach is to perform operations directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [40].
+LiDAR segmentation aims to assign semantic labels to each point in the point cloud [59].
+A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [60].
 However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes.
-To address this issue, many recent works propose projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency.
-Given point clouds with coordinates $(x,y,z)$ and the expected projected resolution $(H,W)$ of range images, each point can be mapped to its correspondence coordinates $(u,v)$ in the range image via the following transformation:
-$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (1)
-where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively.
-However, such a projection process can cause the loss of occluded points and often requires post-processing techniques to reduce the blurry prediction of lost points [24].
-We conduct a pilot study on different range-view methods, including RangeNet++ [35], SalsaNeXt [12], FIDNet [64], and CENet [8], and compare their performance with different KNN post-processing parameters and 2D performance.
-As shown in Fig.
-2, the post-processing procedures are often unsupervised, and the hyperparameters can significantly influence the final performance.
-To address this issue, we propose a scalable Frustum-Range Network (FRNet) that can directly predict semantic labels for each point while maintaining high efficiency.
-3.2 Frustum-Range Architecture
-Given a point cloud $\mathbf{P}$ acquired by the LiDAR sensor, FRNet is expected to learn a feed-forward network $\Phi$ to predict semantic labels $\mathbf{Y}$ for each point:
-$\mathbf{Y}=\Phi(\mathbf{P},\theta)~{},$ (2)
-where $\theta$ denotes the learnable parameters in the network.
-As shown in Fig.
-3, FRNet comprises a Frustum Feature Encoder that carries out per-point feature extraction, a Frustum-Point Fusion Module that fuses per-point features and frustum features hierarchically, and a Fusion Head Module that predicts semantic labels for each point.
+To address this problem, many recent works have proposed projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency.
+Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
+$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})f^{-1}H]}~{},$ (1)
+where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, and $\phi_{\text{up}}$ are $\phi_{\text{down}}$ are the inclination angles at the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image.
+However, such projection often results in the loss of occluded points and requires post-processing techniques to reconstruct the entire semantic information [32].
+We conducted a pilot study on various range-view methods, including RangeNet++ [30], SalsaNext [57], FIDNet [55], and CENet [58].
+We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with various nearest neighbor parameters, as well as their performance on the 2D range images.
+As illustrated in Fig.
+2, the post-processing procedures are often unsupervised, and the choice of hyperparameters can significantly impact the final performance.
+To address this issue, we propose a scalable Frustum-Range Network (FRNet) capable of directly predicting semantic labels for each point while maintaining high efficiency.
+III-B Frustum-Range Architecture
+Problem Definition.
+Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\hat{Y}}$ for each individual point as follows:
+$\mathcal{\hat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
+where $\theta$ represents the learnable parameters within the network.
+As depicted in Fig.
+3, FRNet consists of three key components: 1) a frustum feature encoder responsible for per-point feature extraction; 2) a frustum-point fusion module that efficiently fuses hierarchical point features and frustum features; 3) a fusion head module that integrates point features at different levels for accurate semantic prediction.
+Operator Representation.
+Let $\text{Flat}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{H\times W\times C}$ denotes the function that projects point features onto the frustum image plane with resolution $(H,W)$.
+Within each frustum region, we apply a max-pooling function to obtain the frustum features.
+Let $\text{MLP}(\cdot):\mathbb{R}^{N\times C}\to\mathbb{R}^{N\times D}$ represents the MLPs that take $C$-dimensional point features as input and output $D$-dimensional point features.
+Let $\text{Inflat}(\cdot):\mathbb{R}^{H\times W\times C}\to\mathbb{R}^{N\times C}$ be the operation that back-projects frustum features to point features.
+Notably, points falling into the same frustum region share the same features under this operation.
 Frustum Feature Encoder.
-LiDAR segmentation involves predicting semantic labels for each point, so it is important to extract individual features for each point.
-Before converting point clouds into 2D representations, Frustum Feature Encoder first extracts per-point features via several MLPs, which is crucial for accurate label prediction.
-For points $\mathbf{P}_{i}=\{p_{i}^{1},p_{i}^{2},...,p_{i}^{N}\}$ in the $i$-th frustum region, the points are manually embedded with $(p_{i}^{k},p_{i}^{k}-p_{i}^{mean})$, where $p_{i}^{k}$ is the $k$-th point in the frustum, $N$ is the total number of points in the frustum region, and $p_{i}^{mean}=\frac{1}{N}\sum_{k=1}^{N}p_{i}^{k}$ represents the mean information of the frustum region.
-The per-point features are then obtained as follows:
-$f_{i}^{k}=\text{MLP}([p_{i}^{k};p_{i}^{mean}])~{},$ (3)
+In LiDAR segmentation, the accurate prediction of semantic labels for each point requires the extraction of individual features for each point.
+Preceding the conversion of the point cloud into a 2D representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs.
+This process is crucial for accurate label prediction.
+Firstly, We group the point cloud to yield a set of $M$ frustum regions $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$ according to Eq.
+1, where $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$, consisting of $N_{i}$ points.
+To explicitly embed the frustum structure, we leverage an average pooling function to obtain the cluster points $\widetilde{\mathcal{P}}=\{\widetilde{\mathcal{P}}_{1},\widetilde{\mathcal{P}}_{2},...,\widetilde{\mathcal{P}}_{M}\}$, where $\widetilde{\mathcal{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$.
+Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\widetilde{\mathcal{P}}$ within each frustum region.
+The per-point features are obtained as follows:
+$\mathcal{F}_{p}^{0}=\text{MLP}([\mathcal{P};\mathcal{P}-\widetilde{\mathcal{P}}])~{},$ (3)
 […]
 Frustum-Point Fusion Module.
-As shown in Fig.
-4, the Frustum-Point (FP) Fusion Module comprises two components: 1) A Frustum-to-Point fusion that hierarchically updates per-point features, and 2) A Point-to-Frustum fusion to fuse individual features of each point within its corresponding frustum region.
-Specifically, for the point features $F_{p}$ and frustum features $F_{f}$, $F_{f}$ is back-project to the point-level features $F_{p}^{f}$ based on the projection index.
-This means that points within the same frustum region share the same frustum features.
-Then, the two features are concatenated together, followed by an MLP to update per-point features:
-$\hat{F}_{p}=\text{MLP}([F_{p}^{f};F_{p}])~{}.$ (4)
-Next, the updated per-point features $\hat{F}_{p}$ are pooled into the corresponding frustum region to further enhance the frustum features $F_{f}$ in a residual-attentive manner.
-The pooled features are denoted as $F_{f}^{p}$, which are calculated with the max operation based on the frustum region.
-The pooled features $F_{f}^{p}$ are first concatenated with $F_{f}$, followed by a simple convolutional layer to reduce the number of channels:
-$F^{fuse}=g([F_{f}^{p};F_{f}],\theta_{1})~{},$ (5)
-where $g(\cdot)$ is the mapping function and $\theta_{1}$ is the learnable parameters. Finally, the fused features are used to further enhance the frustum features in a residual-attentive manner:
-$\hat{F}_{f}=F_{f}+\sigma(h(F^{fuse},\theta_{2}))\odot F^{fuse}~{},$ (6)
-where $\sigma$ represents the sigmoid function, $h$ is the mapping function, $\theta_{2}$ is the learnable parameters, and $\odot$ indicates element-wise multiplication.
-The output features $\hat{F}_{f}$ are then used to extract larger local frustum features with subsequent convolutions.
+Following prior works [55, 58, 32], we utilize a 2D backbone composed of multiple convolution blocks that efficiently extract hierarchical frustum features.
+Each stage block is accompanied by a frustum-point fusion module, facilitating the incremental update of point and frustum features.
+As illustrated in Fig.
+4, the frustum-point fusion module consists of two essential components: 1) A frustum-to-point fusion that enables the hierarchical update of per-point features, and 2) A point-to-frustum fusion that fuses individual features of each point within its corresponding frustum region.
+The $i$-th frustum-point fusion module takes $\mathcal{F}_{p}^{i-1}$ and ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ as inputs, where ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is the output of the 2D convolution block that takes $\mathcal{F}_{f}^{i-1}$ as input.
+To efficiently extract context information of the point, we employ a frustum-to-point fusion block to integrate local features of the frustum region into points.
+Firstly, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index.
+Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update per-point features, avoiding the high-computation in neighborhood searches:
+$\mathcal{F}_{p}^{i}=\text{MLP}([\text{Inflat}(\widetilde{\mathcal{F}}_{f}^{i-1});\mathcal{F}_{p}^{i-1}])~{}.$ (4)
+As discussed above, the 2D representation is non-compact due to the sparsity of the point cloud.
+The convolution block inevitably introduces noisy information for the empty frustum, leading to compact feature representation.
+To preserve the sparsity attributes of the 2D representation, we design a point-to-frustum fusion module to fuse the sparse and dense frustum features.
+Specifically, the updated point features $\mathcal{F}_{p}^{i}$ are first pooled into the corresponding frustum region to yield a non-compact 2D representation.
+Subsequently, the non-compact frustum feature is concatenated with the compact frustum feature ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ and passed through a simple convolutional layer to reduce the number of channels:
+$\mathcal{F}_{\text{fuse}}^{i}=\text{Conv}([\text{Flat}(\mathcal{F}_{p}^{i});\widetilde{\mathcal{F}}_{f}^{i-1}])~{},$ (5)
+where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and activation function.
+Finally, the fused features are utilized to further enhance the frustum features in a residual-attentive manner:
+$\mathcal{F}_{f}^{i}=\widetilde{\mathcal{F}}_{f}^{i-1}+\sigma(h(\mathcal{F}_{\text{fuse}}^{i},\theta))\odot\mathcal{F}_{\text{fuse}}^{i}~{},$ (6)
+where $\sigma$ represents the sigmoid function, $h(\cdot)$ is a linear function with trainable parameters $\theta$, and $\odot$ indicates element-wise multiplication.
+Figure 4: Frustum-point fusion module comprises two steps: 1) A Frustum-to-Point fusion to update per-point features. 2) A Point-to-Frustum fusion to update frustum features.
 Fusion Head Module.
-Since the aim of FRNet is to classify categories for each point, the most straightforward method is to directly use the updated per-point features after the FP Fusion Module for prediction.
-Although this can also achieve great performance, we observe that this is not the optimal solution.
-The output of our FP Fusion Module consists of two features: per-point features and frustum features represented in a 2D manner, and our prediction head expects to fully utilize both of these two features.
-Additionally, the per-point features generated from the Frustum Feature Encoder Module directly embed coordinate information, which can also provide more low-level geometric information.
-The Fusion Head Module progressively fuses three different level features.
-Firstly, the frustum features $F_{f}$ are mapped to point-level features $F_{f}^{p}$.
-It is worth noting that points in the same frustum regions share the same features at this stage.
-Then, we map the features $F_{f}^{p}$ with MLPs to ensure they share the same channels as point features.
-Finally, the add operation is applied to fuse the two features.
-The same operation is also used to fuse with low-level point features $F_{p}^{l}$.
+FRNet aims to assign a category for each point in the point cloud, which requires point-level features for end-to-end label prediction.
+Although output point features at the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution.
+Features from different levels have different context-aware information.
+Lower-level features are geometric while higher-level features are semantically rich.
+Leveraging these different level features can enhance performance.
+Drawing inspiration from FIDNet [55] and CENet [58], we first combine all point and frustum features from each stage of the backbone and fuse them to reduce channel dimension:
+$\mathcal{F}_{p}^{\text{out}}=\text{MLP}([\mathcal{F}_{p}^{1};...;\mathcal{F}_{p}^{V}])~{},$ (7)
+$\mathcal{F}_{f}^{\text{out}}=\text{Conv}([\mathcal{F}_{f}^{1};...;\mathcal{F}_{f}^{V}])~{},$ (8)
+where $V$ is the number of stages in the backbone.
+Fusing backbone features with different receptive fields helps obtain more context-aware and semantic information.
+Additionally, point features $\mathcal{F}_{p}^{\text{out}}$ contains local information from neighborhood frustum regions, while frustum features $\mathcal{F}_{f}^{\text{out}}$ contains global information within the corresponding frustum region.
+Meanwhile, $\mathcal{F}_{p}^{0}$ from the frustum feature encoder module constructs graph structures within the frustum region and between neighborhood frustum regions.
+Progressively fusing these features helps extract features from local to global, and from geometric to semantic, resulting in more accurate predictions.
+To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project frustum features to point-level and use an MLP layer ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$.
+Then, an addition operation is applied to fuse the two features.
+The same operation is utilized to fuse the new features with $\mathcal{F}_{p}^{0}$.
 The overall operation can be formulated as:
-$F_{out}=\text{MLP}(\text{MLP}(F_{f}^{p})+F_{p})+F_{p}^{l}~{}.$ (7)
-Here, $F_{out}$ is used to generate the final semantic scores for each point.
+$\mathcal{F}_{\text{logit}}=\text{MLP}(\text{MLP}(\text{Inflat}(\mathcal{F}_{f}^{\text{out}}))+\mathcal{F}_{p}^{\text{out}})+\mathcal{F}_{p}^{0}~{}.$ (9)
+Here, $\mathcal{F}_{\text{logit}}$ is used to generate the final semantic scores with a linear head for the point over the entire point cloud.
 Optimization.
-To improve the optimization of FRNet, we not only utilize the commonly used point-level supervised loss function but also supervise the frustum-level features with generated pseudo frustum labels.
-Typically, a frustum region may contain multiple points, and these points often do not belong to the same category.
-Based on this observation, we generate a pseudo label for each frustum in a statistical manner.
-We count the frequency of each category in the frustum region and choose the category with the highest frequency as the pseudo label for the frustum.
-Finally, the frustum-level features will be supervised using the generated pseudo labels.
-The overall loss can be calculated as:
-$L=L_{p}+\lambda L_{f}~{},$ (8)
-where $L_{p}$ and $L_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision.
-$L_{p}$ is calculated by cross-entropy loss and $L_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42].
-Figure 4: Frustum-Point fusion module comprises two steps: 1) A Frustum-to-Point fusion to update per-point features. 2) A Point-to-Frustum fusion to update frustum features.
-3.3 Frustum-Range Operators
+In the conventional setup, a point-level loss function is commonly employed to supervise the final semantic scores.
+However, in this work, we propose optimizing frustum features with additional auxiliary heads, utilizing the frustum-level loss function.
+This approach requires generating pseudo frustum labels.
+Typically, a frustum region encompasses multiple points, which may not all belong to the same category.
+To address this, we adopt a statistical approach inspired by sparse-voxel methods [16].
+We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label of the frustum.
+Finally, the frustum-level features are supervised using the generated pseudo labels.
+The overall loss function can be formulated as follows:
+$\mathcal{L}=\mathcal{L}_{p}+\lambda\mathcal{L}_{f}~{},$ (10)
+where $\mathcal{L}_{p}$ and $\mathcal{L}_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision.
+$\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [61], and boundary loss [62].
+III-C Frustum-Range Operators
 Data augmentation enables the model to learn more robust representations.
-Commonly used techniques in LiDAR segmentation often operate on the global view, including rotation, flipping, scaling, jittering, etc.
+Common augmentations in LiDAR segmentation typically operate on the global view, encompassing rotation, flipping, scaling, jittering, etc.
 In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
 FrustumMix.
-Different from the previous LaserMix [26] and PolarMix [52] augmentations which mix two LiDAR scenes in the point level, we propose a FrustumMix to mix two scenes based on frustum units.
-Points in the scan will be first grouped into different frustum regions along inclination and azimuth directions.
-We then randomly split frustum regions equally from two directions and switch the corresponding frustum regions from two scans.
+Previous mixing strategies [44, 45] involve swapping two LiDAR scans at the point level, which may disrupt the internal structure within the frustum region.
+In this work, we propose FrustumMix, a method to swap frustum regions from two scenes.
+Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction via Eq.
+1, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$.
+We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\widetilde{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$.
+FrustumMix enhances context-awareness between frustums while preserving the invariant structure within the frustum.
 […]
 Method Venue Representation Param SemanticKITTI nuScenes ScribbleKITTI SemanticPOSS
-FPS normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ Val normal-$\uparrow$ Test normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$ mIoU normal-$\uparrow$ mAcc normal-$\uparrow$
-MinkUNet [10] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M
-$\mathbf{9.1}$
-$63.7$
-$55.0$
-$\mathbf{53.1}$
-$\mathbf{68.3}$
-SPVNAS [47] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$
-$\underline{66.4}$
-$\underline{56.9}$
-$48.4$
-$61.5$
-Cylinder3D [65] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M
-$6.2$
-$\mathbf{65.9}$
-$\mathbf{67.8}$
-$\mathbf{76.1}$
-$\mathbf{77.9}$
-$\mathbf{57.0}$
-$\underline{52.9}$
-$\underline{64.9}$
-RPVNet [54] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$
-$65.5$
-$70.3$
-$77.6$
-2DPASS [56] ECCV’22 Multi-View $\bullet$ $26.5$ M
-$\mathbf{8.4}$
-$69.3$
-$72.2$
-$\mathbf{79.4}$
-$80.8$
-SphereFormer [27] CVPR’23 Multi-View $\bullet$ $32.3$ M
-$4.9$
-$67.8$
-$74.8$ $78.4$
-$\underline{81.9}$
-UniSeg [32] ICCV’23 Multi-View $\bullet$ $147.6$ M
-$6.9$
-$\mathbf{71.3}$
-$\mathbf{75.2}$
-$\mathbf{83.5}$
-RangeNet++ [35] IROS’19 Range View $\bullet$ $50.4$ M
-$12.0$
-$50.3$
-$52.2$
-$65.5$
-$44.6$
-$57.8$
-$30.9$
-FIDNet [64] IROS’20 Range View $\bullet$ $6.1$ M
-$31.8$
-$58.9$
-$59.5$
-$54.1$
-$65.4$
-$46.4$
-CENet [8] ICME’22 Range View $\bullet$ $6.8$ M $33.4$
-$62.6$
-$64.7$
-$55.7$
-$66.8$
-$50.3$
-RangeViT [4] CVPR’23 Range View $\bullet$ $23.7$ M
-$10.0$
-$60.7$
-$64.0$
-$75.2$
-$53.6$
-$66.5$
-RangeFormer [24] ICCV’23 Range View $\bullet$ $24.3$ M
-$6.2$
-$67.6$
-$\mathbf{73.3}$
-$78.1$
-$80.1$
-$63.0$
-Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M
-$\mathbf{33.8}$
-$67.1$
-$\underline{72.5}$
-$78.8$ $82.1$
-$62.4$
-$71.2$ $52.5$ $67.1$
-FRNet Ours Frustum-Range $\bullet$ $10.0$ M
-$29.1$
-$\mathbf{68.7}$
-$\mathbf{73.3}$
-$\mathbf{79.0}$
-$\mathbf{82.5}$
-$\mathbf{63.1}$
-$\mathbf{72.3}$
-$\mathbf{53.5}$
-$\mathbf{68.1}$
+FPS $\uparrow$ Val $\uparrow$ Test $\uparrow$ Val $\uparrow$ Test $\uparrow$ mIoU $\uparrow$ mAcc $\uparrow$ mIoU $\uparrow$ mAcc $\uparrow$
+MinkUNet [17] CVPR’19 Sparse Voxel $\bullet$ $21.7$ M $9.1$ $62.8$ $63.7$ $75.8$ $55.0$ $\mathbf{53.1}$ $\mathbf{68.3}$
+SPVNAS [15] ECCV’20 Sparse Voxel $\bullet$ $21.8$ M $8.9$ $62.5$ $66.4$ $74.4$ $56.9$ $48.4$ $61.5$
+Cylinder3D [16] CVPR’21 Sparse Voxel $\bullet$ $55.9$ M $6.2$ $65.9$ $67.8$ $\mathbf{76.1}$ $\mathbf{77.9}$ $\mathbf{57.0}$ $52.9$ $64.9$
+PVKD [65] CVPR’22 Sparse Voxel $\bullet$ $14.1$ M $\mathbf{13.2}$ $\mathbf{66.4}$ $\mathbf{71.2}$ $76.0$
+KPConv [7] ICCV’19 Raw Points $\bullet$ $18.3$ M $3.2$ $61.3$ $58.8$
+RandLA-Net [8] CVPR’20 Raw Points $\bullet$ $1.2$ M $1.9$ $57.1$ $53.9$
+PTv2 [66] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $\mathbf{70.3}$ $\mathbf{72.6}$ $\mathbf{80.2}$ $\mathbf{82.6}$
+WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $\mathbf{7.4}$ $68.0$ $70.8$ $79.1$
+RPVNet [21] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$ $65.5$ $70.3$ $77.6$
+2DPASS [67] ECCV’22 Multi-View $\bullet$ $26.5$ M $\mathbf{8.4}$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
+SphereFormer [68] CVPR’23 Multi-View $\bullet$ $32.3$ M $4.9$ $67.8$ $74.8$ $78.4$ $81.9$
+UniSeg [23] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $\mathbf{71.3}$ $\mathbf{75.2}$ $\mathbf{83.5}$
+RangeNet++ [30] IROS’19 Range View $\bullet$ $50.4$ M $12.0$ $50.3$ $52.2$ $65.5$ $44.6$ $57.8$ $30.9$
+FIDNet [55] IROS’20 Range View $\bullet$ $6.1$ M $31.8$ $58.9$ $59.5$ $71.4$ $54.1$ $65.4$ $46.4$
+CENet [58] ICME’22 Range View $\bullet$ $6.8$ M $33.4$ $62.6$ $64.7$ $73.3$ $55.7$ $66.8$ $50.3$
+RangeViT [31] CVPR’23 Range View $\bullet$ $23.7$ M $10.0$ $60.7$ $64.0$ $75.2$ $53.6$ $66.5$
+RangeFormer [32] ICCV’23 Range View $\bullet$ $24.3$ M $6.2$ $67.6$ $\mathbf{73.3}$ $78.1$ $80.1$ $63.0$
+Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M $\mathbf{33.8}$ $67.1$ $72.5$ $78.8$ $82.1$ $62.4$ $71.2$ $52.5$ $67.1$
+FRNet Ours Frustum-Range $\bullet$ $10.0$ M $29.1$ $\mathbf{68.7}$ $\mathbf{73.3}$ $\mathbf{79.0}$ $\mathbf{82.5}$ $\mathbf{63.1}$ $\mathbf{72.3}$ $\mathbf{53.5}$ $\mathbf{68.1}$
+RangeInterpolation.
+Previous range-view methods [58, 55, 32] often encounter a large number of empty pixels due to the sparsity of LiDAR points, resulting in noisy information.
+To mitigate this problem, we introduce RangeInterpolation, a technique that utilizes the range image to reconstruct surfaces in the LiDAR scan.
+Specifically, the point cloud $\mathcal{P}$ is first projected onto the range image $\mathcal{R}\in\mathbb{R}^{H\times W\times(3+L)}$.
+For empty pixel position $(u,v)$ in the range image $\mathcal{R}$, we establish a window with size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd.
+Our aim is to interpolate the point falling into $(u,v)$ using the surrounding range information within the predefined window.
+Considering that not all pixels in the window contain valid values, we aggregate the information from non-empty pixels within the window for point interpolation using an average operation.
+During the training procedure, the semantic label of the interpolated point is determined via threshold-guided voting within the window.
+Specifically, we identify the most frequent category within the window and its frequency.
+If the frequency is below a threshold, we classify the interpolated point as located at the boundary and assign the label as ignored, which will not contribute to loss calculation.
+Otherwise, the label is determined by the category with the highest frequency.
+IV Experiments
+In this section, we demonstrate the scalability and robustness of the proposed method.
+We first introduce the datasets, metrics, and detailed implementations we used.
+Subsequently, we show that FRNet achieves an optimal trade-off between efficiency and accuracy when compared to state-of-the-art methods across various task setups.
+Finally, we conduct a series of ablation studies to analyze each component in FRNet.
 […]
@@
 FrustumMix Ours $55.7$ $62.5$ $63.0$ $64.9$ $60.0$ $70.0$ $72.6$ $74.1$ $45.6$ $55.7$ $58.2$ $60.8$
-Sup.-only - FRNet (Ours) $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
+Sup.-only - FRNet $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
+PolarMix [45] NeurIPS’22 $50.1$ $60.9$ $62.0$ $63.8$ $55.6$ $69.6$ $71.0$ $73.8$ $43.2$ $55.0$ $56.1$ $57.3$
 LaserMix [44] CVPR’23 $52.9$ $62.9$ $63.2$ $65.0$ $58.7$ $71.5$ $72.3$ $75.0$ $45.8$ $56.8$ $57.7$ $59.0$
 FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $\mathbf{65.2}$ $\mathbf{65.4}$ $\mathbf{61.2}$ $\mathbf{72.2}$ $\mathbf{74.6}$ $\mathbf{75.4}$ $\mathbf{46.6}$ $57.0$ $59.5$ $\mathbf{61.2}$
-Range-Interpolation.
-Previous range-view methods often suffer from large number of empty pixels, which still exist in FRNet.
-To alleviate this problem, we propose to interpolate some potential points based on the range image.
-Specifically, LiDAR points will be first projected to range images as previous works do.
-We then scan all pixels in the range images and find the empty pixels.
-For each empty pixel, we will judge whether there are projected points in the pixels along its left and right directions.
-When both the two pixels have projected points, the interpolated points of this pixel can be calculated by averaging the coordinates of two points.
-During the training procedure, we also need to generate semantic labels for the interpolated points, which will face two cases.
-If both right and left points have the same semantic labels, we will set the semantic label of the interpolated point to be the same as them.
-For another case with different semantic labels, we consider the interpolated points as boundaries of the object, and it will be set to an ignored label, which will not be used to calculate loss.
-4 Experiments
-4.1 Settings
+Figure 5: Class-wise LiDAR segmentation results of FRNet and the baseline model on the test set of SemanticKITTI [28].
+IV-A Experimental Settings
 Datasets.
-We conduct experiments on four popular LiDAR segmentation datasets.
-SemanticKITTI [5] consists of $22$ sequences with a total of $43,552$ LiDAR scans.
-Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing with $19$ semantic classes.
-nuScenes [13] provides $1,000$ driving scenes captured with much sparser points by 32-beam LiDAR sensors, and 16 classes are evaluated during inference.
-SemanticPOSS [37] is a more challenging benchmark collected by Peking University with much smaller and sparser point clouds.
-It consists of $2,988$ LiDAR scenes with a large number of dynamic instances.
-Officially, it is divided into $6$ parts, and part $2$ is used to evaluate the performance with $13$ categories.
-ScribbleKITTI [49] shares the same scenes with SemanticKITTI [5] but only provides weak annotations with line scribbles, which means around $8.06\%$ labeled points available during training.
+We conduct a comprehensive experiment on four popular LiDAR segmentation datasets.
+SemanticKITTI [28] is a large-scale outdoor dataset that comprises $22$ sequences collected from various scenes in Karlsruhe, Germany, using a Velodyne HDL-64E LiDAR sensor.
+Officially, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) for validation, and sequences $11$ to $21$ ($20,351$ scans) for online testing.
+The original dataset is annotated with $28$ classes, though only $19$ merged classes are used for single-scan evaluation.
+The vertical FOV is $[-25^{\circ},3^{\circ}]$.
+nuScenes [43] is a multimodal dataset that is widely used in autonomous driving scenarios.
+It contains LiDAR point clouds captured around streets in Boston and Singapore, provided by 32-beam LiDAR sensors.
+The dataset includes $1,000$ driving scenes with sparser points.
+It is annotated with $32$ classes and only $16$ semantic categories are used for evaluation.
+The vertical FOV is $[-30^{\circ},10^{\circ}]$.
+SemanticPOSS [64] is a more challenging benchmark collected at Peking University and features much smaller and sparser point clouds.
+It consists of $2,988$ LiDAR scenes with numerous dynamic instances.
+Officially, it is divided into $6$ sequences, with sequence $2$ used for evaluation with $13$ merged categories.
+The vertical FOV is $[-16^{\circ},7^{\circ}]$.
+ScribbleKITTI [63] shares the same scenes with SemanticKITTI [28] but only provides weak annotations with line scribbles.
+It contains $189$ million labeled points, with approximately $8.06\%$ labeled points available during training.
+Additionally, we conduct a comprehensive out-of-distribution robustness evaluation experiment on SemanticKITTI-C and nuScenes-C from the Robo3D [71] benchmark.
+Each dataset contains eight corruption types, including “fog”, “wet ground”, “snow”, “motion blur”, “beam missing”, “crosstalk”, “incomplete echo”, and “cross sensor”.
 Evaluation Metrics.
-In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to access the performance.
-IoU and Acc can be calculated as:
-$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},~{}~{}\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (9)
-where $\text{TP}_{i},\text{TP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively.
+In line with standard protocols, we utilize Intersection-over-Union (IoU) and Accuracy (Acc) for each class $i$ and compute the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) to assess performance.
+IoU and Acc can be calculated as follows:
+$\text{IoU}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}+\text{FN}_{i}}~{},\text{Acc}_{i}=\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}~{},$ (11)
+where $\text{TP}_{i},\text{FP}_{i},\text{FN}_{i}$ are true-positives, false-positives, and false-negatives for class $i$, respectively.
 For robustness evaluation, we follow the practice of Robo3D [71] to use Corruption Error (CE) and Resilience Rate (RR) to evaluate the robustness of FRNet.
-The CE and RR can be calculated as:
-$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},~{}~{}\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (10)
-where $\text{IoU}_{i,l}$ is the task-specific IoU; $\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ are scores of the baseline model and scores on the clean evaluation set.
+The CE and RR are calculated as follows:
+$\text{CE}_{i}=\frac{\sum_{l=1}^{3}(1-\text{IoU}_{i,l})}{\sum_{l=1}^{3}(1-\text{IoU}_{i,l}^{\text{base}})}~{},\text{RR}_{i}=\frac{\sum_{l=1}^{3}\text{IoU}_{i,l}}{3\times\text{IoU}_{\text{clean}}}~{},$ (12)
+where $\text{IoU}_{i,l}$ represents the task-specific IoU.
+$\text{IoU}_{i,l}^{\text{base}}$ and $\text{IoU}_{\text{clean}}$ denote scores of the baseline model and scores on the “clean” evaluation set, respectively.
 Implementation Details.
-We implement FRNet using the MMDetection3D [11] framework and utilize CENet [8] as our backbone to extract frustum features.
-We set the shape of the 2D representation to $64\times 512$ for SemanticKITTI [5] and ScribbelKITTI [49] and $32\times 480$ for nuScenes [13] and SemanticPOSS [37], respectively.
-We choose AdamW [34] as our default optimizer and utilize the OneCycle scheduler [45] with $lr=0.01$.
-The models are trained for $50$ epochs on SemanticKITTI [5] and ScribbelKITTI [49], and $80$ epochs on nuScenes [13] and SemanticPOSS [37], respectively, with a batch size of $2$ with $4$ GPUs.
-4.2 Comparative Study
+We implement FRNet using the popular MMDetection3D [72] codebase and employ CENet [58] as the 2D backbone to extract frustum features.
+The 2D representation shape is set to $64\times 512$ for SemanticKITTI [28] and ScribbelKITTI [63], and $32\times 480$ for nuScenes [43] and SemanticPOSS [64], respectively.
+For optimization, We choose AdamW [73] as our default optimizer with an initial learning rate of $0.01$.
+We utilized the OneCycle scheduler [74] to dynamically adjust the learning rate during training.
+All models are trained using four GPUs for $50$ epochs on SemanticKITTI [28] and ScribbelKITTI [63], and $80$ epochs on nuScenes [43] and SemanticPOSS [64], respectively.
+The batch size is set to $4$ for each GPU.
+Furthermore, we implement a faster variant, dubbed Fast-FRNet, which features a smaller 2D representation resolution and scaled-down 2D backbone.
+Specifically, the size of the 2D representation is adjusted to $32\times 360$ for all datasets, and the 2D backbone is a variant of ResNet18 [75], following the design of CENet [58].
+IV-B Comparative Study
 Comparisons to State of the Arts.
-Firstly, we compare the proposed FRNet with state-of-the-art LiDAR segmentation methods from different representations on both fully-supervised and weakly-supervised benchmarks.
+We begin by comparing FRNet with state-of-the-art LiDAR segmentation methods across different representations, including sparse-voxel, point-view, multi-view, and range-view representations, on both fully-supervised and weakly-supervised benchmarks.
 Tab.
-1 summarizes all the results and the FPS of all the models is evaluated with a single GeForce RTX 2080Ti GPU for fair comparison.
-Fast-FRNet takes resnet18 [19] as the 2D backbone and the shapes of 2D representation on four datasets are $32\times 360$.
-Compared with multi-view methods such as SphereFormer [27] and UniSeg [32], although FRNet does not have a significant advantage in performance, its parameters and FPS are far superior to theirs.
-Compared with voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency, and Fast-FRNet achieves the fastest inference speed with only a little drop in performance.
-Fig.
-6 also shows qualitative results among state-of-the-art methods, including CENet [8] and RangeViT [4] on the val set of SemanticKITTI [5].
-FRNet reduces prediction errors over large areas, especially for vegetation and parking.
-Label-Efficient LiDAR Segmentation.
+I provides a summary of all the results.
+We evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for fair comparison.
+For SemanticKITTI [28] and nuScenes [43], we report mIoU scores on both the val and test sets.
+For ScribbleKITTI [63] and SemanticPOSS [64], which share the same data on val and test sets, we report both mIoU and mAcc scores on the val set.
+Compared with multi-view methods, although FRNet does not exhibit a significant advantage in performance, its parameters and FPS are notably superior.
+In comparison with sparse-voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency.
+Additionally, Fast-FRNet demonstrates the fastest inference speed with only a minor drop in performance.
+[…]
+# Method Venue mCE $\downarrow$ mRR $\uparrow$ mIoU $\uparrow$ Fog Weg Snow Motion Beam Cross Echo Sensor
+SemanticKITTI-C MinkUNet [17] CVPR’19 $100.0$ $81.9$ $62.8$ $\mathbf{55.9}$ $54.0$ $53.3$ $32.9$ $56.3$ $\mathbf{58.3}$ $54.4$ $46.1$
+SPVCNN [15] ECCV’20 $100.3$ $\mathbf{82.2}$ $62.5$ $55.3$ $54.0$ $51.4$ $34.5$ $56.7$ $58.1$ $54.6$ $46.0$
+FIDNet [55] IROS’20 $113.8$ $77.0$ $58.8$ $43.7$ $51.6$ $49.7$ $40.4$ $49.3$ $49.5$ $48.2$ $29.9$
+Cylinder3D [16] CVPR’21 $103.3$ $80.1$ $63.4$ $37.1$ $57.5$ $46.9$ $52.5$ $57.6$ $56.0$ $52.5$ $46.2$
+2DPASS [67] ECCV’22 $106.1$ $77.5$ $64.6$ $40.5$ $60.7$ $48.5$ $\mathbf{57.8}$ $58.8$ $28.5$ $55.8$ $50.0$
+CENet [58] ICME’22 $103.4$ $81.3$ $62.6$ $42.7$ $57.3$ $53.6$ $52.7$ $55.8$ $45.4$ $53.4$ $45.8$
+WaffleIron [11] ICCV’23 $109.5$ $72.2$ $66.0$ $45.5$ $58.6$ $49.3$ $33.0$ $59.3$ $22.5$ $\mathbf{58.6}$ $\mathbf{54.6}$
+FRNet Ours $\mathbf{96.8}$ $80.0$ $\mathbf{67.6}$ $47.6$ $\mathbf{62.2}$ $\mathbf{57.1}$ $56.8$ $\mathbf{62.5}$ $40.9$ $58.1$ $47.3$
+nuScenes-C MinkUNet [17] CVPR’19 $100.0$ $74.4$ $75.8$ $53.6$ $73.9$ $40.4$ $\mathbf{73.4}$ $\mathbf{68.5}$ $26.6$ $\mathbf{63.8}$ $\mathbf{51.0}$
+SPVCNN [15] ECCV’20 $106.7$ $74.7$ $74.4$ $59.0$ $72.5$ $41.1$ $58.4$ $65.4$ $36.8$ $62.3$ $49.2$
+FIDNet [55] IROS’20 $122.4$ $73.3$ $71.4$ $64.8$ $68.0$ $59.0$ $48.9$ $48.1$ $57.5$ $48.8$ $23.7$
+Cylinder3D [16] CVPR’21 $111.8$ $72.9$ $76.2$ $59.9$ $72.7$ $58.1$ $42.1$ $64.5$ $44.4$ $60.5$ $42.2$
+2DPASS [67] ECCV’22 $\mathbf{98.6}$ $75.2$ $\mathbf{77.9}$ $64.5$ $\mathbf{76.8}$ $54.5$ $62.0$ $67.8$ $34.4$ $63.2$ $45.8$
+CENet [58] ICME’22 $112.8$ $76.0$ $73.3$ $67.0$ $69.9$ $61.6$ $58.3$ $50.0$ $\mathbf{60.9}$ $53.3$ $24.8$
+WaffleIron [11] ICCV’23 $106.7$ $72.8$ $76.1$ $56.1$ $73.9$ $49.6$ $59.5$ $65.2$ $33.1$ $61.5$ $44.0$
+FRNet Ours $\mathbf{98.6}$ $\mathbf{77.5}$ $77.7$ $\mathbf{69.1}$ $76.6$ $\mathbf{69.5}$ $54.5$ $68.3$ $41.4$ $58.7$ $43.1$
+Label-Efficient LiDAR Semantic Segmentation.
 Recently, semi-supervised learning has been explored in LiDAR segmentation.
-In this work, we follow the paradigm of LaserMix [26] to conduct FRNet on the semi-supervised segmentation task.
-Here, we use FRNet as the backbone and the proposed FrustumMix to mix labeled and unlabeled.
-Tab.
-2 summarizes the results of recent methods.
-Results show that our FRNet backbone achieves appealing performance among state-of-the-art methods.
-Meanwhile, we compare the proposed FrustumMix with LaserMix among different backbones, including Cylinder3D [65] and FRNet.
-Results show that the proposed FrustumMix outperforms LaserMix, even in voxel representation.
-Robustness.
-Robustness reflects the model’s ability to generalization to a certain extent, which is particularly important in autonomous driving, especially in extreme weather conditions.
-To this end, we also evaluate the robustness of FRNet using the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [25], which include 8 different conditions and each condition has 3 different levels.
-The models trained with SemanticKITTI [5] and nuScenes [13] are used to evaluate the performance on SemanticKITTI-C and nuScenes-C [25], respectively.
+In this work, we adopt the approach of LaserMix [44] to apply FRNet to the semi-supervised segmentation task, as shown in Tab.
+II.
+Firstly, we utilize Cylinder3D [16] as the backbone and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning.
+Notably, FrustumMix demonstrates superior performance compared to the LaserMix [44] method.
+Subsequently, we employ FRNet as the backbone and assess various strategies for mixing labeled and unlabeled scans, including PolarMix [45], LaserMix [44], and our FrustumMix.
+The Sup.-only baseline results are obtained without FrustumMix and RangeInterpolation as data augmentation techniques.
+Results show that FrustumMix consistently achieves the best scores across all data splits, particularly with limited annotations.
+Out-of-Distribution Robustness.
+Robustness, which reflects a model’s ability to generalize under different conditions, is crucial for automotive security systems, particularly in extreme weather conditions [76, 77].
+To evaluate the out-of-distribution robustness of FRNet, we utilize the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [71].
+The models trained on SemanticKITTI [28] and nuScenes [43] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [71], respectively.
 As shown in Tab.
-3, FRNet demonstrates strong robustness, especially under extreme weather conditions.
-[…]
-Method Venue SemKITTI-C nuScenes-C
-mCE normal-$\downarrow$ mRR normal-$\uparrow$ mCE normal-$\downarrow$ mRR normal-$\uparrow$
-MinkUNet [10] CVPR’19 $100.0$
-$\mathbf{81.9}$
-$100.0$
-$74.4$
-Cylinder3D [65] CVPR’21
-$103.3$
-$80.1$
-$111.8$
-$72.9$
-2DPASS [56] ECCV’22
-$106.1$
-$77.5$
-$\mathbf{98.6}$
-$75.2$
-FIDNet [64] IROS’20
-$113.8$
-$77.0$
-$122.4$
-$73.3$
-CENet [8] ICME’22
-$103.4$
-$81.3$
-$112.8$
-$76.0$
-FRNet Ours
-$\mathbf{96.8}$
-$80.0$
-$\mathbf{98.6}$
-$\mathbf{77.5}$
-Figure 5: Ablation study on different configurations in FRNet.
-Tables show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [5].
-b) Different mixing strategies on the val set of nuScenes [13].
-c) Different mixing strategies on the val set of SemanticKITTI [5].
-Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c).
-Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [8, 4].
-To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively.
-Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
-4.3 Ablation Study
-In this section, we will discuss each component in FRNet on the val set of SemanticKITTI [5] and nuScenes [13].
-Model Analysis.
-We first analyze different components in FRNet, including the Frustum Feature Encoder, Frustum-Point (FP) Fusion, Fusion Head, Frustum-level Supervision, and Range-Interpolation.
-The results are summarized in Tab.
-4.
-As the baseline of FRNet, the Frustum Feature Encoder module extracts per-point features in the frustum region and predicts semantic labels of each point with the supervision of point-level labels, and the performance is $62.7\%$ and $72.8\%$ on SemanticKITTI [5] and nuScenes [13], respectively.
-When we use the FP Fusion module to update per-point features hierarchically, the performance improves by $1.6\%$ and $0.8\%$ on the two datasets, respectively.
-With the frustum-level supervision, the performance further improves by $1.9\%$ and $2.8\%$.
-Next, we utilize a Fusion Head module to fuse different level features for predicting, and it improves by $0.8\%$ on both datasets.
-To alleviate the empty pixels in 2D representations, we propose a Range-Interpolation method to generate some potential points.
-Results show that this augmentation can also help to improve the performance by $0.6\%$ and $0.5\%$.
-Finally, following most works, we also use Test Time Augmentation (TTA) during inference, and it will bring an improvement of $1.1\%$ and $1.3\%$, respectively.
+III, FRNet achieves promising performance across most corruption types and demonstrates superiority over recent works employing various LiDAR representations, including sparse voxel [17, 15, 16], range view [55, 58], multi-view fusion [67], and raw points [11].
+Figure 6: Qualitative results among state-of-the-art LiDAR segmentation methods [55, 58, 31] on the val set of SemanticKITTI [28].
+To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and red, respectively.
+Best viewed in colors and zoomed-in for details.
 […]
@@
 mIoU mAcc mIoU mAcc
-$62.7$
-$72.3$
-$72.8$
-$82.7$
-$64.3$
-$73.5$
-$73.6$
-$83.4$
-$66.2$
-$73.9$
-$76.4$
-$84.6$
-$67.0$
-$74.1$
-$77.2$
-$84.8$
-$67.6$
-$74.2$
-$77.7$
-$\mathbf{85.2}$
-$\mathbf{68.7}$
-$\mathbf{74.9}$
-$\mathbf{79.0}$
-$85.0$
-Mixing Strategy.
-Although we have validated the effectiveness of our proposed FrustumMix in semi-supervised segmentation, we still compare FrustumMix with different popular outdoor mix strategies, including LaserMix [26], PolarMix [52].
-As shown in Fig.
-5 (b) and (c), the proposed FrustumMix significantly improves the performance and outperforms existing mixing strategies on both datasets.
-Range-Interpolation Strategy.
-Range-Interpolation synthesizes possible points to reduce empty pixels in 2D representations.
-For the empty pixel in the range image, we initially use a $3\times 3$ window to pick its neighboring pixels and average valid pixel information as the interpolated point.
-Unfortunately, this results in a $0.4\%$ drop in performance.
-When we simply average the effective pixels, the generated points might fall out of the expected frustum region due to the empty surroundings.
-Based on this observation, we re-think the interpolation strategy.
-For an empty pixel, only when there are projection points on both the left and right or above and below, the calculated interpolation point can fall in the expected region.
-To this end, we propose two interpolation strategies along horizontal and vertical directions and improve the results by $0.4\%$ and $0.6\%$, respectively.
-It seems that points along horizontal directions share similar attributes as they line in the same laser beams.
-Backbones.
-To validate the versatility of our method, we select different range-view methods as our 2D backbone to extract frustum features, including SalsaNeXt [12], FIDNet [64], and CENet [8].
-As shown in Fig.
-5 (a), our proposed frustum-range representation can bring an improvement of $5.0\%$ to $6.8\%$ on different range-view methods.
-5 Conclusion
-In this work, we presented a scalable Frustum-Range network (FRNet) for LiDAR segmentation, which can be seamlessly integrated into any range-view method.
-FRNet comprises three main components: the Frustum Feature Encoder Module to extract per-point features in low-level, the Frustum-Point Fusion Module to update per-point features hierarchically, and the Fusion Head Module to fuse different level features to predict more accurate labels.
-We also introduced two efficient augmentations, FurstumMix and Range-Interpolation, to improve the generalization of FRNet.
-Extensive experiments showed that our FRNet achieves competitive performance while maintaining high efficiency.
+$62.7$ $72.3$ $72.8$ $82.7$
+$64.3$ $73.5$ $73.6$ $83.4$
+$66.2$ $73.9$ $76.4$ $84.6$
+$67.0$ $74.1$ $77.2$ $84.8$
+$67.6$ $74.2$ $77.7$ $\mathbf{85.2}$
+$\mathbf{68.7}$ $\mathbf{74.9}$ $\mathbf{79.0}$ $85.0$
+Quantitative Assessments.
+We compare the class-wise IoU scores of FRNet with those of the baseline method [58] in Fig.
+5.
+Notably, FRNet demonstrates significant improvement across most semantic classes, particularly for dynamic classes with intricate structures, where it achieves mIoU gains ranging from $15\%$ to $24\%$.
+In Fig.
+6, we present the prediction results on SemanticKITTI [28], compared with those of state-of-the-art range view methods, including FIDNet [55], CENet [58], and RangeViT [31].
+It is evident that FRNet delivers more accurate predictions, particularly on large planar objects such as “vegetation”, “terrain”, and “sidewalk”, which are often challenging to distinguish accurately.
+IV-C Ablation Study
+In this section, we will discuss the efficacy of each design in our FRNet architecture.
+Unless otherwise specified, all experiments are conducted and reported on the val set of SemanticKITTI [28] and nuScenes [43], respectively.
+Component Designs.
+Tab.
+IV summarizes the ablation results of each component in the FRNet architecture.
+Firstly, We directly apply the frustum feature encoder to CENet [58], incorporating KNN post-processing to predict semantic labels over the entire point cloud, yielding $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [28] and nuScenes [43], respectively.
+Subsequently, by integrating the point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in an improvement of $1.6\%$ and $0.8\%$ mIoU, respectively.
+To regularize the frustum features, the frustum-level supervision is introduced and it leads to significant gains of $1.9\%$ and $2.8\%$ mIoU, respectively.
+The fusion head module, which fuses multiple features at different levels for more accurate prediction, brings an additional $0.8\%$ mIoU gain on both datasets.
+Furthermore, the proposed RangeInterpolation, aimed at alleviating empty pixels in the 2D representation by reconstructing semantic surfaces via surrounding range information, provides a performance boost of approximately $0.5\%$ mIoU.
+Finally, adopting Test Time Augmentation during inference, following prior works, brings an improvement of $1.1\%$ and $1.3\%$ mIoU, respectively.
+Figure 7: Ablation study on different FRNet configurations.
+Plots show the comparisons of: a) Different range-view backbones on the val set of SemanticKITTI [28].
+b) Different mixing strategies on the val set of nuScenes [43].
+c) Different mixing strategies on the val set of SemanticKITTI [28].
+Color-filled boxes denote mAcc while empty boxes denote mIoU in (b) and (c), respectively.
+Mixing Strategies.
+We further explore the effectiveness of the proposed FrustumMix.
+As shown in Tab.
+II, FrustumMix significantly improves performance on semi-supervised learning tasks and demonstrates generalizability to other modalities.
+To contextualize FrustumMix within common data augmentation techniques, we compare it with recent popular outdoor mixing strategies, including LaserMix [44] and PolarMix [45].
+As illustrated in Fig.
+7(b) and Fig.
+7(c), FrustumMix consistently outperforms these prior strategies.
+TABLE V: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [28].
+Green pixel denotes the empty pixel needed to be interpolated.
+Violet pixels are the surrounding pixels used to generate the point.
+All mIoU scores are given in percentage (%).
+Baseline $1\times 3$ $3\times 1$ $1\times 5$ $5\times 1$
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$67.0$ $67.6_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.6)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$ $67.5_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.5)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$
+cross $3\times 3$ cross $5\times 5$ $3\times 5$ $3\times 3$ $5\times 5$
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+$67.5_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.5)}}}}$ $67.4_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.4)}}}}$ $67.1_{\mathbf{{{\color[rgb]{0.3046875,0.62890625,0.34375}\definecolor[named]{pgfstrokecolor}{rgb}{0.3046875,0.62890625,0.34375}(+0.1)}}}}$ $66.8_{\mathbf{{{\color[rgb]{0.546875,0.33203125,0.609375}\definecolor[named]{pgfstrokecolor}{rgb}{0.546875,0.33203125,0.609375}(-0.2)}}}}$ $66.6_{\mathbf{{{\color[rgb]{0.546875,0.33203125,0.609375}\definecolor[named]{pgfstrokecolor}{rgb}{0.546875,0.33203125,0.609375}(-0.4)}}}}$
+RangeInterpolation Window Settings.
+RangeInterpolation reconstructs semantic surfaces based on the range image, facilitating a compact 2D representation for coherent semantic learning.
+Tab.
+V showcases the performance of RangeInterpolation with different window settings.
+Notably, We observe consistent performance improvements when considering interpolation along the horizontal or vertical direction exclusively.
+Specifically, configurations employing horizontal windows, such as the $1\times 3$ and $1\times 5$ settings, demonstrate superior results compared to those focusing on the vertical direction ($3\times 1$ and $5\times 1$ windows).
+We conjecture that points aligned along the same laser beams often share similar attributes, leading to more accurate and coherent semantic surfaces.
+Conversely, larger window sizes, such as the $3\times 5$, $3\times 3$, and $5\times 5$ configurations, tend to exhibit non-uniform point distributions within the window, resulting in noisy points that fail to accurately represent the expected region and consequently leading to a drop in performance.
+2D Backbones.
+To validate the versatility of our method, we employ various range-view methods as our 2D backbone for extracting frustum features, including SalsaNext [57], FIDNet [55], and CENet [58].
+As depicted in Fig.
+7(a), our proposed frustum-range representation consistently delivers performance gains ranging from $5.0\%$ to $6.8\%$ in mIoU across different range-view methods.
+V Conclusion
+In this work, we presented FRNet, a Frustum-Range network designed for efficient and effective LiDAR segmentation, which can be seamlessly integrated into range-view approaches.
+FRNet comprises three key components: the Frustum Feature Encoder Module for extracting per-point features at lower levels, the Frustum-Point Fusion Module for hierarchical updating of per-point features, and the Fusion Head Module for fusing features from different levels to predict more accurate labels.
+Additionally, We introduced two efficient augmentation methods, FurstumMix and RangeInterpolation, to enhance the generalizability of range-view LiDAR segmentation.
+Extensive experiments across diverse driving perception datasets verified that FRNet achieves competitive performance while maintaining high efficiency.
+We hope this work can lay a solid foundation for future works targeted at real-time, in-vehicle 3D perception.
+Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction.
 Appendix
-[…]
-[…]
-Appendix A Additional Implementation Details
-In this section, we provide additional details regarding the implementation and reproduction details of the proposed FRNet and Fast-FRNet.
-A.1 Datasets
-We adopt the following four LiDAR segmentation datasets to validate the effectiveness and superiority of the proposed approach.
-  • SemanticKITTI [5] is a large-scale outdoor dataset which is based on KITTI odometry benchmark [14].
-It consists of $22$ sequences, collected from different scenes in Karlsruhe, Germany by Velodyne HDL-64E.
-Commonly, sequences $00$ to $07$ and $09$ to $10$ ($19,130$ scans) are used for training, sequence $08$ ($4,071$ scans) is used for validation, and sequences $11$ to $21$ ($20,351$ scans) are used for testing.
-The original dataset is annotated with $28$ classes and only $19$ semantic classes are used for single-scan evaluation.
-The vertical field-of-view is $[-25^{\circ},3^{\circ}]$.
-  • nuScenes [13] is a multimodal dataset that is widely used in autonomous driving.
-The LiDAR point clouds are collected around the streets in Boston and Singapore by Velodyne HDL-32E.
-It comprises $1,000$ driving scenes, in which $700$ scenes ($28,130$ scans) for training, $150$ scenes ($6,019$ scans) for validation, and $150$ scenes ($6,008$ scans) for testing.
-The dataset is annotated with $32$ classes and only $16$ semantic classes are used for evaluation.
-The vertical field-of-view is $[-30^{\circ},10^{\circ}]$.
-  • ScribbleKITTI [49] is a scribble annotated dataset with line drawings, which shares the same LiDAR point clouds with SemanticKITTI [5].
-It contains $189$ million labeled points, around $8.06\%$ of the total training points.
-This dataset is suitable for conducting weakly-supervised LiDAR segmentation, which is useful for reducing the annotation cost in existing LiDAR segmentation systems.
-  • SemanticPOSS [37] is a more challenging dataset collected in Peking University by Pandora.
-The dataset includes $6$ sequences.
-Commonly, sequence $02$ ($500$ scans) is used for validation, and others ($2,488$ scans) are used for training.
-The dataset is annotated with $14$ classes and only $13$ semantic classes are used for evaluation.
-The vertical field-of-view is $[-16^{\circ},7^{\circ}]$.
-A.2 Hyperparameters
-For each point with coordinate $(x,y,z)$, its corresponding frustum region $(u,v)$ in the 2D representation with shape of $(H,W)$ is calculated as:
-$\binom{u}{v}=\binom{\frac{1}{2}[1-\arctan(y,x)\pi^{-1}]W}{[1-(\arcsin(zd^{-1})+f_{up})f^{-1}]H}~{},$ (11)
-where $d=\sqrt{x^{2}+y^{2}+z^{2}}$ represents the depth of a point, $f=|f_{up}|+|f_{down}|$ represents the vertical field-of-view of the sensor, and $f_{up}$ and $f_{down}$ are the pitch angles at the upward and downward directions, respectively.
-Points with the same $(u,v)$ are grouped into the same frustum region.
-The hyperparameter configuration of each module in our proposed FRNet framework is summarized as follows.
-  • Frustum Feature Encoder.
-The input of frustum feature encoder for each point is $(x,y,z,i,d,x-x_{m},y-y_{m},z-z_{m})$, where $(x_{m},y_{m},z_{m})$ is the mean coordinate in the corresponding frustum region.
-The features are fed into four MLPs that map the channel from $8$ to $64$, $128$, $256$, and $256$, respectively, with batch normalization and ReLU function on the first three layers.
-Then the point features are pooled into frustum features, which are fed into an MLP that maps the channel from $256$ to $16$ with a ReLU function.
-  • 2D Backbone.
-We adopt a variant of ResNet18 and ResNet34 used in CENet [8] as the backbone of Fast-FRNet and FRNet, respectively.
-The spatial shape of each stage is $[(H,W),(\frac{H}{2},\frac{W}{2}),(\frac{H}{4},\frac{W}{4}),(\frac{H}{8},\frac{W}{8})]$ and the output channel is $[128,128,128,128]$.
-The four features are interpolated to the shape of $(H,W)$ for feature concatenation, with two convolutional layers to reduce the channel dimension to $[256,128]$ progressively.
-  • Frustum-Point Fusion Module.
-Point features are updated hierarchically at each stage of the 2D backbone.
-Accordingly, the channel dimension of point features at each stage is $[128,128,128,128]$.
-We also concatenate point features at each stage and reduce the channel dimension to $[256,128]$ with two MLPs progressively.
-  • Optimization.
-For point-level supervision, we use the cross-entropy loss.
-For frustum-level supervision, we use cross-entropy loss, Lovasz-Softmax loss [6], and boundary loss [42], and the weights for each loss are $1.0$, $1.5$, and $1.0$, respectively.
-The frustum-level loss is also used at each stage of the 2D backbone to supervise the middle frustum features.
-The weights for point-level loss and frustum-level loss are $1.0$ and $1.0$, respectively.
-Algorithm 1 FrustumMix, NumPy-style
-# lidar_list: the whole dataset.
-# h: height of the 2D representation.
-# w: width of the 2D representation.
-# f_u: upward pitch angles of the field-of-view.
-# f_d: downward pitch angles of the field-of-view.
-# num_areas: List of split regions along inclination direction. We set [3, 4, 5, 6] here.
-# sample: function to sample the lidar points and labels from the whole dataset.
-# group: function to calculate frustum regions for points, refer to Eq. (1).
-def FrustumMix(xa, ya, xb, yb):
-    # group frustum regions
-    ua, va = group(xa, h, w, f_u, f_d)
-    ub, vb = group(xb, h, w, f_u, f_d)
-    out_points = []
-    out_labels = []
-    if np.random.rand() > 0.5:
-        # mix two scans along inclination direction
-        n = np.random.choice(num_areas, size=1)[0]
-        inc_list = np.linspace(0, h, n+1, dtype=np.int64)
-        for i in range(n):
-            start = inc_list[i]
-            end = inc_list[i+1]
-            if i % 2 == 0:
-                idx = (va>=start) & (va<end)
-                out_points.append(xa[idx])
-                out_labels.append(ya[idx])
-            else:
-                idx = (vb>=start) & (vb<end)
-                out_points.append(xb[idx])
-                out_labels.append(yb[idx])
-    else:
-        # mix two scans along azimuth direction
-        start = np.random.randint(0, w//2)
-        end = start + w//2
-        idxa = (ua<start) | (ua>=end)
-        idxb = (ub>=start) & (ub<end)
-        out_points = [xa[idxa], xb[idxb]]
-        out_labels = [ya[idxa], yb[idxb]]
-    out_points = np.concatenate(out_points)
-    out_labels = np.concatenate(out_labels)
-    return out_points, out_labels
-# sample two sans
-idx_a = np.random.randint(0, len(lidar_list))
-idx_b = np.random.randint(0, len(lidar_list))
-points_a, labels_a = sample(idx_a)
-points_b, labels_b = sample(idx_b)
-# mix the two scans
-points, labels = FrustumMix(points_a, labels_a, points_b, labels_b)
-Algorithm 2 Range-Interpolation, NumPy-style
-# h: height of the range image.
-# w: width of the range image.
-# f_u: upward pitch angles of the field-of-view.
-# f_d: downward pitch angles of the field-of-view.
-# points: lidar points with shape [m, c].
-# labels: semantic labels with shape: [m,].
-# ignore_label: ignored label that is not used to calculate loss.
-# laser_scan: function to convert LiDAR points to range image.
-def RangeInterpolation(p, l):
-    # convert to range images
-    # r_m means the valid mask for each pixel. 0 is invalid, while 1 is valid.
-    r_p, r_l, r_m = laser_scan(p, l, h, w, f_u, f_d)
-    out_points = [p]
-    out_labels = [l]
-    for y in range(h):
-        for x in range(w):
-            if r_m[y, x]:
-                continue
-            if (x-1>=0) and (x+1<w):
-                if r_m[y, x-1] and r_m[y, x+1]:
-                    # both right and left pixels are valid
-                    point = (r_p[y, x-1]+r_p[y, x+1]) / 2
-                    if r_l[y, x-1] == r_l[y, x+1]:
-                        label = r_l[y, x-1]
-                    else:
-                        label = ignore_label
-                    out_points.append(point)
-                    out_labels.append(label)
-                    r_p[y, x] = point
-                    r_l[y, x] = label
-                    r_m[y, x] = 1
-    out_points = np.concatenate(out_points)
-    out_labels = np.concatenate(out_labels)
-    return out_points, out_labels
-points, labels = RangeInterpolation(points, labels)
-A.3 Training Configurations
-The common augmentation methods, including random dropping, random flipping, global rotation, global scale, global translation, and instance copy, are used in FRNet.
-Additionally, we also use the following proposed augmentation methods to further improve the performance.
-  • FrustumMix.
-Based on the calculated frustum region $(u,v)$, the two scans are randomly mixed along inclination and azimuth directions.
-For inclination direction, the frustum regions are randomly split into several subsets sampled from $[3,4,5,6]$, while for azimuth direction, the frustum regions are split into $2$ subsets.
-  • Range-Interpolation.
-Range-Interpolation aims to synthesize some possible points based on range images, which helps to fill empty pixels in the 2D representations.
-We first scan all the pixels in the range image.
-For each empty pixel, the interpolated points are calculated only when its right and left pixels both have projected points.
-During the validation and testing phase, only the Range-Interpolation is kept to fill the empty pixels in 2D representations.
-To make the final results correspond to the original points, we set the first $N$ labels as the predicted labels as the last multiple points are interpolated points.
-Moreover, we apply the test time augmentation to improve the performance.
-Specifically, the input points are flipped along different directions, with randomly rotated, scaled, and jittered.
-The total number of voted point cloud scenes is $4$.
-For each scene, FRNet generates confidence scores for each class.
-We add the confidence scores of the $4$ scenes and then predict the final semantic labels.
-Appendix B Additional Experimental Results
-In this section, we provide the complete experimental results and qualitative visual examples among the four popular LiDAR segmentation benchmarks used in this work.
+VI Additional Experimental Results
+In this section, we provide the complete experimental results among the four popular LiDAR segmentation benchmarks used in this work.
 (a) Speed vs. Accuracy
@@
 corruption error (the lower the better).
-B.1 Scalability and Robustness
+VI-A Scalability and Robustness
 In this section, we provide additional analysis of existing LiDAR segmentation approaches regarding their trade-offs in-between the inference speed (FPS), segmentation accuracy (mIoU), and out-of-training-distribution robustness (mCE).
 As shown in Fig.
-7, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed.
+8, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed.
 Those more accurate LiDAR segmentation models often contain larger parameter sets and are less efficient during the inference stage.
@@
 Our models achieve competitive segmentation results and superior robustness over the voxel-based and point-voxel fusion counterparts, while still maintaining high efficiency for real-time LiDAR segmentation.
-B.2 Per-Class Results
-In this section, we supplement the complete (class-wise) segmentation results of the baselines, prior works, and our proposed FRNet and Fast-FRNet.
-B.2.1 SemanticKITTI
-Tab.
-5 shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [5].
-We compare FRNet with different LiDAR representations, including point-view, range-view, voxel-view, and multi-view methods.
-Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods.
-Compared with SphereFormer [27] and UniSeg [32], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy.
-More results about efficiency and accuracy can be found in the main body.
-It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
-B.2.2 nuScenes
-Tab.
-6 and Tab.
-7 show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [13], respectively.
-The results demonstrate the great advantages of Fast-FRNet and FRNet.
-For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters.
-Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
-B.2.3 ScribbleKITTI
-To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [49], including RangeNet++ [35], SalsaNext [12], FIRNet [64], CENet [8], and RangeViT [4].
-All the models are trained with their official settings.
-As shown in Tab.
-8, FRNet achieves state-of-the-art performance.
-However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
-B.2.4 SemanticPOSS
-We also re-implement some popular voxel-view methods, including SPVCNN [47], Cylinder3D [65], and MinkNet [10], on SemanticPOSS [37].
-All the experiments are conducted based on MMDetection3D framework [11].
-Tab.
-9 summarizes the class-wise IoU scores among these methods and some other popular range-view methods.
-Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods.
-For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
-B.3 Qualitative Results
-In this section, we provide additional qualitative examples to further validate the effectiveness and superiority of the proposed FRNet framework.
-B.3.1 Visual Comparisons
-We supplement visual comparison examples of the proposed FRNet with prior works [4] and show the results in Fig.
-9.
-Compared with range-view methods, FRNet can handle better for objects with complex structures, such as vegetation and terrain.
-In addition, the objects have much clearer boundaries.
-B.3.2 Failure Cases
-Although FRNet achieves great improvements among popular range-view methods, there are still some limitations that need to be considered.
-As shown in Fig.
-8, FRNet tends to misclassify objects with similar structures to a large extent, i.e., bicyclist and motorcyclist, sidewalk, and other-ground.
-Spatial-temporal methods will help to alleviate this problem and yield better performance.
-B.3.3 Video Demos
-In addition to the figures shown in the main body and supplementary material, we also provide some video demos to support the superiority of FRNet.
-Specifically, we attached three video clips in the supplementary file, named demo1.mp4, demo2.mp4, and demo3.mp4.
-The first two demos include the error maps of the proposed FRNet and the prior art RangeViT [4].
-The third demo contains the error maps and semantic predictions of FRNet.
-Each video demo shows LiDAR semantic segmentation sequences with hundreds of frames.
-All three demos validate again the effectiveness and superiority of FRNet in tackling the challenging LiDAR semantic segmentation task.
-We hope these video demos can provide a more comprehensive qualitative assessment of our proposed approach.
-Table 5: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [5] leaderboard.
+TABLE VI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [28] leaderboard.
 All IoU scores are given in percentage (%).
@@
 Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign PointNet [78] $14.6$ $46.3$ $1.3$ $0.3$ $0.1$ $0.8$ $0.2$ $0.2$ $0.0$ $61.6$ $15.8$ $35.7$ $1.4$ $41.4$ $12.9$ $31.0$ $4.6$ $17.6$ $2.4$ $3.7$ PointNet++ [60] $20.1$ $53.7$ $1.9$ $0.2$ $0.9$ $0.2$ $0.9$ $1.0$ $0.0$ $72.0$ $18.7$ $41.8$ $5.6$ $62.3$ $16.9$ $46.5$ $13.8$ $30.0$ $6.0$ $8.9$ SqSeg [29] $30.8$ $68.3$ $18.1$ $5.1$ $4.1$ $4.8$ $16.5$ $17.3$ $1.2$ $84.9$ $28.4$ $54.7$ $4.6$ $61.5$ $29.2$ $59.6$ $25.5$ $54.7$ $11.2$ $36.3$ SqSegV2 [50] $39.6$ $82.7$ $21.0$ $22.6$ $14.5$ $15.9$ $20.2$ $24.3$ $2.9$ $88.5$ $42.4$ $65.5$ $18.7$ $73.8$ $41.0$ $68.5$ $36.9$ $58.9$ $12.9$ $41.0$ RandLA-Net [8] $50.3$ $94.0$ $19.8$ $21.4$ $42.7$ $38.7$ $47.5$ $48.8$ $4.6$ $90.4$ $56.9$ $67.9$ $15.5$ $81.1$ $49.7$ $78.3$ $60.3$ $59.0$ $44.2$ $38.1$ RangeNet++ [30] $52.2$ $91.4$ $25.7$ $34.4$ $25.7$ $23.0$ $38.3$ $38.8$ $4.8$ $91.8$ $65.0$ $75.2$ $27.8$ $87.4$ $58.6$ $80.5$ $55.1$ $64.6$ $47.9$ $55.9$ PolarNet [33] $54.3$ $93.8$ $40.3$ $30.1$ $22.9$ $28.5$ $43.2$ $40.2$ $5.6$ $90.8$ $61.7$ $74.4$ $21.7$ $90.0$ $61.3$ $84.0$ $65.5$ $67.8$ $51.8$ $57.5$ MPF [22] $55.5$ $93.4$ $30.2$ $38.3$ $26.1$ $28.5$ $48.1$ $46.1$ $18.1$ $90.6$ $62.3$ $74.5$ $30.6$ $88.5$ $59.7$ $83.5$ $59.7$ $69.2$ $49.7$ $58.1$ 3D-MiniNet [79] $55.8$ $90.5$ $42.3$ $42.1$ $28.5$ $29.4$ $47.8$ $44.1$ $14.5$ $91.6$ $64.2$ $74.5$ $25.4$ $89.4$ $60.8$ $82.8$ $60.8$ $66.7$ $48.0$ $56.6$ SqSegV3 [52] $55.9$ $92.5$ $38.7$ $36.5$ $29.6$ $33.0$ $45.6$ $46.2$ $20.1$ $91.7$ $63.4$ $74.8$ $26.4$ $89.0$ $59.4$ $82.0$ $58.7$ $65.4$ $49.6$ $58.9$ KPConv [7] $58.8$ $96.0$ $32.0$ $42.5$ $33.4$ $44.3$ $61.5$ $61.6$ $11.8$ $88.8$ $61.3$ $72.7$ $31.6$ $95.0$ $64.2$ $84.8$ $69.2$ $69.1$ $56.4$ $47.4$ SalsaNext [57] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $19.4$ $91.7$ $63.7$ $75.8$ $29.1$ $90.2$ $64.2$ $81.8$ $63.6$ $66.5$ $54.3$ $62.1$ FIDNet [55] $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $23.7$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$ FusionNet [80] $61.3$ $95.3$ $47.5$ $37.7$ $41.8$ $34.5$ $59.5$ $56.8$ $11.9$ $91.8$ $68.8$ $77.1$ $30.8$ $92.5$ $69.4$ $84.5$ $69.8$ $68.5$ $60.4$ $66.5$ PCSCNet [81] $62.7$ $95.7$ $48.8$ $46.2$ $36.4$ $40.6$ $55.5$ $68.4$ $55.9$ $89.1$ $60.2$ $72.4$ $23.7$ $89.3$ $64.3$ $84.2$ $68.2$ $68.1$ $60.5$ $63.9$ KPRNet [56] $63.1$ $95.5$ $54.1$ $47.9$ $23.6$ $42.6$ $65.9$ $65.0$ $16.5$ $93.2$ $73.9$ $80.6$ $30.2$ $91.7$ $68.4$ $85.7$ $69.8$ $71.2$ $58.7$ $64.1$ TornadoNet [82] $63.1$ $94.2$ $55.7$ $48.1$ $40.0$ $38.2$ $63.6$ $60.1$ $34.9$ $89.7$ $66.3$ $74.5$ $28.7$ $91.3$ $65.6$ $85.6$ $67.0$ $71.5$ $58.0$ $65.9$ LiteHDSeg [83] $63.8$ $92.3$ $40.0$ $55.4$ $37.7$ $39.6$ $59.2$ $71.6$ $54.3$ $93.0$ $68.2$ $78.3$ $29.3$ $91.5$ $65.0$ $78.2$ $65.8$ $65.1$ $59.5$ $67.7$ RangeViT [31] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $38.1$ $93.1$ $70.2$ $80.0$ $32.5$ $92.0$ $69.0$ $85.3$ $70.6$ $71.2$ $60.8$ $64.7$ CENet [58] $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $43.5$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $70.0$ $61.5$ $67.6$ SVASeg [18] $65.2$ $96.7$ $56.4$ $57.0$ $49.1$ $56.3$ $70.6$ $67.0$ $15.4$ $92.3$ $65.9$ $76.5$ $23.6$ $91.4$ $66.1$ $85.2$ $72.9$ $67.8$ $63.9$ $65.2$ AMVNet [20] $65.3$ $96.2$ $59.9$ $54.2$ $48.8$ $45.7$ $71.0$ $65.7$ $11.0$ $90.1$ $71.0$ $75.8$ $32.4$ $92.4$ $69.1$ $85.6$ $71.7$ $69.6$ $62.7$ $67.2$ GFNet [84] $65.4$ $96.0$ $53.2$ $48.3$ $31.7$ $47.3$ $62.8$ $57.3$ $44.7$ $\mathbf{93.6}$ $72.5$ $\mathbf{80.8}$ $31.2$ $\mathbf{94.0}$ $\mathbf{73.9}$ $85.2$ $71.1$ $69.3$ $61.8$ $68.0$ JS3C-Net [85] $66.0$ $95.8$ $59.3$ $52.9$ $54.3$ $46.0$ $69.5$ $65.4$ $39.9$ $88.9$ $61.9$ $72.1$ $31.9$ $92.5$ $70.8$ $84.5$ $69.8$ $67.9$ $60.7$ $68.7$ MaskRange [86] $66.1$ $94.2$ $56.0$ $55.7$ $59.2$ $52.4$ $67.6$ $64.8$ $31.8$ $91.7$ $70.7$ $77.1$ $29.5$ $90.6$ $65.2$ $84.6$ $68.5$ $69.2$ $60.2$ $66.6$ SPVNAS [15] $66.4$ $97.3$ $51.5$ $50.8$ $59.8$ $58.8$ $65.7$ $65.2$ $43.7$ $90.2$ $67.6$ $75.2$ $16.9$ $91.3$ $65.9$ $86.1$ $73.4$ $71.0$ $64.2$ $66.9$ MSSNet [87] $66.7$ $96.8$ $52.2$ $48.5$ $54.4$ $56.3$ $67.0$ $70.9$ $49.3$ $90.1$ $65.5$ $74.9$ $30.2$ $90.5$ $64.9$ $84.9$ $72.7$ $69.2$ $63.2$ $65.1$ Cylinder3D [16] $68.9$ $97.1$ $67.6$ $63.8$ $50.8$ $58.5$ $73.7$ $69.2$ $48.0$ $92.2$ $65.0$ $77.0$ $32.3$ $90.7$ $66.5$ $85.6$ $72.5$ $69.8$ $62.4$ $66.2$ AF2S3Net [88] $69.7$ $94.5$ $65.4$ $\mathbf{86.8}$ $39.2$ $41.1$ $\mathbf{80.7}$ $80.4$ $74.3$ $91.3$ $68.8$ $72.5$ $\mathbf{53.5}$ $87.9$ $63.2$ $70.2$ $68.5$ $53.7$ $61.5$ $71.0$ RPVNet [21] $70.3$ $97.6$ $68.4$ $68.7$ $44.2$ $61.1$ $75.9$ $74.4$ $73.4$ $93.4$ $70.3$ $80.7$ $33.3$ $93.5$ $72.1$ $86.5$ $75.1$ $71.7$ $64.8$ $61.4$ SDSeg3D [89] $70.4$ $97.4$ $58.7$ $54.2$ $54.9$ $65.2$ $70.2$ $74.4$ $52.2$ $90.9$ $69.4$ $76.7$ $41.9$ $93.2$ $71.1$ $86.1$ $74.3$ $71.1$ $65.4$ $70.6$ GASN [90] $70.7$ $96.9$ $65.8$ $58.0$ $59.3$ $61.0$ $80.4$ $\mathbf{82.7}$ $46.3$ $89.8$ $66.2$ $74.6$ $30.1$ $92.3$ $69.6$ $87.3$ $73.0$ $72.5$ $66.1$ $71.6$ PVKD [65] $71.2$ $97.0$ $67.9$ $69.3$ $53.5$ $60.2$ $75.1$ $73.5$ $50.5$ $91.8$ $70.9$ $77.5$ $41.0$ $92.4$ $69.4$ $86.5$ $73.8$ $71.9$ $64.9$ $65.8$ Fast-FRNet $72.5$ $97.1$ $66.0$ $73.1$ $59.3$ $65.7$ $76.0$ $78.9$ $54.5$ $91.8$ $72.6$ $77.6$ $42.0$ $92.4$ $70.6$ $86.5$ $71.9$ $72.5$ $63.1$ $66.7$ 2DPASS [67] $72.9$ $97.0$ $63.6$ $63.4$ $61.1$ $61.5$ $77.9$ $81.3$ $74.1$ $89.7$ $67.4$ $74.7$ $40.0$ $93.5$ $72.9$ $86.2$ $73.9$ $71.0$ $65.0$ $70.4$ RangeFormer [32] $73.3$ $96.7$ $69.4$ $73.7$ $59.9$ $66.2$ $78.1$ $75.9$ $58.1$ $92.4$ $73.0$ $78.8$ $42.4$ $92.3$ $70.1$ $86.6$ $73.3$ $72.8$ $66.4$ $66.6$ FRNet $73.3$ $97.3$ $67.9$ $74.6$ $59.4$ $66.3$ $78.1$ $79.2$ $57.3$ $92.1$ $73.0$ $78.1$ $41.8$ $92.7$ $71.0$ $86.7$ $73.2$ $72.5$ $64.7$ $67.3$ SphereFormer [68] $74.8$ $97.5$ $70.1$ $70.5$ $59.6$ $67.7$ $79.0$ $80.4$ $\mathbf{75.3}$ $91.8$ $69.7$ $78.2$ $41.3$ $93.8$ $72.8$ $86.7$ $75.1$ $72.4$ $66.8$ $\mathbf{72.9}$ UniSeg [23] $\mathbf{75.2}$ $\mathbf{97.9}$ $\mathbf{71.9}$ $75.2$ $\mathbf{63.6}$ $\mathbf{74.1}$ $78.9$ $74.8$ $60.6$ $92.6$ $\mathbf{74.0}$ $79.5$ $46.1$ $93.4$ $72.7$ $\mathbf{87.5}$ $\mathbf{76.3}$ $\mathbf{73.1}$ $\mathbf{68.3}$ $68.5$
-Table 6: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [13].
+TABLE VII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [43].
 All IoU scores are given in percentage (%).
 The best and second best scores for each class are highlighted in bold and underline.
-Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [9] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [35] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [61] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [38] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [12] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [63] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [4] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [65] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [31] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [54] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [24] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [27] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
-Table 7: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [13].
+Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [88] [’21] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [30] [’19] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [33] [’20] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [81] [’22] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [57] [’20] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [18] [’22] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [31] [’23] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [16] [’21] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [20] [’20] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [21] [’21] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [32] [’23] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [68] [’23] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
+TABLE VIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [43].
 All IoU scores are given in percentage (%).
 The best and second best scores for each class are highlighted in bold and underline.
-Method mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [61] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [55] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [66] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [65] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [31] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [47] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [9] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [15] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [24] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [58] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [56] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [57] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [27] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [32] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
-Table 8: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [49] leaderboard.
+Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [33] [’20] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [85] [’21] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [91] [’21] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [16] [’21] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [20] [’20] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [15] [’20] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [88] [’21] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [92] [’21] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [32] [’23] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [90] [’22] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [67] [’22] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [93] [’23] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [68] [’23] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [23] [’23] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
+TABLE IX: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [63] leaderboard.
 All IoU scores are given in percentage (%).
@@
 Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign RangeNet++ [30] $44.6$ $84.6$ $24.6$ $38.9$ $13.7$ $12.9$ $29.0$ $51.4$ $0.0$ $86.0$ $35.2$ $72.4$ $4.5$ $80.2$ $41.6$ $80.1$ $51.2$ $66.9$ $43.9$ $34.5$ SalsaNext [57] $50.3$ $85.6$ $33.8$ $42.7$ $8.4$ $20.8$ $64.3$ $71.5$ $0.2$ $85.6$ $31.6$ $72.0$ $1.1$ $84.0$ $39.9$ $81.0$ $62.0$ $63.6$ $60.5$ $46.3$ RangeViT [31] $53.6$ $85.6$ $31.6$ $50.1$ $40.3$ $36.3$ $57.6$ $68.7$ $0.0$ $86.1$ $32.6$ $75.2$ $0.3$ $87.9$ $49.3$ $83.6$ $62.8$ $67.5$ $59.6$ $43.7$ FIDNet [55] $54.1$ $85.6$ $36.7$ $48.7$ $60.8$ $38.4$ $63.3$ $68.2$ $0.0$ $84.1$ $25.9$ $71.2$ $0.4$ $85.6$ $41.3$ $81.7$ $64.1$ $62.7$ $61.5$ $48.0$ MinkNet [17] $55.0$ $88.1$ $13.2$ $55.1$ $72.3$ $36.9$ $61.3$ $77.1$ $0.0$ $83.4$ $32.7$ $71.0$ $0.3$ $90.0$ $50.0$ $84.1$ $66.6$ $65.8$ $61.6$ $35.2$ CENet [58] $55.7$ $86.1$ $39.4$ $53.2$ $61.0$ $46.1$ $69.2$ $72.2$ $0.0$ $85.7$ $28.7$ $72.6$ $1.1$ $85.8$ $43.1$ $81.8$ $64.2$ $63.8$ $59.6$ $45.0$ SPVCNN [15] $56.9$ $88.6$ $25.7$ $55.9$ $67.4$ $48.8$ $65.0$ $78.2$ $0.0$ $82.6$ $30.4$ $70.1$ $0.3$ $\mathbf{90.5}$ $49.6$ $84.4$ $\mathbf{67.6}$ $66.1$ $61.6$ $\mathbf{48.7}$ Cylinder3D [16] $57.0$ $88.5$ $39.9$ $58.0$ $58.4$ $48.1$ $68.6$ $77.0$ $\mathbf{0.5}$ $84.4$ $30.4$ $72.2$ $2.5$ $89.4$ $48.4$ $81.9$ $64.6$ $59.8$ $61.2$ $\mathbf{48.7}$ Fast-FRNet $62.4$ $90.8$ $41.0$ $\mathbf{66.8}$ $81.7$ $\mathbf{64.0}$ $70.5$ $84.2$ $0.0$ $91.3$ $35.1$ $78.3$ $0.0$ $89.4$ $64.4$ $85.0$ $65.5$ $69.5$ $60.5$ $47.7$ RangeFormer [32] $63.0$ $\mathbf{92.6}$ $\mathbf{51.6}$ $65.7$ $74.4$ $49.6$ $71.6$ $82.1$ $0.0$ $\mathbf{94.8}$ $\mathbf{44.4}$ $\mathbf{80.6}$ $\mathbf{11.4}$ $85.6$ $56.9$ $\mathbf{87.2}$ $64.1$ $\mathbf{77.0}$ $\mathbf{62.7}$ $45.1$ FRNet $\mathbf{63.1}$ $90.5$ $42.3$ $\mathbf{66.8}$ $\mathbf{82.4}$ $63.0$ $\mathbf{73.4}$ $\mathbf{86.2}$ $0.0$ $92.1$ $35.1$ $79.1$ $0.3$ $90.4$ $\mathbf{64.7}$ $85.4$ $65.8$ $70.7$ $61.8$ $48.6$
-Appendix C Broader Impact
-In this section, we elaborate on the positive societal influence and potential limitations of the proposed FRNet.
-C.1 Positive Societal Impacts
-Conducting efficient and accurate range view LiDAR semantic segmentation for autonomous driving has several positive societal impacts, particularly in terms of enhancing road safety, improving transportation efficiency, and contributing to broader technological advancements.
-  • Enhanced Road Safety.
-One of the most significant benefits is the improvement in road safety.
-Autonomous vehicles equipped with advanced LiDAR semantic segmentation can accurately detect and classify objects in their environment, such as other vehicles, pedestrians, cyclists, and road obstacles.
-This precise perception capability allows for safer navigation and decision-making, potentially reducing accidents caused by human error.
-  • Reduced Traffic Congestion.
-Autonomous vehicles with advanced perception systems can optimize driving patterns, leading to smoother traffic flow.
-This can significantly reduce traffic congestion, especially in urban areas, thereby saving time for commuters and reducing stress associated with driving.
-  • Emergency Response and Healthcare.
-In emergency situations, autonomous vehicles can be used to quickly and safely transport patients or deliver medical supplies.
-The precision of LiDAR segmentation ensures that these vehicles can navigate through complex environments effectively.
-  • Advancements in Smart City Infrastructure.
-The integration of autonomous vehicles with smart city initiatives can lead to more efficient urban planning and infrastructure development.
-LiDAR data can be used not just for navigation but also for gathering urban data, which can inform city planning and maintenance.
-C.2 Potential Limitations
-Although the proposed FRNet and Fast-FRNet are capable of providing a better trade-off between LiDAR segmentation network efficiency and accuracy, there are several limitations within the current framework.
-First, frustum-level supervision counts the high-frequency semantic labels in the frustum region, which will cover the objects with few points.
-Such supervision weakens the performance of FRNet on tiny objects.
-Second, FRNet cannot handle objects with similar structures well.
-Some objects share a similar appearance with different semantic attributes, which limits the ability of FRNet to distinguish them in a 2D manner.
-Appendix D Public Resources Used
-In this section, we acknowledge the use of public resources, during the course of this work.
-D.1 Public Codebase Used
-We acknowledge the use of the following public codebase, during the course of this work:
-D.2 Public Datasets Used
-We acknowledge the use of the following public datasets, during the course of this work:
-D.3 Public Implementations Used
-We acknowledge the use of the following public implementations, during the course of this work:
-Table 9: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [37] leaderboard.
+TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [64] leaderboard.
 All IoU scores are given in percentage (%).
@@
 FRNet $\mathbf{53.5}$ $\underline{77.7}$ $28.7$ $81.2$ $28.9$ $74.3$ $\mathbf{30.4}$ $34.5$ $29.8$ $81.1$ $\mathbf{45.6}$ $47.9$ $56.4$ $79.2$
-Figure 8: Failure case examples in FRNet.
-To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively.
-Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
-Figure 9: Qualitative comparisons with state-of-the-art LiDAR segmentation methods [4] on the val set of nuScenes [13].
-To highlight the differences compared with groundtruth, the correct and incorrect predictions are painted in gray and violet, respectively.
-Each point cloud scene is visualized under both the bird’s-eye-view and range-view.
+VI-B Per-Class Results
+In this section, we supplement the complete (class-wise) segmentation results of the baselines, prior works, and our proposed FRNet and Fast-FRNet.
+VI-B1 SemanticKITTI
+Tab.
+VI shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [28].
+We compare FRNet with different LiDAR representations, including point-view, range-view, sparse-voxel-view, and multi-view methods.
+Results demonstrate that FRNet achieves appealing performance among state-of-the-art methods.
+Compared with SphereFormer [68] and UniSeg [23], although FRNet obtains sub-par performance, it maintains satisfactory efficiency for real-time processing, achieving the balance between efficiency and accuracy.
+More results about efficiency and accuracy can be found in the main body.
+It is worth noting that Fast-FRNet achieves the fastest speed of inference with only a little sacrifice of performance.
+VI-B2 nuScenes
+Tab.
+VII and Tab.
+VIII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [43], respectively.
+The results demonstrate the great advantages of Fast-FRNet and FRNet.
+For the much sparser LiDAR points, the performance of Fast-FRNet is very close to FRNet with fewer parameters.
+Compared with popular range-view methods, FRNet achieves great improvement in some dynamic instances, such as car, bicycle, motorcycle, etc.
+VI-B3 ScribbleKITTI
+To prove the great advantages of FRNet among range-view methods, we also reimplement the popular LiDAR segmentation methods on ScribbleKITTI [63], including RangeNet++ [30], SalsaNext [57], FIRNet [55], CENet [58], and RangeViT [31].
+All the models are trained with their official settings.
+As shown in Tab.
+IX, FRNet achieves state-of-the-art performance.
+However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
+VI-B4 SemanticPOSS
+We also reimplement some popular sparse-voxel-view methods, including SPVCNN [15], Cylinder3D [16], and MinkNet [17], on SemanticPOSS [64].
+All the experiments are conducted based on MMDetection3D framework [72].
+Tab.
+X summarizes the class-wise IoU scores among these methods and some other popular range-view methods.
+Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods.
+For some tiny objects, such as traffic sign and cone/stone, FRNet achieves superior improvements.
+VII Broader Impact
+In this section, we elaborate on the positive societal influence and potential limitations of the proposed FRNet.
+VII-A Positive Societal Impacts
+Conducting efficient and accurate range view LiDAR semantic segmentation for autonomous driving has several positive societal impacts, particularly in terms of enhancing road safety, improving transportation efficiency, and contributing to broader technological advancements.
+  • Enhanced Road Safety.
+One of the most significant benefits is the improvement in road safety.
+Autonomous vehicles equipped with advanced LiDAR semantic segmentation can accurately detect and classify objects in their environment, such as other vehicles, pedestrians, cyclists, and road obstacles.
+This precise perception capability allows for safer navigation and decision-making, potentially reducing accidents caused by human error.
+  • Reduced Traffic Congestion.
+Autonomous vehicles with advanced perception systems can optimize driving patterns, leading to smoother traffic flow.
+This can significantly reduce traffic congestion, especially in urban areas, thereby saving time for commuters and reducing stress associated with driving.
+  • Emergency Response and Healthcare.
+In emergency situations, autonomous vehicles can be used to quickly and safely transport patients or deliver medical supplies.
+The precision of LiDAR segmentation ensures that these vehicles can navigate through complex environments effectively.
+  • Advancements in Smart City Infrastructure.
+The integration of autonomous vehicles with smart city initiatives can lead to more efficient urban planning and infrastructure development.
+LiDAR data can be used not just for navigation but also for gathering urban data, which can inform city planning and maintenance.
+VII-B Potential Limitations
+Although the proposed FRNet and Fast-FRNet are capable of providing a better trade-off between LiDAR segmentation network efficiency and accuracy, there are several limitations within the current framework.
+First, frustum-level supervision counts the high-frequency semantic labels in the frustum region, which will cover the objects with few points.
+Such supervision weakens the performance of FRNet on tiny objects.
+Second, FRNet cannot handle objects with similar structures well.
+Some objects share a similar appearance with different semantic attributes, which limits the ability of FRNet to distinguish them in a 2D manner.
+VIII Public Resources Used
+In this section, we acknowledge the use of public resources, during the course of this work.
+VIII-A Public Codebase Used
+We acknowledge the use of the following public codebase, during the course of this work:
+VIII-B Public Datasets Used
+We acknowledge the use of the following public datasets, during the course of this work:
+VIII-C Public Implementations Used
+We acknowledge the use of the following public implementations, during the course of this work:

### v3

@@
 Abstract
-LiDAR segmentation has become a crucial component in advanced autonomous driving systems.
+LiDAR segmentation has become a crucial component of advanced autonomous driving systems.
 Recent range-view LiDAR segmentation approaches show promise for real-time processing.
@@
 In this work, we propose FRNet, a simple yet powerful method aimed at restoring the contextual information of range image pixels using corresponding frustum LiDAR points.
-Firstly, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is crucial for point-level predictions.
-Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information via the frustum features.
-Finally, a head fusion module is used to fuse features at different levels for final semantic prediction.
+First, a frustum feature encoder module is used to extract per-point features within the frustum region, which preserves scene consistency and is critical for point-level predictions.
+Next, a frustum-point fusion module is introduced to update per-point features hierarchically, enabling each point to extract more surrounding information through the frustum features.
+Finally, a head fusion module is used to fuse features at different levels for final semantic predictions.
 Extensive experiments conducted on four popular LiDAR segmentation benchmarks under various task setups demonstrate the superiority of FRNet.
@@
 I Introduction
-LiDAR segmentation, an crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4].
+LiDAR segmentation, a crucial and indispensable component in modern autonomous driving, robotics, and other safety-critical applications, has witnessed substantial advancements recently [1, 2, 3, 4].
 The principal challenge lies in devising a scalable LiDAR segmentation system that strikes a delicate balance between efficiency and accuracy, especially in resource-constrained operational scenarios [5, 6].
-Existing LiDAR segmentation methods adopt various data representation perspectives, each with its own set of trade-offs in terms of accuracy and computational efficiency.
-The point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve raw spatial information, but entail computationally expensive neighborhood searches to construct local structures.
-This demand results in considerable computation overhead and limits their capability to deal with large-scale scene point clouds [13, 14, 10].
-Sparse-voxel-based methods [15, 16, 17, 18] transform scattered LiDAR points into regular voxel representations and leverage popular sparse convolutions to extract features.
+Here, “scalable” refers to the system’s ability to maintain high performance while managing increased data loads without compromising real-time processing capabilities.
+Existing LiDAR segmentation methods adopt various data representation perspectives, each with inherent trade-offs in terms of accuracy and computational efficiency.
+Point-based methods [7, 8, 9, 10, 11, 12] manipulate the original point cloud to preserve spatial granularity but entail computationally expensive neighborhood searches to construct local structures.
+This demand results in significant computational overhead and limits their capability to deal with large-scale point clouds [13, 14, 10].
+Sparse-voxel-based methods [15, 16, 17, 18, 19] transform scattered LiDAR points into regular voxel grids and leverage popular sparse convolutions to extract features.
 However, this process also demands heavy computations, especially for applications that require high voxel resolutions [16, 15, 20].
@@
 While this strategy can improve performance, it aggregates the computation required for all representations.
-Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [27, 23].
+Albeit these representation methods achieved satisfactory performance, their computation overhead hinders them from real-time online applications in real life [28, 24, 29, 30, 31].
 Figure 1: A study on the scalability of state-of-the-art LiDAR segmentation models on the SemanticKITTI [32] leaderboard.
@@
 In essence, although pseudo-image methods excel in terms of real-time processing capabilities, their reduced dimensional perspective can overlook crucial 3D details, leading to degraded scalability.
-Observing the above issues, in this work, we propose a scalable Frustum-Range Network (FRNet) for LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy.
+Observing the above issues, in this work, we propose a Frustum-Range Network (FRNet) for scalable LiDAR segmentation, which incorporates points into the range image, achieving a superior balance between efficiency and accuracy.
 FRNet consists of three main components.
@@
 Data augmentation plays an important role in training robust and generalizable models.
-Inspired by the success of mixing strategies [44, 45, 46], we introduce a novel mixing method dubbed FrustumMix, which operates based on frustum units.
+Inspired by the success of mixing strategies [48, 49, 50], we introduce a fine-grained mixing method dubbed FrustumMix, which operates based on frustum units.
 Specifically, we randomly divide point cloud scenes into several frustum regions along inclination and azimuth directions and swap the corresponding regions from different scenes, with the neighborhood of each frustum region coming from the other scene.
@@
 RangeInterpolation is able to assist in creating a more compact 2D representation with fewer empty pixels, enabling convolutions to learn more semantically coherent information.
-In summary, the key contributions of this work can be listed as follows:
+The key contributions of this work are summarized as:
   • We introduce a novel Frustum-Range representation for scalable LiDAR segmentation.
 In our framework, the point-level geometric information is integrated into the pseudo-image representation, which not only retains the complete geometric structure of the point cloud, but also takes advantage of the efficiency of the 2D network.
-  • We propose two novel 3D data augmentation techniques to assist in training robust and generalizable models.
+  • We propose two 3D data augmentation techniques to assist in training robust and generalizable models.
 FrustumMix generates more complex scenes by mixing two different scans, while RangeInterpolation reconstructs the semantic surfaces based on surrounding range information.
@@
 To address this problem, WaffleIron [11] proposes to mix point and image features with residual connections.
-Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predicting over the entire point cloud.
-In this work, we propose to integrate point and image features efficiently, preserving both 3D geometric information from raw point features and efficiency from 2D convolutions, without the need for any pre- or post-processing techniques, which yields superior scalability for real-time LiDAR segmentation.
+Although achieving promising performance, it requires downsampling points based on voxel grid and space neighborhood search for point feature embedding, which hinders it from real-time processing and end-to-end predictions over the entire point cloud.
+In this work, we propose an efficient integration of point and image features that preserves both the 3D geometric information from raw point features and the efficiency of 2D convolutions, eliminating the need for pre- or post-processing techniques and yielding superior scalability for real-time LiDAR segmentation.
 II-B LiDAR Data Augmentation
-Recent works have explored various data augmentation techniques on point clouds [53, 54], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes.
-Mix3D [46] creates new training samples by combining two scenes out of context, which requires a larger memory cost.
-PolarMix [45] proposes to mix two LiDAR scenes based on azimuth angles and cropping instances from one scene to another, while LaserMix [44] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning.
+Recent works have explored various data augmentation techniques for point clouds [57, 58], which have shown promise in improving performance for indoor scene understanding but exhibit limited generalization to outdoor scenes.
+Mix3D [50] creates new training samples by combining two out-of-context scenes, which incurs higher memory costs.
+PolarMix [49] proposes to mix two LiDAR scenes based on azimuth angles and crop instances from one scene to another, while LaserMix [48] suggests exchanging labeled scans with unlabeled scans along the inclination direction for efficient semi-supervised learning.
 RangeFormer [36] presents a lightweight augmentation method that operates only on projected points in range images, disregarding occluded points.
@@
 RangeFormer [36] divides full point clouds into sub-clouds and infers each subset in a supervised manner.
-However, these methods heavily rely on 2D results and fail to extract point features over entire points in a learnable manner.
+However, these methods heavily rely on 2D results and fail to extract point features over the entire point cloud in a supervised manner.
 In contrast, KPRNet [60] and RangeViT [35] incorporate a KPConv [7] block to extract point features in an end-to-end manner.
@@
 LiDAR segmentation aims to assign semantic labels to each point in the point cloud [62].
-A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm of PointNet++ [60].
+A common approach is to operate directly on the LiDAR points without any pre-processing, following the paradigm established by PointNet++ [63].
 However, point-based operations often require extensive computations and can yield sub-optimal performance due to the sparsity of outdoor scenes.
 To address this problem, many recent works have proposed projecting LiDAR points onto a regular intermediate representation, known as range images, based on spherical coordinates to enhance efficiency.
-Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
-$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})f^{-1}H]}~{},$ (1)
-where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, and $\phi_{\text{up}}$ are $\phi_{\text{down}}$ are the inclination angles at the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image.
+Specifically, given a point cloud $\mathcal{P}=\{\mathbf{p}_{i}\}_{i=1}^{N}$ with $N$ points, where $\mathbf{p}_{i}\in\mathbb{R}^{3+L}$ includes the coordinate $(x_{i},y_{i},z_{i})$ and additional $L$-dimensional features (intensity, elongation, etc.), we map the point to its corresponding 2D position $(u_{i},v_{i})$ in the range image via the following transformation:
+$\binom{u_{i}}{v_{i}}=\binom{\frac{1}{2}[1-\arctan(y_{i},x_{i})\pi^{-1}]W}{[1-(\arcsin(z_{i}d_{i}^{-1})+\phi_{\text{down}})\phi^{-1}]H}~{},$ (1)
+where $d_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}+z_{i}^{2}}$ is the depth of the point; $\phi=|\phi_{\text{up}}|+|\phi_{\text{down}}|$ represents the vertical field-of-view (FOV) of the sensor, with $\phi_{\text{up}}$ and $\phi_{\text{down}}$ being the inclination angles in the upward and downward directions, respectively; $H$ and $W$ are the height and width of the range image.
 However, such projection often results in the loss of occluded points and requires post-processing techniques to reconstruct the entire semantic information [36].
 We conducted a pilot study on various range-view methods, including RangeNet++ [34], SalsaNext [64], FIDNet [59], and CENet [61].
-We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with various nearest neighbor parameters, as well as their performance on the 2D range images.
+We compared their performances on the original points (i.e., after the 2D-to-3D projection) using the widely-used KNN post-processing with different nearest neighbor parameters, as well as their performance on the 2D range images.
 As illustrated in Fig.
@@
 Problem Definition.
-Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\hat{Y}}$ for each individual point as follows:
-$\mathcal{\hat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
+Given a point cloud $\mathcal{P}$ acquired by the LiDAR sensor, the objective of FRNet is to employ a feed-forward network $\mathcal{G}$ to predict semantic labels $\mathcal{\widehat{Y}}$ for each individual point as follows:
+$\mathcal{\widehat{Y}}=\mathcal{G}(\mathcal{P},\theta)~{},$ (2)
 where $\theta$ represents the learnable parameters within the network.
@@
 Frustum Feature Encoder.
-In LiDAR segmentation, the accurate prediction of semantic labels for each point requires the extraction of individual features for each point.
-Preceding the conversion of the point cloud into a 2D representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs.
-This process is crucial for accurate label prediction.
-Firstly, We group the point cloud to yield a set of $M$ frustum regions $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$ according to Eq.
-1, where $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$, consisting of $N_{i}$ points.
-To explicitly embed the frustum structure, we leverage an average pooling function to obtain the cluster points $\widetilde{\mathcal{P}}=\{\widetilde{\mathcal{P}}_{1},\widetilde{\mathcal{P}}_{2},...,\widetilde{\mathcal{P}}_{M}\}$, where $\widetilde{\mathcal{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$.
-Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\widetilde{\mathcal{P}}$ within each frustum region.
-The per-point features are obtained as follows:
+Accurate prediction of semantic labels for each point requires the extraction of individual features.
+Before converting the point cloud into a 2D frustum representation, the frustum feature encoder plays a pivotal role in extracting per-point features through a series of MLPs, which is crucial for accurate label prediction.
+The frustum feature encoder follows three steps.
+① We divide the point cloud into $H$ and $W$ frustum regions along the vertical and horizontal directions, respectively.
+According to Eq.
+1, for each point $\mathbf{p}_{i}$ in the point cloud $\mathcal{P}$, we calculate its frustum position $(u_{i},v_{i})$.
+Points sharing the same $(u,v)$ coordinates are grouped into the same frustum region.
+This results in a set of $M$ frustum regions, defined as $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},...,\mathcal{P}_{M}\}$, where $M=H\times W$ and $\mathcal{P}_{i}\in\mathbb{R}^{N_{i}\times(3+L)}$ consists of $N_{i}$ points within the $i$-th frustum region.
+② To explicitly embed the frustum structure, we leverage an average pooling function within each frustum region to obtain the cluster points $\mathcal{\widetilde{P}}=\{\mathcal{\widetilde{P}}_{1},\mathcal{\widetilde{P}}_{2},...,\mathcal{\widetilde{P}}_{M}\}$, where $\mathcal{\widetilde{P}}_{i}\in\mathbb{R}^{1\times(3+L)}$.
+Subsequently, we construct a frustum graph between the point cloud $\mathcal{P}$ and the cluster points $\mathcal{\widetilde{P}}$ within each frustum region.
+The per-point features are embedded as follows:
 $\mathcal{F}_{p}^{0}=\text{MLP}([\mathcal{P};\mathcal{P}-\widetilde{\mathcal{P}}])~{},$ (3)
 […]
-Frustum-Point Fusion Module.
-Following prior works [55, 58, 32], we utilize a 2D backbone composed of multiple convolution blocks that efficiently extract hierarchical frustum features.
-Each stage block is accompanied by a frustum-point fusion module, facilitating the incremental update of point and frustum features.
-As illustrated in Fig.
-4, the frustum-point fusion module consists of two essential components: 1) A frustum-to-point fusion that enables the hierarchical update of per-point features, and 2) A point-to-frustum fusion that fuses individual features of each point within its corresponding frustum region.
-The $i$-th frustum-point fusion module takes $\mathcal{F}_{p}^{i-1}$ and ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ as inputs, where ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is the output of the 2D convolution block that takes $\mathcal{F}_{f}^{i-1}$ as input.
-To efficiently extract context information of the point, we employ a frustum-to-point fusion block to integrate local features of the frustum region into points.
-Firstly, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index.
-Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update per-point features, avoiding the high-computation in neighborhood searches:
-$\mathcal{F}_{p}^{i}=\text{MLP}([\text{Inflat}(\widetilde{\mathcal{F}}_{f}^{i-1});\mathcal{F}_{p}^{i-1}])~{}.$ (4)
-As discussed above, the 2D representation is non-compact due to the sparsity of the point cloud.
-The convolution block inevitably introduces noisy information for the empty frustum, leading to compact feature representation.
-To preserve the sparsity attributes of the 2D representation, we design a point-to-frustum fusion module to fuse the sparse and dense frustum features.
-Specifically, the updated point features $\mathcal{F}_{p}^{i}$ are first pooled into the corresponding frustum region to yield a non-compact 2D representation.
-Subsequently, the non-compact frustum feature is concatenated with the compact frustum feature ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ and passed through a simple convolutional layer to reduce the number of channels:
-$\mathcal{F}_{\text{fuse}}^{i}=\text{Conv}([\text{Flat}(\mathcal{F}_{p}^{i});\widetilde{\mathcal{F}}_{f}^{i-1}])~{},$ (5)
-where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and activation function.
+[…]
+To efficiently extract context information for the points, we employ a frustum-to-point fusion block to integrate local features from the frustum region into the points.
+First, ~$\widetilde{\mathcal{F}}_{f}^{i-1}$ is back-projected to the point-level features according to the projection index.
+Subsequently, the back-projected features are concatenated with $\mathcal{F}_{p}^{i-1}$ and passed through an MLP to update the per-point features, avoiding high computational costs from neighborhood searches:
+[…]
+[…]
+[…]
+where $\text{Conv}(\cdot)$ represents a convolution layer with batch normalization and an activation function.
 Finally, the fused features are utilized to further enhance the frustum features in a residual-attentive manner:
-$\mathcal{F}_{f}^{i}=\widetilde{\mathcal{F}}_{f}^{i-1}+\sigma(h(\mathcal{F}_{\text{fuse}}^{i},\theta))\odot\mathcal{F}_{\text{fuse}}^{i}~{},$ (6)
+[…]
 where $\sigma$ represents the sigmoid function, $h(\cdot)$ is a linear function with trainable parameters $\theta$, and $\odot$ indicates element-wise multiplication.
 Figure 4: Frustum-point fusion module comprises two steps: 1) A Frustum-to-Point fusion to update per-point features. 2) A Point-to-Frustum fusion to update frustum features.
+(a) Scene 1
+(b) Scene 2
+(c) Mixed Scene
+Figure 5: FrustumMix illustration.
+(a) and (b) show the original two LiDAR scenes.
+(c) presents the mixed scenes generated by the FrustumMix strategy, where scene 1 is colored green and scene 2 is colored purple.
+(a) Before RangeInterpolation
+(b) After RangeInterpolation
+Figure 6: RangeInterpolation illustration.
+(a) shows the raw range image with numerous empty pixels.
+(b) illustrates the range image after applying RangeInterpolation, resulting in a smoother and more compact representation.
 Fusion Head Module.
-FRNet aims to assign a category for each point in the point cloud, which requires point-level features for end-to-end label prediction.
-Although output point features at the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution.
+FRNet aims to assign a category to each point in the point cloud, which requires point-level features for end-to-end label prediction.
+Although the output point features from the last layer of the frustum-point fusion module can yield promising performance, we observe this is not the optimal solution.
 Features from different levels have different context-aware information.
@@
 Progressively fusing these features helps extract features from local to global, and from geometric to semantic, resulting in more accurate predictions.
-To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project frustum features to point-level and use an MLP layer ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$.
+To fuse $\mathcal{F}_{f}^{\text{out}}$ and $\mathcal{F}_{p}^{\text{out}}$, we first back-project the frustum features to point-level and use an MLP layer to ensure the same channels as $\mathcal{F}_{p}^{\text{out}}$.
 Then, an addition operation is applied to fuse the two features.
@@
 To address this, we adopt a statistical approach inspired by sparse-voxel methods [16].
-We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label of the frustum.
+We generate a pseudo label for each frustum by first counting the frequency of each category existing in the given frustum region, and then choosing the category with the highest frequency as the pseudo label for the frustum.
 Finally, the frustum-level features are supervised using the generated pseudo labels.
@@
 where $\mathcal{L}_{p}$ and $\mathcal{L}_{f}$ represent the point-level and frustum-level loss functions, respectively, and $\lambda$ is a hyperparameter that controls the weight of frustum supervision.
-$\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovasz-Softmax loss [61], and boundary loss [62].
+$\mathcal{L}_{p}$ is calculated by cross-entropy loss, while $\mathcal{L}_{f}$ consists of cross-entropy loss, Lovász-Softmax loss [65], and boundary loss [66].
 III-C Frustum-Range Operators
 Data augmentation enables the model to learn more robust representations.
-Common augmentations in LiDAR segmentation typically operate on the global view, encompassing rotation, flipping, scaling, jittering, etc.
+Common augmentations in LiDAR segmentation typically operate from a global perspective, including rotation, flipping, scaling, jittering, etc.
 In this work, we introduce two novel augmentations based on our proposed frustum-range view representation.
 FrustumMix.
-Previous mixing strategies [44, 45] involve swapping two LiDAR scans at the point level, which may disrupt the internal structure within the frustum region.
-In this work, we propose FrustumMix, a method to swap frustum regions from two scenes.
-Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction via Eq.
-1, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$.
-We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\widetilde{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$.
-FrustumMix enhances context-awareness between frustums while preserving the invariant structure within the frustum.
+Previous mixing strategies [48, 49] perform point-level swapping between two LiDAR scans.
+While these methods achieve notable improvements in various LiDAR representations, they are not well aligned with the frustum representation, potentially causing two distinct scenes to mix within the same frustum region.
+This misalignment disrupts both the semantic coherence and geometric structure within the frustum region.
+To this end, we propose FrustumMix, a fine-grained extension of LaserMix [48] that aligns better with the frustum representation.
+Specifically, given two LiDAR point clouds $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, we randomly split the point clouds into $M$ non-overlapping frustum regions along the inclination or azimuth direction, denoted as $\mathcal{P}_{1}=\{\mathcal{A}_{1}^{1},...,\mathcal{A}_{1}^{M}\}$ and $\mathcal{P}_{2}=\{\mathcal{A}_{2}^{1},...,\mathcal{A}_{2}^{M}\}$.
+We then generate a new LiDAR scan $\hat{\mathcal{P}}$ by alternately swapping frustum regions from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, as follows: $\hat{\mathcal{P}}=\mathcal{A}_{1}^{1}\cup\mathcal{A}_{2}^{2}\cup\mathcal{A}_{1}^{3}\cup\mathcal{A}_{2}^{4}\cup\cdots$.
+This approach ensures that mixing occurs at the level of frustum units, preserving semantic coherence and structural details entirely within each frustum region, thereby maintaining alignment with the frustum representation.
+Fig.
+5 provides an illustration of FrustumMix.
+By enhancing context-awareness between frustums while preserving intra-frustum structural invariance, FrustumMix achieves improved segmentation performance and better representation learning.
 […]
@@
 RandLA-Net [8] CVPR’20 Raw Points $\bullet$ $1.2$ M $1.9$ $57.1$ $53.9$
-PTv2 [66] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $\mathbf{70.3}$ $\mathbf{72.6}$ $\mathbf{80.2}$ $\mathbf{82.6}$
-WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $\mathbf{7.4}$ $68.0$ $70.8$ $79.1$
+PTv2 [70] NeurIPS’22 Raw Points $\bullet$ $12.8$ M $4.7$ $70.3$ $72.6$ $80.2$ $82.6$
+WaffleIron [11] ICCV’23 Raw Points $\bullet$ $6.8$ M $7.4$ $68.0$ $70.8$ $79.1$
+PTv3 [12] CVPR’24 Raw Points $\bullet$ $46.2$ M $\mathbf{20.2}$ $\mathbf{70.8}$ $\mathbf{74.2}$ $\mathbf{80.4}$ $\mathbf{82.7}$
 RPVNet [22] ICCV’21 Multi-View $\bullet$ $24.8$ M $7.9$ $65.5$ $70.3$ $77.6$
-2DPASS [67] ECCV’22 Multi-View $\bullet$ $26.5$ M $\mathbf{8.4}$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
+2DPASS [71] ECCV’22 Multi-View $\bullet$ $26.5$ M $8.4$ $69.3$ $72.2$ $\mathbf{79.4}$ $80.8$
 SphereFormer [72] CVPR’23 Multi-View $\bullet$ $32.3$ M $4.9$ $67.8$ $74.8$ $78.4$ $81.9$
-UniSeg [23] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $\mathbf{71.3}$ $\mathbf{75.2}$ $\mathbf{83.5}$
+UniSeg [24] ICCV’23 Multi-View $\bullet$ $147.6$ M $6.9$ $71.3$ $75.2$ $83.5$
+TASeg [73] CVPR’24 Multi-View $\bullet$ $46.7$ M $6.1$ $\mathbf{72.7}$ $\mathbf{76.5}$ $\mathbf{84.6}$
 RangeNet++ [34] IROS’19 Range View $\bullet$ $50.4$ M $12.0$ $50.3$ $52.2$ $65.5$ $44.6$ $57.8$ $30.9$
@@
 CENet [61] ICME’22 Range View $\bullet$ $6.8$ M $33.4$ $62.6$ $64.7$ $73.3$ $55.7$ $66.8$ $50.3$
-RangeViT [31] CVPR’23 Range View $\bullet$ $23.7$ M $10.0$ $60.7$ $64.0$ $75.2$ $53.6$ $66.5$
-RangeFormer [32] ICCV’23 Range View $\bullet$ $24.3$ M $6.2$ $67.6$ $\mathbf{73.3}$ $78.1$ $80.1$ $63.0$
+[…]
+[…]
 Fast-FRNet Ours Frustum-Range $\bullet$ $7.5$ M $\mathbf{33.8}$ $67.1$ $72.5$ $78.8$ $82.1$ $62.4$ $71.2$ $52.5$ $67.1$
@@
 Specifically, the point cloud $\mathcal{P}$ is first projected onto the range image $\mathcal{R}\in\mathbb{R}^{H\times W\times(3+L)}$.
-For empty pixel position $(u,v)$ in the range image $\mathcal{R}$, we establish a window with size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd.
-Our aim is to interpolate the point falling into $(u,v)$ using the surrounding range information within the predefined window.
-Considering that not all pixels in the window contain valid values, we aggregate the information from non-empty pixels within the window for point interpolation using an average operation.
+For empty pixel positions $(u,v)$ in the range image $\mathcal{R}$, we establish a window of size $m\times n$ centered at $(u,v)$, where $m$ and $n$ are odd.
+We aim to interpolate the point falling into $(u,v)$ using the surrounding range information within the pre-defined window.
+Considering that not all pixels in the window contain valid values, we aggregate information from non-empty pixels within the window for point interpolation using an average operation.
 During the training procedure, the semantic label of the interpolated point is determined via threshold-guided voting within the window.
@@
 Otherwise, the label is determined by the category with the highest frequency.
+As shown in Fig.
+6, RangeInterpolation produces coherent semantic information in the range image, resulting in a more compact representation.
 IV Experiments
@@
 FrustumMix Ours $55.7$ $62.5$ $63.0$ $64.9$ $60.0$ $70.0$ $72.6$ $74.1$ $45.6$ $55.7$ $58.2$ $60.8$
+Sup.-only - PTv3 [12] $42.7$ $60.9$ $62.6$ $64.1$ $48.0$ $67.8$ $71.5$ $75.9$ $40.7$ $54.6$ $56.2$ $58.9$
+PolarMix [49] NeurIPS’22 $47.2$ $61.6$ $63.1$ $64.8$ $53.4$ $69.3$ $72.0$ $76.1$ $42.5$ $55.8$ $56.9$ $59.6$
+LaserMix [48] CVPR’23 $51.7$ $63.3$ $64.8$ $65.3$ $57.4$ $71.8$ $74.2$ $77.4$ $45.0$ $56.9$ $58.3$ $60.7$
+FrustumMix Ours $54.2$ $64.7$ $\mathbf{65.6}$ $\mathbf{66.1}$ $60.3$ $\mathbf{72.3}$ $\mathbf{75.6}$ $\mathbf{77.7}$ $46.2$ $57.5$ $60.0$ $\mathbf{61.6}$
 Sup.-only - FRNet $44.9$ $60.4$ $61.8$ $63.1$ $51.9$ $68.1$ $70.9$ $74.6$ $42.4$ $53.5$ $55.1$ $57.0$
@@
 LaserMix [48] CVPR’23 $52.9$ $62.9$ $63.2$ $65.0$ $58.7$ $71.5$ $72.3$ $75.0$ $45.8$ $56.8$ $57.7$ $59.0$
-FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $\mathbf{65.2}$ $\mathbf{65.4}$ $\mathbf{61.2}$ $\mathbf{72.2}$ $\mathbf{74.6}$ $\mathbf{75.4}$ $\mathbf{46.6}$ $57.0$ $59.5$ $\mathbf{61.2}$
+FrustumMix Ours $\mathbf{55.8}$ $\mathbf{64.8}$ $65.2$ $65.4$ $\mathbf{61.2}$ $72.2$ $74.6$ $75.4$ $\mathbf{46.6}$ $57.0$ $59.5$ $61.2$
 Figure 7: Class-wise LiDAR segmentation results of FRNet and the baseline model on the test set of SemanticKITTI [32].
@@
 Datasets.
-We conduct a comprehensive experiment on four popular LiDAR segmentation datasets.
+We conduct comprehensive experiments on four popular LiDAR segmentation datasets.
 SemanticKITTI [32] is a large-scale outdoor dataset that comprises $22$ sequences collected from various scenes in Karlsruhe, Germany, using a Velodyne HDL-64E LiDAR sensor.
@@
 We implement FRNet using the popular MMDetection3D [77] codebase and employ CENet [61] as the 2D backbone to extract frustum features.
-The 2D representation shape is set to $64\times 512$ for SemanticKITTI [28] and ScribbelKITTI [63], and $32\times 480$ for nuScenes [43] and SemanticPOSS [64], respectively.
+For SemanticKITTI [32] and ScribbelKITTI [67], the point cloud is divided into $64\times 512$ frustum regions, while for nuScenes [47] and SemanticPOSS [68], it is divided into $32\times 480$.
 For optimization, we choose AdamW [78] as our default optimizer with an initial learning rate of $0.01$.
 We utilized the OneCycle scheduler [79] to dynamically adjust the learning rate during training.
-All models are trained using four GPUs for $50$ epochs on SemanticKITTI [28] and ScribbelKITTI [63], and $80$ epochs on nuScenes [43] and SemanticPOSS [64], respectively.
+All models are trained using four GPUs for $50$ epochs on SemanticKITTI [32] and ScribbelKITTI [67], and for $80$ epochs on nuScenes [47] and SemanticPOSS [68], respectively.
 The batch size is set to $4$ for each GPU.
-Furthermore, we implement a faster variant, dubbed Fast-FRNet, which features a smaller 2D representation resolution and scaled-down 2D backbone.
-Specifically, the size of the 2D representation is adjusted to $32\times 360$ for all datasets, and the 2D backbone is a variant of ResNet18 [75], following the design of CENet [58].
+Additionally, we implement a faster variant, dubbed Fast-FRNet, which features a larger frustum area and a scaled-down 2D backbone.
+Specifically, the point cloud is divided into $32\times 360$ frustum regions for all datasets, and the 2D backbone is a variant of ResNet-18 [80], following the design of CENet [61].
 IV-B Comparative Study
@@
 I provides a summary of all the results.
-We evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for fair comparison.
-For SemanticKITTI [28] and nuScenes [43], we report mIoU scores on both the val and test sets.
-For ScribbleKITTI [63] and SemanticPOSS [64], which share the same data on val and test sets, we report both mIoU and mAcc scores on the val set.
-Compared with multi-view methods, although FRNet does not exhibit a significant advantage in performance, its parameters and FPS are notably superior.
-In comparison with sparse-voxel-view and range-view methods, FRNet achieves appealing performance while still maintaining high efficiency.
+In this benchmark, both FrustumMix and RangeInterpolation are employed as data augmentation techniques to enhance model performance.
+To ensure a fair comparison, we evaluate the Frame Per Second (FPS) on a single GeForce RTX 2080Ti GPU for all models.
+For SemanticKITTI [32] and nuScenes [47], we report mIoU scores on both the validation and test sets.
+For ScribbleKITTI [67] and SemanticPOSS [68], which share the same data across the val and test sets, we report both mIoU and mAcc scores on the val set.
+Compared with recent multi-view methods [24], although FRNet does not exhibit a significant advantage in performance, it shows notable superiority in terms of parameters and FPS.
+In comparison with range-view, sparse-voxel, and raw-points methods, FRNet achieves appealing performance while still maintaining high efficiency.
 Additionally, Fast-FRNet demonstrates the fastest inference speed with only a minor drop in performance.
@@
 Label-Efficient LiDAR Semantic Segmentation.
-Recently, semi-supervised learning has been explored in LiDAR segmentation.
+Recently, semi-supervised learning has gained traction in LiDAR segmentation.
 In this work, we adopt the approach of LaserMix [48] to apply FRNet to the semi-supervised segmentation task, as shown in Tab.
 II.
-Firstly, we utilize Cylinder3D [16] as the backbone and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning.
-Notably, FrustumMix demonstrates superior performance compared to the LaserMix [44] method.
-Subsequently, we employ FRNet as the backbone and assess various strategies for mixing labeled and unlabeled scans, including PolarMix [45], LaserMix [44], and our FrustumMix.
-The Sup.-only baseline results are obtained without FrustumMix and RangeInterpolation as data augmentation techniques.
-Results show that FrustumMix consistently achieves the best scores across all data splits, particularly with limited annotations.
+For fair comparisons with other methods, FrustumMix and RangeInterpolation are not employed as data augmentation techniques in this benchmark, and TTA is excluded during inference.
+We first utilize Cylinder3D [16] and PTv3 [12] as backbones and leverage the proposed FrustumMix technique to mix labeled and unlabeled data for consistency learning, which aligns with the LaserMix framework [48].
+Notably, FrustumMix demonstrates superior performance compared to the original LaserMix [48] strategy.
+Next, we employ FRNet as the backbone and evaluate various strategies for mixing labeled and unlabeled scans, including PolarMix [49], LaserMix [48], and our FrustumMix.
+The Sup.-only baseline results are also obtained without FrustumMix and RangeInterpolation as data augmentation techniques.
+The results indicate that FrustumMix consistently achieves the best scores across all data splits when compared with other mixing strategies, particularly with limited annotations.
 Out-of-Distribution Robustness.
@@
 To evaluate the out-of-distribution robustness of FRNet, we utilize the noisy datasets SemanticKITTI-C and nuScenes-C introduced in Robo3D [76].
-The models trained on SemanticKITTI [28] and nuScenes [43] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [71], respectively.
+In this benchmark, the models trained on SemanticKITTI [32] and nuScenes [47] are directly applied to evaluate the performance on SemanticKITTI-C and nuScenes-C [76], respectively, without any fine-tunning.
 As shown in Tab.
@@
 Best viewed in colors and zoomed-in for details.
-[…]
+Quantitative Assessments.
+We compare the class-wise IoU scores of FRNet with those of the baseline method [61] in Fig.
+7.
+Notably, FRNet demonstrates significant improvement across most semantic classes, particularly for dynamic classes with intricate structures, where it achieves mIoU gains ranging from $15\%$ to $24\%$.
+In Fig.
+8, we present the prediction results on SemanticKITTI [32], compared with those of state-of-the-art range view methods, including FIDNet [59], CENet [61], and RangeViT [35].
+It is evident that FRNet delivers more accurate predictions, particularly on large planar objects such as “vegetation”, “terrain”, and “sidewalk”, which are often challenging to distinguish accurately.
+IV-C Ablation Study
+In this section, we will discuss the efficacy of each design in our FRNet architecture.
+Unless otherwise specified, all experiments are conducted and reported on the val set of SemanticKITTI [32] and nuScenes [47], respectively.
+TABLE IV: Ablation study of each component in FRNet on the val set of SemanticKITTI [32] and nuScenes [47].
+FFE: Frustum Feature Encoder.
+FP: Frustum-Point Fusion.
+FH: Fusion Head.
+FS: Frustum-level Supervision.
+RI: RangeInterpolation.
+TTA: Test Time Augmentation.
+All mIoU and mAcc scores are given in percentage (%).
 FFE FP FH FS RI TTA SemKITTI nuScenes
@@
 $\mathbf{68.7}$ $\mathbf{74.9}$ $\mathbf{79.0}$ $85.0$
-Quantitative Assessments.
-We compare the class-wise IoU scores of FRNet with those of the baseline method [58] in Fig.
-5.
-Notably, FRNet demonstrates significant improvement across most semantic classes, particularly for dynamic classes with intricate structures, where it achieves mIoU gains ranging from $15\%$ to $24\%$.
-In Fig.
-6, we present the prediction results on SemanticKITTI [28], compared with those of state-of-the-art range view methods, including FIDNet [55], CENet [58], and RangeViT [31].
-It is evident that FRNet delivers more accurate predictions, particularly on large planar objects such as “vegetation”, “terrain”, and “sidewalk”, which are often challenging to distinguish accurately.
-IV-C Ablation Study
-In this section, we will discuss the efficacy of each design in our FRNet architecture.
-Unless otherwise specified, all experiments are conducted and reported on the val set of SemanticKITTI [28] and nuScenes [43], respectively.
 Component Designs.
@@
 IV summarizes the ablation results of each component in the FRNet architecture.
-Firstly, We directly apply the frustum feature encoder to CENet [58], incorporating KNN post-processing to predict semantic labels over the entire point cloud, yielding $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [28] and nuScenes [43], respectively.
-Subsequently, by integrating the point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in an improvement of $1.6\%$ and $0.8\%$ mIoU, respectively.
+Notably, FrustumMix is employed across all setups in this ablation study.
+Firstly, we directly apply the frustum feature encoder to CENet [61], incorporating KNN post-processing to predict semantic labels over the entire point cloud, achieving $62.7\%$ and $72.8\%$ mIoU on SemanticKITTI [32] and nuScenes [47], respectively.
+Next, by integrating point-frustum fusion to hierarchically update point features, the semantic results are predicted with the updated point features in an end-to-end manner, resulting in improvements of $1.6\%$ and $0.8\%$ mIoU, respectively.
 To regularize the frustum features, the frustum-level supervision is introduced and it leads to significant gains of $1.9\%$ and $2.8\%$ mIoU, respectively.
@@
 Furthermore, the proposed RangeInterpolation, aimed at alleviating empty pixels in the 2D representation by reconstructing semantic surfaces via surrounding range information, provides a performance boost of approximately $0.5\%$ mIoU.
-Finally, adopting Test Time Augmentation during inference, following prior works, brings an improvement of $1.1\%$ and $1.3\%$ mIoU, respectively.
+Finally, adopting Test Time Augmentation during inference, as demonstrated in prior works, leads to improvements of $1.1\%$ and $1.3\%$ mIoU, respectively.
+[…]
+Number of Frustums Height $(H)$ Width $(W)$ mIoU mAcc
+$131,072$ $64$ $2048$ $66.8$ $73.8$
+$65,536$ $64$ $1024$ $67.1$ $73.4$
+$65,536$ $32$ $2048$ $65.5$ $72.7$
+$32,768$ $64$ $512$ $\mathbf{67.6}$ $\mathbf{74.2}$
+$32,768$ $32$ $1024$ $66.0$ $72.4$
+$16,384$ $64$ $256$ $64.3$ $70.5$
+$16,384$ $32$ $512$ $65.5$ $72.1$
+$8,192$ $32$ $256$ $64.3$ $70.8$
+Frustum Representation Resolution.
+Finding a suitable frustum resolution is crucial for balancing accuracy and efficiency.
+In Tab.
+V, we conduct a series of experiments with various frustum resolutions.
+While higher resolution can capture more detailed features, it may not always lead to optimal results.
+The limited area of the frustum region can result in a lack of meaningful structure, causing the model to struggle with noisy or sparse data.
+Conversely, although lower resolution can improve inference efficiency and provide a more coherent representation, the larger area of the frustum region may introduce excessive noise information, particularly at object boundaries.
+Thus, finding an optimal frustum resolution is crucial for balancing the richness of features with computational demands and maintaining overall model performance.
 Figure 9: Ablation study on different FRNet configurations.
@@
 As illustrated in Fig.
-7(b) and Fig.
-7(c), FrustumMix consistently outperforms these prior strategies.
-TABLE V: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [28].
+9(b) and Fig.
+9(c), FrustumMix consistently outperforms these prior strategies.
+TABLE VI: Ablation study on window settings in RangeInterpolation on the val set of SemanticKITTI [32].
 Green pixel denotes the empty pixel needed to be interpolated.
@@
 Tab.
-V showcases the performance of RangeInterpolation with different window settings.
+VI showcases the performance of RangeInterpolation with different window settings.
 Notably, we observe consistent performance improvements when considering interpolation along the horizontal or vertical direction exclusively.
@@
 As depicted in Fig.
-7(a), our proposed frustum-range representation consistently delivers performance gains ranging from $5.0\%$ to $6.8\%$ in mIoU across different range-view methods.
+9(a), our proposed frustum-range representation consistently delivers performance gains ranging from $5.0\%$ to $6.8\%$ in mIoU across different range-view methods.
+TABLE VII: Inference efficiency comparison of FRNet and MinkUNet [17] implemented with various codebases and sparse convolution backends, as well as PTv3 [12] on the val set of SemanticKITTI [32].
+Codebase Backbone Backend FPS
+MMDetection3D MinkUNet MinkowskiEngine $9.1$
+MinkUNet SpConv $8.6$
+MinkUNet TorchSparse $10.2$
+MinkUNet TorchSparse++ $21.3$
+FRNet Convolution $29.1$
+Fast-FRNet Convolution $\mathbf{33.8}$
+PCSeg MinkUNet TorchSparse $10.6$
+PointCept MinkUNet MinkowskiEngine $8.9$
+PTv3 FlashAttention $20.2$
+Inference Efficiency.
+Multiple sparse convolution backends, including MinknowskiEngine, SpConv, TorchSparse, and TorchSparse++, offer efficient processing for voxel-based methods.
+To compare FRNet with these backends, we implement MinkUNet [17] using the MMDetection3D codebase across various backends.
+Additionally, we evaluate the FPS of MinkUNet implemented with the PCSeg and PointCept codebases, as well as the recent efficient method PTv3 [12].
+All evaluations are conducted on a single NVIDIA 2080Ti GPU.
+As shown in Tab.
+VII, while TorchSparse++ delivers satisfactory efficiency for LiDAR segmentation, FRNet achieves the fastest inference speed, making it suitable for real-time applications in autonomous driving systems.
+(a) car
+(b) person
+(c) vegetation
+(d) drivable surface
+(e) sidewalk
+(f) trailer
+Figure 10: The open-vocabulary LiDAR semantic segmentation experiments of FRNet.
+Our model demonstrates the ability to perform open-vocabulary segmentation with various textual inputs.
+The segmented points corresponding to the provided text descriptions are highlighted in blue, showcasing its potential for open-world perception tasks.
+TABLE VIII: The multi-source training results of the proposed Fast-FRNet on the val set of the SemanticKITTI [32], nuScenes [47], and SemanticPOSS [68] datasets.
+All mIoU and mAcc scores are given in percentage (%).
+Method SemKITTI nuScenes SemPOSS
+mIoU mAcc mIoU mAcc mIoU mAcc
+Single-Source $67.1$ $73.8$ $78.8$ $84.4$ $52.5$ $67.1$
+Multi-Source $67.4$ $74.2$ $79.1$ $84.7$ $53.2$ $68.0$
+Fast-FRNet (w/ PPT [83]) $\mathbf{68.2}$ $\mathbf{74.7}$ $\mathbf{79.3}$ $\mathbf{84.9}$ $\mathbf{53.9}$ $\mathbf{68.4}$
+TABLE IX: Zero-shot segmentation performance on the val set of nuScenes [47]. All mIoU and mAcc scores are given in percentage (%).
+Method nuScenes
+mIoU mAcc
+OpenScene [84] $42.1$ $61.8$
++ FRNet $\mathbf{43.5}$ $\mathbf{62.0}$
+Multi-Dataset Fusion.
+Recent studies [83, 51] highlight the importance of large-scale benchmarks achieved through multi-dataset fusion strategies across heterogeneous datasets.
+In this work, We extend Fast-FRNet to support multi-dataset joint training, as shown in Tab.
+VIII.
+Initially, we combine multi-source datasets using mean-variance normalization to align intensity values, which provides slight performance gains compared to single-source training.
+Additionally, inspired by PPT [83], we incorporate point-prompt techniques to facilitate multi-source joint training.
+This approach yields further improvements, demonstrating the effectiveness of prompt-guided strategies in multi-dataset fusion.
+Open-Vocabulary Semantic Segmentation.
+To investigate the potential of FRNet in open-vocabulary segmentation, we adopt an approach inspired by OpenScene [84].
+Specifically, we integrate text and image features into the framework, aligning them with point features.
+During inference, text inputs are used to guide the segmentation of point clouds with similar features.
+As shown in Tab.
+IX, FRNet demonstrates promising performance in zero-shot segmentation.
+Additionally, in Fig.
+10, we present qualitative results of segmentation using corresponding text inputs, highlighting the capability of FRNet to handle open-world perception tasks effectively.
 V Conclusion
@@
 We hope this work can lay a solid foundation for future works targeted at real-time, in-vehicle 3D perception.
-Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction.
+Moving forward, we will further enhance the representation learning capability of FRNet and incorporate such a method into other 3D perception tasks, such as 3D object detection and semantic occupancy prediction, and comprehensive vehicle-road coordination perception tasks.
 Appendix
@@
 As shown in Fig.
-8, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed.
+11, current LiDAR segmentation models pursue either the segmentation accuracy or inference speed.
 Those more accurate LiDAR segmentation models often contain larger parameter sets and are less efficient during the inference stage.
@@
 Our models achieve competitive segmentation results and superior robustness over the voxel-based and point-voxel fusion counterparts, while still maintaining high efficiency for real-time LiDAR segmentation.
-TABLE VI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [28] leaderboard.
+TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticKITTI [32] leaderboard.
 All IoU scores are given in percentage (%).
@@
 Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign PointNet [85] $14.6$ $46.3$ $1.3$ $0.3$ $0.1$ $0.8$ $0.2$ $0.2$ $0.0$ $61.6$ $15.8$ $35.7$ $1.4$ $41.4$ $12.9$ $31.0$ $4.6$ $17.6$ $2.4$ $3.7$ PointNet++ [63] $20.1$ $53.7$ $1.9$ $0.2$ $0.9$ $0.2$ $0.9$ $1.0$ $0.0$ $72.0$ $18.7$ $41.8$ $5.6$ $62.3$ $16.9$ $46.5$ $13.8$ $30.0$ $6.0$ $8.9$ SqSeg [33] $30.8$ $68.3$ $18.1$ $5.1$ $4.1$ $4.8$ $16.5$ $17.3$ $1.2$ $84.9$ $28.4$ $54.7$ $4.6$ $61.5$ $29.2$ $59.6$ $25.5$ $54.7$ $11.2$ $36.3$ SqSegV2 [54] $39.6$ $82.7$ $21.0$ $22.6$ $14.5$ $15.9$ $20.2$ $24.3$ $2.9$ $88.5$ $42.4$ $65.5$ $18.7$ $73.8$ $41.0$ $68.5$ $36.9$ $58.9$ $12.9$ $41.0$ RandLA-Net [8] $50.3$ $94.0$ $19.8$ $21.4$ $42.7$ $38.7$ $47.5$ $48.8$ $4.6$ $90.4$ $56.9$ $67.9$ $15.5$ $81.1$ $49.7$ $78.3$ $60.3$ $59.0$ $44.2$ $38.1$ RangeNet++ [34] $52.2$ $91.4$ $25.7$ $34.4$ $25.7$ $23.0$ $38.3$ $38.8$ $4.8$ $91.8$ $65.0$ $75.2$ $27.8$ $87.4$ $58.6$ $80.5$ $55.1$ $64.6$ $47.9$ $55.9$ PolarNet [37] $54.3$ $93.8$ $40.3$ $30.1$ $22.9$ $28.5$ $43.2$ $40.2$ $5.6$ $90.8$ $61.7$ $74.4$ $21.7$ $90.0$ $61.3$ $84.0$ $65.5$ $67.8$ $51.8$ $57.5$ MPF [23] $55.5$ $93.4$ $30.2$ $38.3$ $26.1$ $28.5$ $48.1$ $46.1$ $18.1$ $90.6$ $62.3$ $74.5$ $30.6$ $88.5$ $59.7$ $83.5$ $59.7$ $69.2$ $49.7$ $58.1$ 3D-MiniNet [86] $55.8$ $90.5$ $42.3$ $42.1$ $28.5$ $29.4$ $47.8$ $44.1$ $14.5$ $91.6$ $64.2$ $74.5$ $25.4$ $89.4$ $60.8$ $82.8$ $60.8$ $66.7$ $48.0$ $56.6$ SqSegV3 [56] $55.9$ $92.5$ $38.7$ $36.5$ $29.6$ $33.0$ $45.6$ $46.2$ $20.1$ $91.7$ $63.4$ $74.8$ $26.4$ $89.0$ $59.4$ $82.0$ $58.7$ $65.4$ $49.6$ $58.9$ KPConv [7] $58.8$ $96.0$ $32.0$ $42.5$ $33.4$ $44.3$ $61.5$ $61.6$ $11.8$ $88.8$ $61.3$ $72.7$ $31.6$ $95.0$ $64.2$ $84.8$ $69.2$ $69.1$ $56.4$ $47.4$ SalsaNext [64] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $19.4$ $91.7$ $63.7$ $75.8$ $29.1$ $90.2$ $64.2$ $81.8$ $63.6$ $66.5$ $54.3$ $62.1$ FIDNet [59] $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $23.7$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$ FusionNet [87] $61.3$ $95.3$ $47.5$ $37.7$ $41.8$ $34.5$ $59.5$ $56.8$ $11.9$ $91.8$ $68.8$ $77.1$ $30.8$ $92.5$ $69.4$ $84.5$ $69.8$ $68.5$ $60.4$ $66.5$ PCSCNet [88] $62.7$ $95.7$ $48.8$ $46.2$ $36.4$ $40.6$ $55.5$ $68.4$ $55.9$ $89.1$ $60.2$ $72.4$ $23.7$ $89.3$ $64.3$ $84.2$ $68.2$ $68.1$ $60.5$ $63.9$ KPRNet [60] $63.1$ $95.5$ $54.1$ $47.9$ $23.6$ $42.6$ $65.9$ $65.0$ $16.5$ $93.2$ $73.9$ $80.6$ $30.2$ $91.7$ $68.4$ $85.7$ $69.8$ $71.2$ $58.7$ $64.1$ TornadoNet [89] $63.1$ $94.2$ $55.7$ $48.1$ $40.0$ $38.2$ $63.6$ $60.1$ $34.9$ $89.7$ $66.3$ $74.5$ $28.7$ $91.3$ $65.6$ $85.6$ $67.0$ $71.5$ $58.0$ $65.9$ LiteHDSeg [66] $63.8$ $92.3$ $40.0$ $55.4$ $37.7$ $39.6$ $59.2$ $71.6$ $54.3$ $93.0$ $68.2$ $78.3$ $29.3$ $91.5$ $65.0$ $78.2$ $65.8$ $65.1$ $59.5$ $67.7$ RangeViT [35] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $38.1$ $93.1$ $70.2$ $80.0$ $32.5$ $92.0$ $69.0$ $85.3$ $70.6$ $71.2$ $60.8$ $64.7$ CENet [61] $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $43.5$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $70.0$ $61.5$ $67.6$ SVASeg [18] $65.2$ $96.7$ $56.4$ $57.0$ $49.1$ $56.3$ $70.6$ $67.0$ $15.4$ $92.3$ $65.9$ $76.5$ $23.6$ $91.4$ $66.1$ $85.2$ $72.9$ $67.8$ $63.9$ $65.2$ AMVNet [21] $65.3$ $96.2$ $59.9$ $54.2$ $48.8$ $45.7$ $71.0$ $65.7$ $11.0$ $90.1$ $71.0$ $75.8$ $32.4$ $92.4$ $69.1$ $85.6$ $71.7$ $69.6$ $62.7$ $67.2$ GFNet [90] $65.4$ $96.0$ $53.2$ $48.3$ $31.7$ $47.3$ $62.8$ $57.3$ $44.7$ $\mathbf{93.6}$ $72.5$ $\mathbf{80.8}$ $31.2$ $\mathbf{94.0}$ $\mathbf{73.9}$ $85.2$ $71.1$ $69.3$ $61.8$ $68.0$ JS3C-Net [91] $66.0$ $95.8$ $59.3$ $52.9$ $54.3$ $46.0$ $69.5$ $65.4$ $39.9$ $88.9$ $61.9$ $72.1$ $31.9$ $92.5$ $70.8$ $84.5$ $69.8$ $67.9$ $60.7$ $68.7$ MaskRange [92] $66.1$ $94.2$ $56.0$ $55.7$ $59.2$ $52.4$ $67.6$ $64.8$ $31.8$ $91.7$ $70.7$ $77.1$ $29.5$ $90.6$ $65.2$ $84.6$ $68.5$ $69.2$ $60.2$ $66.6$ SPVNAS [15] $66.4$ $97.3$ $51.5$ $50.8$ $59.8$ $58.8$ $65.7$ $65.2$ $43.7$ $90.2$ $67.6$ $75.2$ $16.9$ $91.3$ $65.9$ $86.1$ $73.4$ $71.0$ $64.2$ $66.9$ MSSNet [93] $66.7$ $96.8$ $52.2$ $48.5$ $54.4$ $56.3$ $67.0$ $70.9$ $49.3$ $90.1$ $65.5$ $74.9$ $30.2$ $90.5$ $64.9$ $84.9$ $72.7$ $69.2$ $63.2$ $65.1$ Cylinder3D [16] $68.9$ $97.1$ $67.6$ $63.8$ $50.8$ $58.5$ $73.7$ $69.2$ $48.0$ $92.2$ $65.0$ $77.0$ $32.3$ $90.7$ $66.5$ $85.6$ $72.5$ $69.8$ $62.4$ $66.2$ AF2S3Net [94] $69.7$ $94.5$ $65.4$ $\mathbf{86.8}$ $39.2$ $41.1$ $\mathbf{80.7}$ $80.4$ $74.3$ $91.3$ $68.8$ $72.5$ $\mathbf{53.5}$ $87.9$ $63.2$ $70.2$ $68.5$ $53.7$ $61.5$ $71.0$ RPVNet [22] $70.3$ $97.6$ $68.4$ $68.7$ $44.2$ $61.1$ $75.9$ $74.4$ $73.4$ $93.4$ $70.3$ $80.7$ $33.3$ $93.5$ $72.1$ $86.5$ $75.1$ $71.7$ $64.8$ $61.4$ SDSeg3D [95] $70.4$ $97.4$ $58.7$ $54.2$ $54.9$ $65.2$ $70.2$ $74.4$ $52.2$ $90.9$ $69.4$ $76.7$ $41.9$ $93.2$ $71.1$ $86.1$ $74.3$ $71.1$ $65.4$ $70.6$ GASN [96] $70.7$ $96.9$ $65.8$ $58.0$ $59.3$ $61.0$ $80.4$ $\mathbf{82.7}$ $46.3$ $89.8$ $66.2$ $74.6$ $30.1$ $92.3$ $69.6$ $87.3$ $73.0$ $72.5$ $66.1$ $71.6$ PVKD [69] $71.2$ $97.0$ $67.9$ $69.3$ $53.5$ $60.2$ $75.1$ $73.5$ $50.5$ $91.8$ $70.9$ $77.5$ $41.0$ $92.4$ $69.4$ $86.5$ $73.8$ $71.9$ $64.9$ $65.8$ Fast-FRNet $72.5$ $97.1$ $66.0$ $73.1$ $59.3$ $65.7$ $76.0$ $78.9$ $54.5$ $91.8$ $72.6$ $77.6$ $42.0$ $92.4$ $70.6$ $86.5$ $71.9$ $72.5$ $63.1$ $66.7$ 2DPASS [71] $72.9$ $97.0$ $63.6$ $63.4$ $61.1$ $61.5$ $77.9$ $81.3$ $74.1$ $89.7$ $67.4$ $74.7$ $40.0$ $93.5$ $72.9$ $86.2$ $73.9$ $71.0$ $65.0$ $70.4$ RangeFormer [36] $73.3$ $96.7$ $69.4$ $73.7$ $59.9$ $66.2$ $78.1$ $75.9$ $58.1$ $92.4$ $73.0$ $78.8$ $42.4$ $92.3$ $70.1$ $86.6$ $73.3$ $72.8$ $66.4$ $66.6$ FRNet $73.3$ $97.3$ $67.9$ $74.6$ $59.4$ $66.3$ $78.1$ $79.2$ $57.3$ $92.1$ $73.0$ $78.1$ $41.8$ $92.7$ $71.0$ $86.7$ $73.2$ $72.5$ $64.7$ $67.3$ SphereFormer [72] $74.8$ $97.5$ $70.1$ $70.5$ $59.6$ $67.7$ $79.0$ $80.4$ $\mathbf{75.3}$ $91.8$ $69.7$ $78.2$ $41.3$ $93.8$ $72.8$ $86.7$ $75.1$ $72.4$ $66.8$ $\mathbf{72.9}$ UniSeg [24] $\mathbf{75.2}$ $\mathbf{97.9}$ $\mathbf{71.9}$ $75.2$ $\mathbf{63.6}$ $\mathbf{74.1}$ $78.9$ $74.8$ $60.6$ $92.6$ $\mathbf{74.0}$ $79.5$ $46.1$ $93.4$ $72.7$ $\mathbf{87.5}$ $\mathbf{76.3}$ $\mathbf{73.1}$ $\mathbf{68.3}$ $68.5$
-TABLE VII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [43].
+TABLE XI: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the val set of nuScenes [47].
 All IoU scores are given in percentage (%).
@@
 Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation AF2S3Net [94] [’21] $62.2$ $60.3$ $12.6$ $82.3$ $80.0$ $20.1$ $62.0$ $59.0$ $49.0$ $42.2$ $67.4$ $94.2$ $68.0$ $64.1$ $68.6$ $82.9$ $82.4$ RangeNet++ [34] [’19] $65.5$ $66.0$ $21.3$ $77.2$ $80.9$ $30.2$ $66.8$ $69.6$ $52.1$ $54.2$ $72.3$ $94.1$ $66.6$ $63.5$ $70.1$ $83.1$ $79.8$ PolarNet [37] [’20] $71.0$ $74.7$ $28.2$ $85.3$ $90.9$ $35.1$ $77.5$ $71.3$ $58.8$ $57.4$ $76.1$ $96.5$ $71.1$ $74.7$ $74.0$ $87.3$ $85.7$ PCSCNet [88] [’22] $72.0$ $73.3$ $42.2$ $87.8$ $86.1$ $44.9$ $82.2$ $76.1$ $62.9$ $49.3$ $77.3$ $95.2$ $66.9$ $69.5$ $72.3$ $83.7$ $82.5$ SalsaNext [64] [’20] $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $63.1$ $61.3$ $76.5$ $96.0$ $70.8$ $71.2$ $71.5$ $86.7$ $84.4$ SVASeg [18] [’22] $74.7$ $73.1$ $44.5$ $88.4$ $86.6$ $48.2$ $80.5$ $77.7$ $65.6$ $57.5$ $82.1$ $96.5$ $70.5$ $74.7$ $74.6$ $87.3$ $86.9$ RangeViT [35] [’23] $75.2$ $75.5$ $40.7$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $66.3$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $89.9$ $87.2$ Cylinder3D [16] [’21] $76.1$ $76.4$ $40.3$ $91.2$ $\mathbf{93.8}$ $51.3$ $78.0$ $78.9$ $64.9$ $62.1$ $84.4$ $96.8$ $71.6$ $76.4$ $75.4$ $90.5$ $87.4$ AMVNet [21] [’20] $76.1$ $\mathbf{79.8}$ $32.4$ $82.2$ $86.4$ $\mathbf{62.5}$ $81.9$ $75.3$ $\mathbf{72.3}$ $\mathbf{83.5}$ $65.1$ $\mathbf{97.4}$ $67.0$ $\mathbf{78.8}$ $74.6$ $90.8$ $87.9$ RPVNet [22] [’21] $77.6$ $78.2$ $43.4$ $92.7$ $93.2$ $49.0$ $85.7$ $80.5$ $66.0$ $66.9$ $84.0$ $96.9$ $73.5$ $75.9$ $70.6$ $90.6$ $88.9$ RangeFormer [36] [’23] $78.1$ $78.0$ $\mathbf{45.2}$ $94.0$ $92.9$ $58.7$ $83.9$ $77.9$ $69.1$ $63.7$ $85.6$ $96.7$ $74.5$ $75.1$ $75.3$ $89.1$ $87.5$ SphereFormer [72] [’23] $78.4$ $77.7$ $43.8$ $94.5$ $93.1$ $52.4$ $\mathbf{86.9}$ $\mathbf{81.2}$ $65.4$ $73.4$ $85.3$ $97.0$ $73.4$ $75.4$ $75.0$ $\mathbf{91.0}$ $\mathbf{89.2}$ Fast-FRNet $78.8$ $78.7$ $42.3$ $\mathbf{95.6}$ $93.1$ $58.9$ $86.3$ $77.9$ $66.9$ $72.1$ $85.4$ $97.0$ $\mathbf{76.3}$ $76.5$ $76.2$ $89.7$ $87.8$ FRNet $\mathbf{79.0}$ $78.5$ $43.9$ $95.4$ $93.2$ $56.3$ $85.8$ $79.0$ $68.5$ $72.8$ $\mathbf{86.5}$ $97.1$ $75.9$ $77.0$ $\mathbf{76.4}$ $89.7$ $88.0$
-TABLE VIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [43].
+TABLE XII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the test set of nuScenes [47].
 All IoU scores are given in percentage (%).
@@
 Method (year) mIoU barrier bicycle bus car construction-vehicle motorcycle pedestrian traffic-cone trailer truck driveable-surface other-ground sidewalk terrain manmade vegetation PolarNet [37] [’20] $69.4$ $72.2$ $16.8$ $77.0$ $86.5$ $51.1$ $69.7$ $64.8$ $54.1$ $69.7$ $63.5$ $96.6$ $67.1$ $77.7$ $72.1$ $87.1$ $84.5$ JS3C-Net [91] [’21] $73.6$ $80.1$ $26.2$ $87.8$ $84.5$ $55.2$ $72.6$ $71.3$ $66.3$ $76.8$ $71.2$ $96.8$ $64.5$ $76.9$ $74.1$ $87.5$ $86.1$ PMF [97] [’21] $77.0$ $82.0$ $40.0$ $81.0$ $88.0$ $64.0$ $79.0$ $80.0$ $76.0$ $81.0$ $67.0$ $97.0$ $68.0$ $78.0$ $74.0$ $90.0$ $88.0$ Cylinder3D [16] [’21] $77.2$ $82.8$ $29.8$ $84.3$ $89.4$ $63.0$ $79.3$ $77.2$ $73.4$ $84.6$ $69.1$ $97.7$ $70.2$ $80.3$ $75.5$ $90.4$ $87.6$ AMVNet [21] [’20] $77.3$ $80.6$ $32.0$ $81.7$ $88.9$ $67.1$ $84.3$ $76.1$ $73.5$ $84.9$ $67.3$ $97.5$ $67.4$ $79.4$ $75.5$ $91.5$ $88.7$ SPVCNN [15] [’20] $77.4$ $80.0$ $30.0$ $91.9$ $90.8$ $64.7$ $79.0$ $75.6$ $70.9$ $81.0$ $74.6$ $97.4$ $69.2$ $80.0$ $76.1$ $89.3$ $87.1$ AF2S3Net [94] [’21] $78.3$ $78.9$ $52.2$ $89.9$ $84.2$ $77.4$ $74.3$ $77.3$ $72.0$ $83.9$ $73.8$ $97.1$ $66.5$ $77.5$ $74.0$ $87.7$ $86.8$ 2D3DNet [98] [’21] $80.0$ $83.0$ $59.4$ $88.0$ $85.1$ $63.7$ $84.4$ $82.0$ $76.0$ $84.8$ $71.9$ $96.9$ $67.4$ $79.8$ $76.0$ $92.1$ $89.2$ RangeFormer [36] [’23] $80.1$ $85.6$ $47.4$ $91.2$ $90.9$ $70.7$ $84.7$ $77.1$ $74.1$ $83.2$ $72.6$ $97.5$ $70.7$ $79.2$ $75.4$ $91.3$ $88.9$ GASN [96] [’22] $80.4$ $85.5$ $43.2$ $90.5$ $92.1$ $64.7$ $86.0$ $83.0$ $73.3$ $83.9$ $75.8$ $97.0$ $71.0$ $81.0$ $\mathbf{77.7}$ $91.6$ $\mathbf{90.2}$ 2DPASS [71] [’22] $80.8$ $81.7$ $55.3$ $92.0$ $91.8$ $73.3$ $86.5$ $78.5$ $72.5$ $84.7$ $75.5$ $97.6$ $69.1$ $79.9$ $75.5$ $90.2$ $88.0$ LidarMultiNet [99] [’23] $81.4$ $80.4$ $48.4$ $94.3$ $90.0$ $71.5$ $87.2$ $\mathbf{85.2}$ $\mathbf{80.4}$ $86.9$ $74.8$ $97.8$ $67.3$ $80.7$ $76.5$ $92.1$ $89.6$ SphereFormer [72] [’23] $81.9$ $83.3$ $39.2$ $\mathbf{94.7}$ $\mathbf{92.5}$ $77.5$ $84.2$ $84.4$ $79.1$ $\mathbf{88.4}$ $\mathbf{78.3}$ $\mathbf{97.9}$ $69.0$ $\mathbf{81.5}$ $77.2$ $\mathbf{93.4}$ $\mathbf{90.2}$ Fast-FRNet $82.1$ $85.3$ $62.9$ $92.1$ $91.5$ $76.5$ $87.7$ $75.9$ $73.2$ $86.2$ $76.0$ $97.8$ $\mathbf{71.9}$ $80.9$ $77.1$ $90.8$ $88.1$ FRNet $82.5$ $85.8$ $65.4$ $92.1$ $91.6$ $77.4$ $87.9$ $77.4$ $74.3$ $86.0$ $75.7$ $97.8$ $71.8$ $80.8$ $77.0$ $91.0$ $88.3$ UniSeg [24] [’23] $\mathbf{83.5}$ $\mathbf{85.9}$ $\mathbf{71.2}$ $92.1$ $91.6$ $\mathbf{80.5}$ $\mathbf{88.0}$ $80.9$ $76.0$ $86.3$ $76.7$ $97.7$ $71.8$ $80.7$ $76.7$ $91.3$ $88.8$
-TABLE IX: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [63] leaderboard.
+TABLE XIII: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the ScribbleKITTI [67] leaderboard.
 All IoU scores are given in percentage (%).
@@
 Method mIoU car bicycle motorcycle truck other-vehicle person bicyclist motorcyclist road parking sidewalk other-ground building fence vegetation trunk terrain pole traffic-sign RangeNet++ [34] $44.6$ $84.6$ $24.6$ $38.9$ $13.7$ $12.9$ $29.0$ $51.4$ $0.0$ $86.0$ $35.2$ $72.4$ $4.5$ $80.2$ $41.6$ $80.1$ $51.2$ $66.9$ $43.9$ $34.5$ SalsaNext [64] $50.3$ $85.6$ $33.8$ $42.7$ $8.4$ $20.8$ $64.3$ $71.5$ $0.2$ $85.6$ $31.6$ $72.0$ $1.1$ $84.0$ $39.9$ $81.0$ $62.0$ $63.6$ $60.5$ $46.3$ RangeViT [35] $53.6$ $85.6$ $31.6$ $50.1$ $40.3$ $36.3$ $57.6$ $68.7$ $0.0$ $86.1$ $32.6$ $75.2$ $0.3$ $87.9$ $49.3$ $83.6$ $62.8$ $67.5$ $59.6$ $43.7$ FIDNet [59] $54.1$ $85.6$ $36.7$ $48.7$ $60.8$ $38.4$ $63.3$ $68.2$ $0.0$ $84.1$ $25.9$ $71.2$ $0.4$ $85.6$ $41.3$ $81.7$ $64.1$ $62.7$ $61.5$ $48.0$ MinkNet [17] $55.0$ $88.1$ $13.2$ $55.1$ $72.3$ $36.9$ $61.3$ $77.1$ $0.0$ $83.4$ $32.7$ $71.0$ $0.3$ $90.0$ $50.0$ $84.1$ $66.6$ $65.8$ $61.6$ $35.2$ CENet [61] $55.7$ $86.1$ $39.4$ $53.2$ $61.0$ $46.1$ $69.2$ $72.2$ $0.0$ $85.7$ $28.7$ $72.6$ $1.1$ $85.8$ $43.1$ $81.8$ $64.2$ $63.8$ $59.6$ $45.0$ SPVCNN [15] $56.9$ $88.6$ $25.7$ $55.9$ $67.4$ $48.8$ $65.0$ $78.2$ $0.0$ $82.6$ $30.4$ $70.1$ $0.3$ $\mathbf{90.5}$ $49.6$ $84.4$ $\mathbf{67.6}$ $66.1$ $61.6$ $\mathbf{48.7}$ Cylinder3D [16] $57.0$ $88.5$ $39.9$ $58.0$ $58.4$ $48.1$ $68.6$ $77.0$ $\mathbf{0.5}$ $84.4$ $30.4$ $72.2$ $2.5$ $89.4$ $48.4$ $81.9$ $64.6$ $59.8$ $61.2$ $\mathbf{48.7}$ Fast-FRNet $62.4$ $90.8$ $41.0$ $\mathbf{66.8}$ $81.7$ $\mathbf{64.0}$ $70.5$ $84.2$ $0.0$ $91.3$ $35.1$ $78.3$ $0.0$ $89.4$ $64.4$ $85.0$ $65.5$ $69.5$ $60.5$ $47.7$ RangeFormer [36] $63.0$ $\mathbf{92.6}$ $\mathbf{51.6}$ $65.7$ $74.4$ $49.6$ $71.6$ $82.1$ $0.0$ $\mathbf{94.8}$ $\mathbf{44.4}$ $\mathbf{80.6}$ $\mathbf{11.4}$ $85.6$ $56.9$ $\mathbf{87.2}$ $64.1$ $\mathbf{77.0}$ $\mathbf{62.7}$ $45.1$ FRNet $\mathbf{63.1}$ $90.5$ $42.3$ $\mathbf{66.8}$ $\mathbf{82.4}$ $63.0$ $\mathbf{73.4}$ $\mathbf{86.2}$ $0.0$ $92.1$ $35.1$ $79.1$ $0.3$ $90.4$ $\mathbf{64.7}$ $85.4$ $65.8$ $70.7$ $61.8$ $48.6$
-TABLE X: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [64] leaderboard.
+TABLE XIV: The class-wise IoU scores of different LiDAR semantic segmentation approaches on the SemanticPOSS [68] leaderboard.
 All IoU scores are given in percentage (%).
@@
 Tab.
-VI shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [28].
+X shows the class-wise IoU scores among different LiDAR semantic segmentation methods on the test set of SemanticKITTI [32].
 We compare FRNet with different LiDAR representations, including point-view, range-view, sparse-voxel-view, and multi-view methods.
@@
 Tab.
-VII and Tab.
-VIII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [43], respectively.
+XI and Tab.
+XII show the class-wise IoU scores among different LiDAR semantic segmentation methods on the $val$ and $test$ set of nuScenes [47], respectively.
 The results demonstrate the great advantages of Fast-FRNet and FRNet.
@@
 As shown in Tab.
-IX, FRNet achieves state-of-the-art performance.
+XIII, FRNet achieves state-of-the-art performance.
 However, the weakly-annotated labels make the frustum-level pseudo labels largely covered by unannotated labels, which will limit performance improvement, especially for other-ground.
@@
 Tab.
-X summarizes the class-wise IoU scores among these methods and some other popular range-view methods.
+XIV summarizes the class-wise IoU scores among these methods and some other popular range-view methods.
 Results demonstrate that FRNet achieves state-of-the-art performance among both range-view and voxel-view methods.